### v2

# Processes $\gamma\gamma\rightarrow\phi_{i}\phi_{j}$ in Inert Higgs Doublet Models and Two Higgs Doublet Models

Abstract
In this paper, we present the results for one-loop induced processes $\gamma\gamma\rightarrow\phi_{i}\phi_{j}$ with CP-even Higgses $\phi_{i,j}\equiv h,\leavevmode\nobreak\ H$ at high energy photon-photon collision, within the frameworks of Inert Higgs Doublet Models and Two Higgs Doublet Models. Total cross-sections are shown as functions of center-of-mass energies. We find that the cross-sections for the computed processes in all the models under investigations are enhanced at around the threshold of singly charged Higgs pair ($\sim 2M_{H^{\pm}}$). Furthermore, the enhancement factors defined as the ratio of cross-sections of $\gamma\gamma\rightarrow\phi_{i}\phi_{j}$ in the investigated models over the corresponding ones for $\gamma\gamma\rightarrow hh$ in the Standard Model, are examined in the model’s parameter space. In the Inert Higgs Doublet Models, the factors are studied in the parameter space of $(M_{H^{\pm}},\leavevmode\nobreak\ \mu^{2}_{2})$ and $(M_{H^{\pm}},\leavevmode\nobreak\ \lambda_{2})$. In the Two Higgs Doublet Models, the factors are examined in the planes of $(M_{H^{\pm}},\leavevmode\nobreak\ t_{\beta})$ as well as in the space of charged Higgs mass $M_{H^{\pm}}$ and the soft-breaking $Z_{2}$ parameter $m_{12}^{2}$. Two scenarios of $c_{\beta-\alpha}>0$ and $c_{\beta-\alpha}<0$ are studied in further detail. The factors give a different behavior from considering these scenarios. As a result, discriminations for the above-mentioned scenarios can be performed at future colliders.
keywords:
Higgs phenomenology, one-loop corrections, analytic methods for quantum field theory, dimensional regularization.
1 Introduction
Measuring for scalar Higgs self-couplings including Standard Model-like Higgs trilinear- and quadratic-couplings as well as the couplings between scalar Higgses in many physics beyond the Standard Models (BSM) plays a key role for determining the scalar potential. We can subsequently answer for the origin of the electroweak symmetry breaking (EWSB). In this scheme, Higgs boson pair productions and multi-scalar Higgs productions should be measured precisely at future colliders. Recently, search for Higgs boson pair productions in the two bottom quarks associated with two photons, four bottom quarks, etc in the final states in proton-proton collisions have been performed at the Large Hadron Colliders (LHC) as in [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12]. It is well-known that the measurements for Higgs self-couplings are rather challenging at the LHC. The results from the study of [13, 14] show that the expected accuracy in the measurement of trilinear Higgs self-couplings would be about $20\%-30\%$ at the high luminosity of $3000$ fb-1. We know that physics the future lepton colliders (LC) will be complementary to the LHC in many aspects, as studied in [15]. Furthermore, the LC can significantly improve the LHC measurements in many cases and more important photon-photon collision is an option of the LC [16, 17] which the scalar Higgs pair productions ($\phi_{i}\phi_{j}$) can be measured via the channels $f\bar{f}\rightarrow f\bar{f}\gamma^{*}\gamma^{*}\rightarrow f\bar{f}\phi_{i}\phi_{j}$ for $f\equiv e,\mu$. In this aspect, the LC could open a window for probing new physics signals through multi-scalar Higgs productions.
From theoretical calculation side, one-loop corrections to double Higgs productions at the LHC have calculated in Standard Model (SM), the Higgs Extensions of the Standard Models (HESM), as well as other BSM frameworks by many groups. For examples, it is worth to mention typical works as in [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78], for further reviews in [79] and the related-references therein. Additionally, one-loop induced for $hh$ productions in the high-energy $\gamma\gamma$ collisions in the SM, the HESM and other BSMs have computed in Refs. [80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94]. Other computations for one-loop induced for Higgs boson pair productions in the linear lepton colliders including future multi-TeV muon colliders, have performed in Refs. [95, 96, 97, 98] and the additional references therein. Furthermore, double pseudo-scalar Higgs $A^{0}A^{0}$ at a $\gamma\gamma$ collider in the Two Higgs Doublet Model have reported in Ref. [99].
In this paper, we present the results for one-loop induced processes $\gamma\gamma\rightarrow\phi_{i}\phi_{j}$ with CP-even Higgses $\phi_{i,j}\equiv h,\leavevmode\nobreak\ H$ at high-energy $\gamma\gamma$ collision within the Higgs Extensions of the Standard Models (HESM) including the Inert Higgs Doublet Models the Two Higgs Doublet Models. A general analytic formulas for the process amplitudes derived in ’t Hooft-Feynman gauge (HF) are valid for a class of the above-mentioned HESM. Analytic results for the calculated processes are presented via the scalar Passarino-Veltman (PV) functions following the output of the packages LoopTools [100] and Collier [101]. The analytical expressions are tested by several numerical checks , e.g the ultraviolet finiteness, infrared finiteness of the one-loop amplitudes. Furthermore, the amplitudes also obey the ward identity which the indentity is also verified numerically in the works. Additionally, both the packages LoopTools and Collier are used for cross-checking for the final results before generating physical results. In phenomenological results, cross-sections are shown as functions of center-of-mass energies. Furthermore, the enhancement factors defined as the ratio of cross-sections of $\gamma\gamma\rightarrow\phi_{i}\phi_{j}$ in the HESMs over the corresponding ones for $\gamma\gamma\rightarrow hh$ in the Standard Model, are examined in parameter space of the models under consideration.
The paper is structured as follows. We review briefly the HESM in the next section. We then present calculations for $\gamma\gamma\rightarrow\phi_{i}\phi_{j}$ with CP-even Higgses $\phi_{i,j}\equiv h,H$ in the HESM in the section $3$. The phenomenological studies for the HESM are discussed in section $4$. In section $5$, conclusion and outlook for the paper are shown. In appendices $A,\leavevmode\nobreak\ B$ we derive the couplings appear in the calculations.
2 Higgs Extensions of the Standard Model
Two typical Higgs Extensions of the Standard Models are studied in this paper. The first model is to the Inert Higgs Doublet Models which are reviewed in next subsection $2.1$. We then discuss the Two Higgs Doublet Models in subsection $2.2$.
2.1 Inert Higgs Doublet Models
In the IHDM, an inert scalar $SU(2)_{L}$ doublet is included into the potential of the SM. The inert scalar particles will respond for dark matter candidates. For reviewing the theory and phenomenological examinations for the IHDM in concrete, we cite to the following papers [104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116]. The scalar potential of the IHDM read the general expression as follows:
$\mathcal{V}_{\textrm{IHDM}}(\phi_{1},\phi_{2})$ $=$ $\mu_{1}^{2}|\phi_{1}|^{2}+\mu_{2}^{2}|\phi_{2}|^{2}+\lambda_{1}|\phi_{1}|^{4}+\lambda_{2}|\phi_{2}|^{4}+\lambda_{3}|\phi_{1}|^{2}|\phi_{2}|^{2}+\lambda_{4}|\phi_{1}^{\dagger}\phi_{2}|^{2}$ (1)
$+\frac{\lambda_{5}}{2}\left\{(\phi_{1}^{\dagger}\phi_{2})^{2}+{\rm h.c}\right\}.$
The potential is conserved with respect to the so-called the global $Z_{2}$-symmetry, e.g. $\phi_{1}\leftrightarrow+\phi_{1}$, $\phi_{2}\leftrightarrow-\phi_{2}$. In this case, the scalar $\phi_{2}$ is odd and $\phi_{1}$, all particles in the SM are even under the $Z_{2}$-transformation. As mentioned, the $Z_{2}$-symmetry is unbroken after the EWSB, the field $\phi_{2}$ has the zero of vacuum expectation values (VEV). Otherwise, the field $\phi_{1}$ develops non-zero VEV ($v$). Two scalar fields are then expanded around VEV for the EWSB as follows:
$\phi_{1}$ $=$ $\left(\begin{array}[]{c}G^{\pm}\\ \frac{1}{\sqrt{2}}(v+h+iG^{0})\\ \end{array}\right),$ (4)
$\phi_{2}$ $=$ $\left(\begin{array}[]{c}H^{\pm}\\ \frac{1}{\sqrt{2}}(H+iA^{0})\\ \end{array}\right).$ (7)
Where the VEV is fixed at $v\sim 246$ GeV (as the SM case). The Goldstone bosons $G^{\pm},\;G^{0}$ are giving the masses for $W^{\pm}$ boson and $Z$ boson, respectively. There is no mixing between $h$ and $H$. The physical spectrum of the IHDM after the EWSB is consisted of three neutral scalar physical states, e.g two CP-even Higgses $h,\leavevmode\nobreak\ H$ and a CP-odd Higgs $A_{0}$. In further, we have pair of singly charged Higgs bosons $H^{\pm}$ in this model. In the spectrum, neutral scalar Higgs $h$ is identical to the SM-like Higgs boson discovered at the LHC. The masses of scalar bosons are calculated from the pare parameters as follows:
$M_{h}^{2}$ $=$ $-2\mu_{1}^{2}=2\lambda_{1}v^{2},$ (8)
$M_{H}^{2}$ $=$ $\mu_{2}^{2}+\frac{v^{2}}{2}\lambda_{L},$ (9)
$M_{A^{0}}^{2}$ $=$ $\mu_{2}^{2}+\frac{v^{2}}{2}\lambda_{R},$ (10)
$M_{H^{\pm}}^{2}$ $=$ $\mu_{2}^{2}+\frac{v^{2}}{2}\lambda_{3}.$ (11)
Where we have used $\lambda_{L/R}=\lambda_{3}+\lambda_{4}\pm\lambda_{5}$. As we mentioned in the above paragraphs, the global $Z_{2}$-symmetry is unbroken after the EWSB, The "inert" Higgs bosons like $H^{\pm}$, $A^{0}$ and $H$ have odd number under the $Z_{2}$-transformation. Subsequently, the "inert" Higgs bosons don’t couple to the SM particles. Therefore, the lightest neutral scalar bosons may be considered as dark matter candidates.
As a consequent of the unbroken of the $Z_{2}$-symmetry, the Yukawa Lagrangian of the IHDM must be the same as that of the SM. In detail, the Yukawa Lagrangian is expressed as follows:
$\mathcal{L}^{\textrm{IHDM}}_{\rm Yukawa}=-\sum_{f=u,d,\ell}g_{hff}\cdot h\bar{f}f+\cdots,$ (12)
where the Yukawa coupling is given $g_{hff}=m_{f}/v$ for fermion $f$. All the couplings involving to the computed processes $\gamma\gamma\rightarrow\phi_{i}\phi_{j}$ in the IHDM are listed in Table 1 (for all physical couplings) and Table 2 (for unphysical particles). We emphasize that the processes $\gamma\gamma\rightarrow hH$ are forbidden in the IHDM due to the $Z_{2}$-symmetry. Therefore, we have only $\phi_{i}\phi_{j}\equiv hh,\leavevmode\nobreak\ HH$ in this case. The detailed expressions for all the concerned-couplings are derived in the appendix $A$.
Vertices Notations Coupling
$hW_{\mu}W_{\nu}$ $g_{hWW}\cdot g_{\mu\nu}$ $i\left(\dfrac{2M_{W}^{2}}{v}\right)\cdot g_{\mu\nu}$
$hZ_{\mu}Z_{\nu}$ $g_{hZZ}\cdot g_{\mu\nu}$ $i\left(\dfrac{M_{Z}^{2}}{v}\right)\cdot g_{\mu\nu}$
$hH^{\pm}{H^{\mp}}$ $g_{hH^{\pm}H^{\mp}}$ $i\dfrac{2(\mu_{2}^{2}-M_{H^{\pm}}^{2})}{v}$
$Z_{\mu}H^{\pm}(p_{1})H^{\mp}(p_{2})$ $g_{ZH^{\pm}H^{\mp}}\cdot(p_{1}-p_{2})_{\mu}$ $i\left(\dfrac{M_{Z}c_{2W}}{v}\right)\cdot(p_{1}-p_{2})_{\mu}$
$A_{\mu}H^{\pm}(p_{1})H^{\mp}(p_{2})$ $g_{AH^{\pm}H^{\mp}}\cdot(p_{1}-p_{2})_{\mu}$ $i\left(\dfrac{M_{Z}s_{2W}}{v}\right)\cdot(p_{1}-p_{2})_{\mu}$
$hhh$ $g_{hhh}$ $-i\left(\dfrac{3M_{h}^{2}}{v}\right)$
$hHH$ $g_{hHH}$ $i\dfrac{2(\mu_{2}^{2}-M_{H}^{2})}{v}$
$H(p_{2})H^{\pm}(p_{1})W^{\mp}_{\mu}$ $g_{HH^{\pm}W}\cdot(p_{1}-p_{2})_{\mu}$$\pm i\left(\dfrac{M_{W}}{v}\right)\cdot(p_{1}-p_{2})_{\mu}$
$H^{\pm}H^{\mp}A_{\mu}A_{\nu}$ $g_{AAH^{\pm}H^{\mp}}\cdot g_{\mu\nu}$ $i\left(\dfrac{M_{Z}^{2}s_{2W}^{2}}{v^{2}}\right)\cdot g_{\mu\nu}$
$HH^{\pm}W^{\mp}_{\nu}A_{\mu}$ $g_{HH^{\pm}WA}\cdot g_{\mu\nu}$ $i\left(\dfrac{2M_{Z}^{2}c_{W}^{2}s_{W}}{v^{2}}\right)\cdot g_{\mu\nu}$
$hhH^{\pm}H^{\mp}$ $g_{hhH^{\pm}H^{\mp}}$ $i\dfrac{2(\mu_{2}^{2}-M_{H^{\pm}}^{2})}{v^{2}}$
$HHH^{\pm}H^{\mp}$ $g_{HHH^{\pm}H^{\mp}}$ $-2i\lambda_{2}$
$hhW^{\pm}_{\mu}W^{\mp}_{\nu}$ $g_{hhWW}\cdot g_{\mu\nu}$ $i\left(\dfrac{2M_{Z}^{2}c_{W}^{2}}{v^{2}}\right)\cdot g_{\mu\nu}$
$HHW^{\pm}_{\mu}W^{\mp}_{\nu}$ $g_{HHWW}\cdot g_{\mu\nu}$ $i\left(\dfrac{2M_{Z}^{2}c_{W}^{2}}{v^{2}}\right)\cdot g_{\mu\nu}$
Table 1: We list all couplings (physical couplings) contributing to $\gamma\gamma\rightarrow\phi_{i}\phi_{j}$ for the IHDM.
Vertices Notations Coupling
$A_{\mu}W^{\pm}_{\nu}G^{\mp}$ $g_{AW^{\pm}G^{\mp}}\cdot g_{\mu\nu}$ $i\left(\dfrac{2M_{Z}^{2}c_{W}^{2}s_{W}}{v}\right)\cdot g_{\mu\nu}$
$HH^{\pm}G^{\mp}$ $g_{HH^{\pm}G^{\mp}}$ $i\dfrac{M_{H^{\pm}}^{2}-M_{H}^{2}}{v}$
$hhG^{\pm}G^{\mp}$ $g_{hhG^{\pm}G^{\mp}}$ $-i\left(\dfrac{M_{h}^{2}}{v^{2}}\right)$
$HHG^{\pm}G^{\mp}$ $g_{HHG^{\pm}G^{\mp}}$ $i\dfrac{2(\mu_{2}^{2}-M_{H^{\pm}}^{2})}{v^{2}}$
$A_{\mu}A_{\nu}G^{\pm}G^{\mp}$ $g_{AAG^{\pm}G^{\mp}}\cdot g_{\mu\nu}$ $i\left(\dfrac{M_{Z}^{2}s_{2W}^{2}}{v^{2}}\right)\cdot g_{\mu\nu}$
$A_{\mu}G^{\pm}(p_{1})G^{\mp}(p_{2})$ $g_{AG^{\pm}G^{\mp}}\cdot(p_{1}-p_{2})_{\mu}$ $i\left(\dfrac{M_{Z}s_{2W}}{v}\right)\cdot(p_{1}-p_{2})_{\mu}$
Table 2: We list all vertices (unphysical couplings) in processes $\gamma\gamma\rightarrow\phi_{i}\phi_{j}$ in the IHDM.
The parameter space of the IHDM for our analysis is included as follows:
$\mathcal{P}_{\rm IHDM}=\{\mu_{2}^{2},\lambda_{2}^{2},M_{h}^{2}\sim 125.\textrm{GeV},M_{H}^{2},M_{A^{0}}^{2},M^{2}_{H^{\pm}}\}.$ (13)
We are going to review the current constraints on the physical parameter space in the IHDM given in Eq. 13. The constraints for the physical parameter space can be obtained by including the theoretical conditions and the experimental data. In the perspective of the experimental data, we take into account the Electroweak Precision Tests (EWPT) of the IHDM, dark matter search at the LHC, as well as including the LEP data. The topics have studied in Refs. [104, 105, 106, 107]. Additionally, the implications for loop-induced decays of the SM-like Higgs ($h$) to $V\gamma$ with $V\equiv Z,\gamma$ in the IHDM framework, e.g. decay process $h\rightarrow\gamma\gamma$ have reported in [108, 115, 116], and decay chanels $h\rightarrow Z\gamma$ have examined in [115, 116]. Furthermore, searching signals of the IHDM at future colliders have performed in Refs. [109, 110, 111, 112, 113, 114]. In the theoretical side, the most important theoretical constraints are obtained from the conditions that the models follow the tree-level unitarity, the vacuum stability, the perturbative regime. The theoretical constraints give the limitations on the Higgs self-couplings $\lambda_{i}$ for $i=1,2,\cdots,5$ and $\mu_{2}$. Taking theoretical and experimental constraints in the above-mentioned papers, one can select the parameter space for the IHDM as follows: we can take $5$ GeV $\leq M_{H}\leq 150$ GeV, $70$ GeV $\leq M_{H^{\pm}},M_{A^{0}}\leq 1000$ GeV, $|\mu_{2}|\leq 500$ GeV, and $0\leq\lambda_{2}\leq 8\pi$.
2.2 Two Higgs Doublet Models
We next to consider the second kind of the Higgs extension of the SM which is to the Two Higgs Doublet Models (THDM) in this paper. For reviewing the theory and the phenomenological studies for the THDM, we refer the paper Ref. [117] for further information. The model is summarized briefly in this section. An complex Higgs doublet possessing hypercharge $Y=1/2$ is added into the scalar sector of the SM. The scalar potential is read the form of
$\mathcal{V}_{\textrm{THDM}}(\phi_{1},\phi_{2})$ $=$ $m_{11}^{2}\phi_{1}^{\dagger}\phi_{1}+m_{22}^{2}\phi_{2}^{\dagger}\phi_{2}-\Big{[}m_{12}^{2}\phi_{1}^{\dagger}\phi_{2}+{\rm h.c.}\Big{]}+\frac{\lambda_{1}}{2}(\phi_{1}^{\dagger}\phi_{1})^{2}+\frac{\lambda_{2}}{2}(\phi_{2}^{\dagger}\phi_{2})^{2}$ (14)
$+\lambda_{3}(\phi_{1}^{\dagger}\phi_{1})(\phi_{2}^{\dagger}\phi_{2})+\lambda_{4}(\phi_{1}^{\dagger}\phi_{2})(\phi_{2}^{\dagger}\phi_{1})+\frac{1}{2}[\lambda_{5}\leavevmode\nobreak\ (\phi_{1}^{\dagger}\phi_{2})^{2}+\leavevmode\nobreak\ {\rm h.c}].$
In the present work, the CP-conserving case of the THDM is examined. Subsequently, the pare parameters in the above potential are set to be real parameters in this version. Additionally, the potential of the THDM follows the $Z_{2}$-symmetry, e.g. $\phi_{1}\leftrightarrow\phi_{1}$ and $\phi_{2}\leftrightarrow-\phi_{2}$, up to the soft breaking terms as $m_{12}^{2}\phi_{1}^{\dagger}\phi_{2}+{\rm h.c.}$ are added into the potential. Where the parameter $m_{12}^{2}$ is the breaking scale for the $Z_{2}$-symmetry.
Two scalar doublets are exanded around their VEVs for the EWSB as
$\phi_{k}$ $=$ $\begin{bmatrix}\rho_{k}^{+}\\ (v_{k}+\eta_{k}+i\xi_{k})/\sqrt{2}\end{bmatrix}\quad\textrm{for}\quad k=1,2.$ (15)
The vacuum expectation value is then fixed at $v=\sqrt{v_{1}^{2}+v_{2}^{2}}\sim 246$ GeV in agreement with the SM case. After the EWSB, the physical particles in the THDM include of two CP-even Higgs bosons $h$ and $H$ in which one of them $h$ being SM-like Higgs boson found at LHC, a CP-odd Higgs ($A^{0}$) boson, and a pair of charged Higgses ($H^{\pm}$). To obtain the physical masses for the new scalar bosons, we perform the following rotations
$\begin{pmatrix}\eta_{1}\\ \eta_{2}\end{pmatrix}$ $=$ $\begin{pmatrix}c_{\alpha}&-s_{\alpha}\\ s_{\alpha}&c_{\alpha}\end{pmatrix}\begin{pmatrix}H\\ h\end{pmatrix},$ (16)
$\begin{pmatrix}\rho_{1}^{\pm}\\ \rho_{2}^{\pm}\end{pmatrix}$ $=$ $\begin{pmatrix}c_{\beta}&-s_{\beta}\\ s_{\beta}&c_{\beta}\end{pmatrix}\begin{pmatrix}G^{\pm}\\ H^{\pm}\end{pmatrix},$ (17)
$\begin{pmatrix}\xi_{1}\\ \xi_{2}\end{pmatrix}$ $=$ $\begin{pmatrix}c_{\beta}&-s_{\beta}\\ s_{\beta}&c_{\beta}\end{pmatrix}\begin{pmatrix}G^{0}\\ A^{0}\end{pmatrix}.$ (18)
The mixing angle $\beta$ is given by $t_{\beta}\equiv\tan\beta=v_{2}/v_{1}$. The physical masses of Higgs bosons are then presented via the pare parameters as follows:
$M_{H^{\pm}}^{2}$ $=$ $M^{2}-\frac{1}{2}\lambda_{45}v^{2},$ (19)
$M_{A^{0}}^{2}$ $=$ $M^{2}-\lambda_{5}v^{2},$ (20)
$M_{h}^{2}$ $=$ $M_{11}^{2}s_{\beta-\alpha}^{2}+M_{22}^{2}c_{\beta-\alpha}^{2}+M_{12}^{2}s_{2(\beta-\alpha)},$ (21)
$M_{H}^{2}$ $=$ $M_{11}^{2}c_{\beta-\alpha}^{2}+M_{22}^{2}s_{\beta-\alpha}^{2}-M_{12}^{2}s_{2(\beta-\alpha)}$ (22)
where the parameter $M^{2}$ is used as $M^{2}=m_{12}^{2}/(s_{\beta}c_{\beta})$. The elements $M_{ij}$ for $i,j=1,2$ are given by
$M_{11}^{2}$ $=$ $(\lambda_{1}c_{\beta}^{4}+\lambda_{2}s_{\beta}^{4})v^{2}+\frac{v^{2}}{2}\;\lambda_{345}\;s_{2\beta}^{2},$ (23)
$M_{22}^{2}$ $=$ $M^{2}+\frac{v^{2}}{4}\Big{[}\lambda_{12}-2\lambda_{345}\Big{]}s_{2\beta}^{2},$ (24)
$M_{12}^{2}$ $=$ $M_{21}^{2}=-\frac{v^{2}}{2}\Big{[}\lambda_{1}c_{\beta}^{2}-\lambda_{2}s_{\beta}^{2}-\lambda_{345}\;c_{2\beta}\Big{]}s_{2\beta}.$ (25)
Here, the shorten notation have used as $\lambda_{ij\cdots}=\lambda_{i}+\lambda_{j}+\cdots$.
We show the couplings concerning in the amplitude computations for the processes $\gamma\gamma\rightarrow\phi_{i}\phi_{j}$ in Tables 34 (for physical couplings) and in Table 5 (for unphysical couplings). The couplings are derived in the appendix $B$.
Vertices Notations Couplings
$hW_{\mu}W_{\nu}$ $g_{hWW}\cdot g_{\mu\nu}$ $i\left(\dfrac{2M_{W}^{2}}{v}\;s_{\beta-\alpha}\right)\cdot g_{\mu\nu}$
$HW_{\mu}W_{\nu}$ $g_{HWW}\cdot g_{\mu\nu}$ $i\left(\dfrac{2M_{W}^{2}}{v}\;c_{\beta-\alpha}\right)\cdot g_{\mu\nu}$
$hZ_{\mu}Z_{\nu}$ $g_{hZZ}\cdot g_{\mu\nu}$ $i\left(\dfrac{2M_{Z}^{2}}{v}\;s_{\beta-\alpha}\right)\cdot g_{\mu\nu}$
$HZ_{\mu}Z_{\nu}$ $g_{HZZ}\cdot g_{\mu\nu}$ $i\left(\dfrac{2M_{Z}^{2}}{v}\;c_{\beta-\alpha}\right)\cdot g_{\mu\nu}$
$hH^{\pm}H^{\mp}$ $g_{hH^{\pm}H^{\mp}}$ $i\Bigg{[}\dfrac{c_{\alpha+\beta}(4M^{2}-3M_{h}^{2}-2M_{H^{\pm}}^{2})}{2vs_{2\beta}}$
                  $+\dfrac{(2M_{H^{\pm}}^{2}-M_{h}^{2})c_{(\alpha-3\beta)}}{2vs_{2\beta}}\Bigg{]}$
$Z_{\mu}H^{\pm}(p_{1})H^{\mp}(p_{2})$ $g_{ZH^{\pm}H^{\mp}}\cdot(p_{1}-p_{2})_{\mu}$ $i\left(\dfrac{M_{Z}\;c_{2W}}{v}\right)\cdot(p_{1}-p_{2})_{\mu}$
$A_{\mu}H^{\pm}(p_{1})H^{\mp}(p_{2})$ $g_{AH^{\pm}H^{\mp}}\cdot(p_{1}-p_{2})_{\mu}$ $i\left(\dfrac{M_{Z}\;s_{2W}}{v}\right)\cdot(p_{1}-p_{2})_{\mu}$
$hhh$ $g_{hhh}$ $i\dfrac{3e}{4M_{W}s_{W}s_{2\beta}}\Bigg{[}M^{2}c_{\alpha-3\beta}+(M^{2}-M_{h}^{2})c_{3\alpha-\beta}$
                                    $+(2M^{2}-3M_{h}^{2})c_{\alpha+\beta}\Bigg{]}$
$HHH$ $g_{HHH}$ $i\dfrac{3e}{4M_{W}s_{W}s_{2\beta}}\Bigg{[}M^{2}s_{\alpha-3\beta}+(M_{H}^{2}-M^{2})s_{3\alpha-\beta}$
                                    $+(2M^{2}-3M_{H}^{2})s_{\alpha+\beta}\Bigg{]}$
$hHH$ $g_{hHH}$ $i\dfrac{\Big{[}s_{2\alpha}(3M^{2}-M_{h}^{2}-2M_{H}^{2})+M^{2}s_{2\beta}\Big{]}}{v\;s_{2\beta}}s_{\alpha-\beta}$
$Hhh$ $g_{Hhh}$ $i\dfrac{\Big{[}s_{2\alpha}(3M^{2}-M_{H}^{2}-2m_{h}^{2})-M^{2}s_{2\beta}\Big{]}}{2vs_{2\beta}}c_{\alpha-\beta}$
Table 3: The physical couplings contributing to the considered processes $\gamma\gamma\rightarrow\phi_{i}\phi_{j}$ in the THDM.
Vertices Notations Couplings
$HH^{\pm}H^{\mp}$ $g_{HH^{\pm}H^{\mp}}$ $i\Bigg{[}\dfrac{s_{\alpha+\beta}(4M^{2}-3M_{H}^{2}-2M_{H^{\pm}}^{2})}{2v\;s_{2\beta}}$
                  $+\dfrac{(2M_{H^{\pm}}^{2}-M_{H}^{2})s_{\alpha-3\beta}}{2v\;s_{2\beta}}\Bigg{]}$
$H(p_{1})H^{\pm}(p_{2})W^{\mp}_{\mu}$ $g_{HH^{\pm}W}\cdot(p_{1}-p_{2})_{\mu}$$\pm i\left(\dfrac{M_{W}}{v}s_{\beta-\alpha}\right)\cdot(p_{1}-p_{2})_{\mu}$
$h(p_{1})H^{\pm}(p_{2})W^{\mp}_{\mu}$ $g_{hH^{\pm}W}\cdot(p_{1}-p_{2})_{\mu}$ $\mp i\left(\dfrac{M_{W}}{v}c_{\alpha-\beta}\right)\cdot(p_{1}-p_{2})_{\mu}$
$H^{\pm}H^{\mp}A_{\mu}A_{\nu}$ $g_{H^{\pm}H^{\mp}AA}\cdot g_{\mu\nu}$ $i\left(\dfrac{4M_{W}^{2}s_{W}^{2}}{v^{2}}\right)\cdot g_{\mu\nu}$
$HH^{\pm}W^{\mp}_{\mu}A_{\nu}$ $g_{HH^{\pm}WA}\cdot g_{\mu\nu}$ $i\left(\dfrac{2M_{W}^{2}s_{W}}{v^{2}}s_{\alpha-\beta}\right)\cdot g_{\mu\nu}$
$hH^{\pm}W^{\mp}_{\mu}A_{\nu}$ $g_{hH^{\pm}WA}\cdot g_{\mu\nu}$ $i\left(\dfrac{2M_{W}^{2}s_{W}}{v^{2}}c_{\alpha-\beta}\right)\cdot g_{\mu\nu}$
$hHH^{\pm}H^{\mp}$ $g_{HhH^{\pm}H^{\mp}}$ $i\lambda_{HhH^{\pm}H^{\mp}}$ [in Eq. (56)]
$HHH^{\pm}H^{\mp}$ $g_{HHH^{\pm}H^{\mp}}$ $i\lambda_{HHH^{\pm}H^{\mp}}$ [in Eq. (60)]
$hhH^{\pm}H^{\mp}$ $g_{hhH^{\pm}H^{\mp}}$ $i\lambda_{hhH^{\pm}H^{\mp}}$ [in Eq. (61)]
$hhW^{\pm}_{\mu}W^{\mp}_{\nu}$ $g_{hhWW}\cdot g_{\mu\nu}$ $i\left(\dfrac{4M_{W}^{2}}{v^{2}}\right)\cdot g_{\mu\nu}$
$HHW^{\pm}_{\mu}W^{\mp}_{\nu}$ $g_{HHWW}\cdot g_{\mu\nu}$ $i\left(\dfrac{4M_{W}^{2}}{v^{2}}\right)\cdot g_{\mu\nu}$
Table 4: Additional, the physical couplings contributing to the considered processes $\gamma\gamma\rightarrow\phi_{i}\phi_{j}$ in the THDM.
Vertices Notations Couplings
$A_{\mu}W^{\pm}_{\nu}G^{\mp}$ $g_{AW^{\pm}G^{\mp}}\cdot g_{\mu\nu}$ $i\left(\dfrac{2M_{W}^{2}s_{W}}{v}\right)\cdot g_{\mu\nu}$
$HH^{\pm}G^{\mp}$ $g_{HH^{\pm}G^{\mp}}$ $i\left(\dfrac{e}{2M_{W}s_{W}}s_{\alpha-\beta}\right)(M_{H^{\pm}}^{2}-M_{H}^{2})$
$A_{\mu}A_{\nu}G^{\pm}G^{\mp}$ $g_{AAG^{\pm}G^{\mp}}\cdot g_{\mu\nu}$ $i\left(\dfrac{4M_{W}^{2}s_{W}^{2}}{v^{2}}\right)\cdot g_{\mu\nu}$
$A_{\mu}G^{\pm}(p_{1})G^{\mp}(p_{2})$ $g_{AG^{\pm}G^{\mp}}\cdot(p_{1}-p_{2})_{\mu}$ $i\left(\dfrac{2M_{W}s_{W}}{v}\right)\cdot(p_{1}-p_{2})_{\mu}$
$hHG^{\pm}G^{\mp}$ $g_{hHG^{\pm}G^{\mp}}$ $i\lambda_{hHG^{\pm}G^{\mp}}$ [in Eq. (57)]
$hhG^{\pm}G^{\mp}$ $g_{hhG^{\pm}G^{\mp}}$ $i\lambda_{hhG^{\pm}G^{\mp}}$ [in Eq. (62) ]
$HHG^{\pm}G^{\mp}$ $g_{HHG^{\pm}G^{\mp}}$ $i\lambda_{HHG^{\pm}G^{\mp}}$ [in Eq. (63)]
Table 5: Unphysical couplings involving to the processes under investigations are shown.
Finally, we pay attention to the Yukawa sector in the THDM. In order to avoid Tree-level Flavor-Changing Neutral Currents (FCNCs), the discrete $Z_{2}$-symmetry may be proposed in the THDM as in [118]. The $Z_{2}$-parity assignments for all fermions and the definition for four types of the THDM based on the couple of the scalar with fermions are shown in [128]. The Yukawa Lagrangian is then written in the mass eigenstates as in [117]
${\mathcal{L}}_{\rm Yukawa}=-\sum_{f=u,d,\ell}\left(g_{hff}\cdot\bar{f}fh+g_{Hff}\cdot\bar{f}fH-ig_{A^{0}ff}\cdot\bar{f}\gamma_{5}fA^{0}\right)+\cdots,$ (26)
We show the Yukawa couplings of CP-even with fermions for four types of the THDM in Table 6, see [117, 129] for further detail.
Type $g_{huu}$ $g_{hdd}$ $g_{h\ell\ell}$ $g_{Huu}$ $g_{Hdd}$ $g_{H\ell\ell}$
I $\dfrac{m_{u}}{\sqrt{2}v}\dfrac{c_{\alpha}}{s_{\beta}}$ $\dfrac{m_{d}}{\sqrt{2}v}\dfrac{c_{\alpha}}{s_{\beta}}$ $\dfrac{m_{\ell}}{\sqrt{2}v}\dfrac{c_{\alpha}}{s_{\beta}}$ $\dfrac{m_{u}}{\sqrt{2}v}\dfrac{s_{\alpha}}{s_{\beta}}$ $\dfrac{m_{d}}{\sqrt{2}v}\dfrac{s_{\alpha}}{s_{\beta}}$ $\dfrac{m_{\ell}}{\sqrt{2}v}\dfrac{s_{\alpha}}{s_{\beta}}$
II $\dfrac{m_{u}}{\sqrt{2}v}\dfrac{c_{\alpha}}{s_{\beta}}$ $-\dfrac{m_{d}}{\sqrt{2}v}\dfrac{s_{\alpha}}{c_{\beta}}$ $-\dfrac{m_{\ell}}{\sqrt{2}v}\dfrac{s_{\alpha}}{c_{\beta}}$ $\dfrac{m_{u}}{\sqrt{2}v}\dfrac{s_{\alpha}}{s_{\beta}}$ $-\dfrac{m_{d}}{\sqrt{2}v}\dfrac{c_{\alpha}}{c_{\beta}}$ $-\dfrac{m_{\ell}}{\sqrt{2}v}\dfrac{c_{\alpha}}{c_{\beta}}$
X $\dfrac{m_{u}}{\sqrt{2}v}\dfrac{c_{\alpha}}{s_{\beta}}$ $\dfrac{m_{d}}{\sqrt{2}v}\dfrac{c_{\alpha}}{s_{\beta}}$ $-\dfrac{m_{\ell}}{\sqrt{2}v}\dfrac{s_{\alpha}}{c_{\beta}}$ $\dfrac{m_{u}}{\sqrt{2}v}\dfrac{s_{\alpha}}{s_{\beta}}$ $\dfrac{m_{d}}{\sqrt{2}v}\dfrac{s_{\alpha}}{s_{\beta}}$ $-\dfrac{m_{\ell}}{\sqrt{2}v}\dfrac{c_{\alpha}}{c_{\beta}}$
Y $\dfrac{m_{u}}{\sqrt{2}v}\dfrac{c_{\alpha}}{s_{\beta}}$ $-\dfrac{m_{d}}{\sqrt{2}v}\dfrac{s_{\alpha}}{c_{\beta}}$ $\dfrac{m_{\ell}}{\sqrt{2}v}\dfrac{c_{\alpha}}{s_{\beta}}$ $\dfrac{m_{u}}{\sqrt{2}v}\dfrac{s_{\alpha}}{s_{\beta}}$ $-\dfrac{m_{d}}{\sqrt{2}v}\dfrac{c_{\alpha}}{c_{\beta}}$ $\dfrac{m_{\ell}}{\sqrt{2}v}\dfrac{s_{\alpha}}{s_{\beta}}$
Table 6: We show all the Yukawa couplings of CP-even Higges to fermions for all types of THDMs.
The parameter space $\mathcal{P}_{\rm THDM}$ for THDM is include as follows
$\mathcal{P}_{\rm THDM}=\{M_{h}^{2}\sim 125.\textrm{GeV},M_{H}^{2},M_{A^{0}}^{2},M^{2}_{H^{\pm}},m_{12}^{2},t_{\beta},s_{\beta-\alpha}\}.$ (27)
[…]
3 One-loop corrections to $\gamma\gamma\rightarrow\phi_{i}\phi_{j}$ with CP-even Higgses $\phi_{i,j}\equiv h,H$ in HESM
The calculations are performed with the help of FeynArts/FormCalc/LoopTools/Collider packages [102, 100, 101]. We first implement the above-mentioned HESMs into FeynArts [102] model. All one-loop diagrams for the computed processes are then generated automatically by using the program. In the frameworks of the HESM under investigations, all one-loop induced Feynman diagrams for $\gamma\gamma\rightarrow\phi_{i}\phi_{j}$ are listed in the following paragraphs. The computations are performed in the ’t Hooft-Feynman gauge (HF), all Feynman diagrams can be categorized into several groups listed in next paragraphs. We first mention one-loop induced Feynman diagrams with $\phi_{k}^{*}$-poles for $\phi_{k}^{*}=h^{*},H^{*}$. These topologies are connected loop-induced processes $\phi_{k}^{*}\rightarrow\gamma\gamma$ with the vertices of $\phi_{k}^{*}\phi_{i}\phi_{j}$, as plotted in Fig. 1. In Fig. 1, all fermions exchanging in the loop are included in the group $G_{1}$. Feynman diagrams with vector $W$-boson propagating in the loop are concerned. Within the HF gauges, all charged Goldstone $G^{\pm}$ and Ghost particles exchanging in the loop are also taken into consideration this case. These diagrams are putted into the group $G_{2}$. We also have singly charged Higgs appear in the models under concern which they are also exchanged in the loop of diagrams, as noted as the group $G_{3}$.
Figure 1: One-loop induced Feynman diagrams with $\phi_{k}^{*}$-poles for $\phi_{k}^{*}=h^{*},H^{*}$.
One-loop box diagrams contributing to the computed processes are next discussed. Depend on the kind of particles propagating in the loop, we list the box diagrams into several groups as follows. In the group $G_{4}$, as plotted in Fig. 2, all fermions in the loop are concerned.
Figure 2: Box diagrams with fermions in the loop involving to the calculated processes in the HESM.
Additionally, conindering vector $W$-boson, charged Goldstone bosons and Ghost particles propagating in the loop of one-loop box diagrams, as plotted in Figs. 34, are also contributed to the processes under consideration. We classify these diagrams into the group $G_{5}$.
Figure 3: Vector $W$-boson, charged Goldstone bosons and Ghost particles propagating in the loop of one-loop box diagrams.
Figure 4: Vector $W$-boson, charged Goldstone bosons and Ghost particles propagating in the loop of one-loop box diagrams.
Within the frameworks of the HESMs under investigations in the paper, we also have one-loop box diagrams with mixing of vector $W$-boson and charged Higgs in the internal lines, seen Figs. 5,  6. We classify these diagrams into group $G_{6}$.
Figure 5: One-loop box diagrams with both vector $W$-boson, charged Goldstone bosons, Ghost particles and charged Higgs propagating in the loop.
Figure 6: Mixing of vector $W$-boson and charged Higgs in the internal lines of one-loop box diagrams.
Finally, we consider all one-loop box diagrams with singly charged Higgs in the loop, as shown in Fig. 7. These diagrams are then putted into group $G_{7}$.
Figure 7: One-loop box diagrams with charged Higgs in the loop propagating in the loop.
In general, one-loop amplitude for scattering processes $\gamma_{\mu}(q_{1})\,\gamma_{\nu}(q_{2})\rightarrow\phi_{i}(q_{3})\,\phi_{j}(q_{4})$ is expressed in terms of Lorentz structure as follows:
$\mathcal{A}_{\gamma\gamma\rightarrow\phi_{i}\phi_{j}}$ $=$ $\Big{[}F_{00}\,g^{\mu\nu}+F_{12}\,q_{1}^{\nu}q_{2}^{\mu}+F_{13}\,q_{1}^{\nu}q_{3}^{\mu}+F_{23}\,q_{2}^{\mu}q_{3}^{\nu}+F_{33}\,q_{3}^{\mu}q_{3}^{\nu}\Big{]}\varepsilon_{\mu}(q_{1})\varepsilon_{\nu}(q_{2}).$ (28)
In this formulas, the vector $\varepsilon_{\mu}(q)$ is polarization vector of external photon with the $4$-dimension momentum $q$. The invariant masses are given: $q_{1}^{2}=q_{2}^{2}=0,\quad q_{3}^{2}=M_{\phi_{i}}^{2},\quad q_{4}^{2}=M_{\phi_{j}}^{2}$. The factors $F_{ij}$ for $i,j=1,2,3$ are called as one-loop form factors hereafter. They are presented as functions of the following kinematic invariant variables:
$s$ $=$ $(q_{1}+q_{2})^{2}=q_{1}^{2}+2q_{1}\cdot q_{2}+q_{2}^{2}=2q_{1}\cdot q_{2},$ (29)
$t$ $=$ $(q_{1}-q_{3})^{2}=q_{1}^{2}-2q_{1}\cdot q_{3}+q_{3}^{2}=M_{\phi_{i}}^{2}-2q_{1}\cdot q_{3},$ (30)
$u$ $=$ $(q_{2}-q_{3})^{2}=q_{2}^{2}-2q_{2}\cdot q_{3}+q_{3}^{2}=M_{\phi_{i}}^{2}-2q_{2}\cdot q_{3}.$ (31)
The kinematic variables obey the below relation as: $s+t+u=M_{\phi_{i}}^{2}+M_{\phi_{j}}^{2}$. Due to the on-shell photons in initial states, the amplitude follows the ward identity. Subsequently, the above-mentioned factors are related to each others as
$F_{00}$ $=$ $\dfrac{t-M_{\phi_{i}}^{2}}{2}\,F_{13}-\dfrac{s}{2}\,F_{12},$ (32)
$F_{00}$ $=$ $\dfrac{u-M_{\phi_{i}}^{2}}{2}\,F_{23}-\dfrac{s}{2}\,F_{12},$ (33)
$F_{13}$ $=$ $\dfrac{u-M_{\phi_{i}}^{2}}{s}\,F_{33},$ (34)
$F_{23}$ $=$ $\dfrac{t-M_{\phi_{i}}^{2}}{s}\,F_{33}.$ (35)
With the help of the mentioned relations, one-loop amplitude is written by two independent one-loop form factors, e.g $F_{12}$ and $F_{33}$ as chosen in this work. The one-loop amplitude can be finally presented in form of:
$\mathcal{A}_{\gamma\gamma\rightarrow\phi_{i}\phi_{j}}$ $=$ $\Bigg{\{}\Bigg{[}\dfrac{(M_{\phi_{i}}^{2}-t)(M_{\phi_{i}}^{2}-u)}{2s}\cdot g^{\mu\nu}+q_{3}^{\mu}q_{3}^{\nu}+\dfrac{(t-M_{\phi_{i}}^{2})}{s}\cdot q_{2}^{\mu}q_{3}^{\nu}$
$\hskip 14.22636pt+\dfrac{(u-M_{\phi_{i}}^{2})}{s}\cdot q_{3}^{\mu}q_{1}^{\nu}\Bigg{]}\cdot F_{33}+\Big{[}q_{2}^{\mu}q_{1}^{\nu}-\dfrac{s}{2}\cdot g^{\mu\nu}\Big{]}\cdot F_{12}\Bigg{\}}\,\varepsilon_{\mu}(q_{1})\varepsilon_{\nu}(q_{2}).$
From the FeynArts program, we can generate the amplitude automatically which is expressed in term of one-loop tensor integrals. The mentioned tensor integrals appear in the production amplitude can be reduced into the scalar PV-functions by using the FormCalc package [103]. Finally, we collect all factors $F_{12}$ and $F_{33}$ presenting via the scalar PV-functions following the output of the packages LoopTools [100] and Collier [101].
The form factors $F_{ab}$ for $ab={12,33}$ are divided into several parts which are corresponding to the contributions from one-loop triangle and one-loop box diagrams given in the above-paragraphs. In detail, the form factors are decomposed as follows:
$F_{12}$ $=$ $\sum\limits_{\phi_{k}^{*}=h^{*},H^{*}}\dfrac{g_{\phi_{k}^{*}\phi_{i}\phi_{j}}}{\Big{[}s-M_{\phi_{k}}^{2}+i\Gamma_{\phi_{k}}M_{\phi_{k}}\Big{]}}\times$
×$\times\Bigg{\{}\sum\limits_{f}g_{\phi_{k}^{*}ff}\cdot F_{12,f}^{\text{Trig}}+g_{\phi_{k}^{*}WW}\cdot F_{12,W}^{\text{Trig}}+g_{\phi_{k}^{*}SS}\cdot F_{12,S}^{\text{Trig}}\Bigg{\}}$
$+\sum\limits_{f}g_{\phi_{i}ff}\cdot g_{\phi_{j}ff}\cdot F_{12,f}^{\text{Box}}$
+$+\Big{\{}g_{\phi_{i}WW}\cdot g_{\phi_{j}WW}\cdot F_{12,W}^{\text{Box, 1}}+g_{\phi_{i}\phi_{j}WW}\cdot F_{12,W}^{\text{Box, 2}}+g_{\phi_{i}\phi_{j}GG}\cdot F_{12,W}^{\text{Box, 3}}\Big{\}}$
+$+\Big{\{}g_{\phi_{i}SS}\cdot g_{\phi_{j}SS}\cdot F_{12,S}^{\text{Box},1}+g_{\phi_{i}\phi_{j}SS}\cdot F_{12,S}^{\text{Box},2}\Big{\}}$
$+g_{\phi_{i}SW}\cdot g_{\phi_{j}SW}\cdot F_{12,WS}^{\text{Box}},$
$F_{33}$ $=$ $\sum\limits_{f}g_{\phi_{i}ff}\cdot g_{\phi_{j}ff}\cdot F_{33,f}^{\text{Box}}+g_{\phi_{i}WW}\cdot g_{\phi_{j}WW}\cdot F_{33,W}^{\text{Box}}$
$+g_{\phi_{i}SS}\cdot g_{\phi_{j}SS}\cdot F_{33,S}^{\text{Box}}+g_{\phi_{i}SW}\cdot g_{\phi_{j}SW}\cdot F_{33,WS}^{\text{Box}},$
where $S\equiv H^{\pm}$ in this case. In further detail, the first part of form factors $F_{12}$ are calculated from one-loop diagrams of $\phi_{k}^{*}\rightarrow\gamma\gamma$ in connecting with the vertices $\phi_{k}^{*}\phi_{i}\phi_{j}$. These factors are presented in terms of each factor in the bracket, e.g. $F_{12,f}^{\text{Trig}}$ (contributed from the group $G_{1}$ in Fig. 1), $F_{12,W}^{\text{Trig}}$ (evaluated from $G_{2}$ in Fig. 1), $F_{12,S}^{\text{Trig}}$ (attributed from $G_{3}$ in Fig. 1). While the factors $F_{12,f}^{\text{Box}}$ are calculated from the fermion box diagrams in Fig. 2. Another form factors computed from one-loop $W$ box diagrams as in Fig. 3, are divided into the following parts, e.g. $F_{12,W}^{\text{Box, k}}$ for $k=1,2,3$ which are corresponding to the factors factorized out by general trilinear-couplings of $\phi_{i}WW,\leavevmode\nobreak\ \phi_{j}WW$, quadratic-couplings of $\phi_{i}\phi_{j}WW$ and $\phi_{i}\phi_{j}GG$ as in Eq. 3. We have also expressed the factors attributing from one-loop charged Higgs box diagrams into two sub-factors $F_{12,S}^{\text{Box},k}$ for $k=1,2$ which are factorized out by general trilinear-couplings of $\phi_{i}SS,\leavevmode\nobreak\ \phi_{j}SS$, quadratic-couplings of $\phi_{i}\phi_{j}WW$ and $\phi_{i}\phi_{j}SS$ as in Eq. 3. Lastly, from diagrams with mixing $W$ boson and charged Higgs in internal lines, we have the factors $F_{12,WS}^{\text{Box}}$ which can be factorized out in term of the trilinear-couplings $\phi_{i}SW,\leavevmode\nobreak\ \phi_{j}SW$. Otherwise, the form factors $F_{33}$ are only contributed from one-loop box diagrams. They can be factorized out in term of general couplings as in Eq. 3.
The factors $F_{12}$ and $F_{33}$ for the processes $\gamma\gamma\rightarrow hh$ in the HESM can be reduced to the results in the SM. For this case, we set all general couplings related to the HESM as $\phi_{k}SS,\leavevmode\nobreak\ \phi_{k}WS,\leavevmode\nobreak\ \phi_{i}\phi_{j}SS$ being zero and all other couplings is set back to the SM cases. Due to the $Z_{2}$-symmetry, the processes $\gamma\gamma\rightarrow hH$ in the IHDM are forbidden in this case. Reduction for factors $\gamma\gamma\rightarrow hh,\leavevmode\nobreak\ HH$ in the IHDM can be performed by setting appropriately the general couplings to the IHDM case. We can apply the same strategies for getting all one-loop form factors in the THDM. Having all the necessary factors for the computed processes, the tests for the calculations are performed, e.g. the ultraviolet finiteness, infrared finiteness of the one-loop amplitudes. Furthermore, the amplitudes also obey the so-called ward identity. This identity can be verified as follows. We collect all form factors $F_{00},\leavevmode\nobreak\ F_{12},\cdots$ independently and we then confirm the indentities from Eq. 32 to  35 numerically. Additionally, both the packages LoopTools and Collier are used for cross-checking for the final results. We skip showing the numerical results for the tests in this paper. For these topics, we refer our previous paper [129] for examples.
The cross-sections are then calculated as follows
[…]
with $n=2$ if the final particles are identical such as $\gamma\gamma\rightarrow hh,\leavevmode\nobreak\ HH$, and 1 otherwise like $\gamma\gamma\rightarrow hH$. The integration limits are
$t_{\text{min}(\text{max})}$ $=$ $-\dfrac{s}{2}\Bigg{\{}1-\dfrac{M_{\phi_{i}}^{2}+M_{\phi_{j}}^{2}}{s}\pm\Bigg{[}1-2\Bigg{(}\dfrac{M_{\phi_{i}}^{2}+M_{\phi_{j}}^{2}}{s}\Bigg{)}+\Bigg{(}\dfrac{M_{\phi_{i}}^{2}-M_{\phi_{j}}^{2}}{s}\Bigg{)}^{2}\Bigg{]}^{1/2}\Bigg{\}}.$ (40)
The unpolarized amplitude is given
$\frac{1}{4}\sum\limits_{\text{unpol.}}\big{|}\mathcal{A}_{\gamma\gamma\rightarrow\phi_{i}\phi_{j}}\big{|}^{2}$ $=$ $\dfrac{M_{\phi_{i}}^{4}s^{2}+(M_{\phi_{i}}^{2}M_{\phi_{j}}^{2}-tu)^{2}}{8s^{2}}\,\big{|}F_{33}\big{|}^{2}$
$-\dfrac{M_{\phi_{i}}^{2}s}{4}\,\mathcal{R}e\Big{[}F_{33}\cdot\big{(}F_{12}\big{)}^{*}\Big{]}$
$+\dfrac{s^{2}}{8}\,\big{|}F_{12}\big{|}^{2}.$
In phenomenological analyses, we are interested in examining the enhancement factors $\mu_{\phi_{i}\phi_{j}}^{\textrm{NP}}$ with NP standing for the THDM and the IHDM, accordingly, defined as the ratio of cross-sections of $\gamma\gamma\rightarrow\phi_{i}\phi_{j}$ in the HESM over the corresponding ones for $\gamma\gamma\rightarrow hh$ in the SM. The factors are given explicitly by
$\mu_{\phi_{i}\phi_{j}}^{\textrm{NP}}=\frac{\hat{\sigma}^{\textrm{NP}}_{\gamma\gamma\rightarrow\phi_{i}\phi_{j}}}{\hat{\sigma}^{\textrm{SM}}_{\gamma\gamma\rightarrow hh}}(\mathcal{P}_{\textrm{NP}}).$ (42)
In this work, the enhancement factors are examined in the parameter space of the THDM and the IHDM.
4 Phenomenological results
For phenomenological results, all physical input parameters in the SM are given the same as in [128, 129]. Scanning parameters for each HESM will be shown in the next subsections.
4.1 IHDM
Phenomenological studies for the processes $\gamma\gamma\rightarrow\phi_{i}\phi_{j}$ in the IHDM are presented in this subsection. In the IHDM, the process $\gamma\gamma\rightarrow hH$ is forbidden by the $Z_{2}$-symmetry. For this reason, we only concern physical results for the processes $\gamma\gamma\rightarrow hh,\leavevmode\nobreak\ HH$ in the IHDM.
4.1.1 Production cross-sections
In Fig. 8, we show cross-sections for $\gamma\gamma\rightarrow hh,\leavevmode\nobreak\ HH$ in the IHDM, together with the ones for $hh$ production in the SM, as functions of center-of-mass energy (CoM, or $\sqrt{\hat{s}_{\gamma\gamma}}$). For the generated data, we select the following parameter space in the IHDM as $M_{H^{\pm}}=200$ GeV, $M_{H}=150$ GeV and fix $\lambda_{2}=0.8$ for all cases. We vary $350$ GeV $\leq\sqrt{\hat{s}_{\gamma\gamma}}\leq 1500$ GeV in the plots. Cross-sections are presented for $\mu_{2}^{2}=0$ GeV2 on the left panel and for $200^{2}$ GeV2 on the right panel, respectively. In the plots, the black line shows for cross-sections for $\gamma\gamma\rightarrow hh$ in the SM and the blue (green) line presents for $\gamma\gamma\rightarrow hh\leavevmode\nobreak\ (HH)$ in THDM, respectively.
We first comment on the results in the case of $\mu_{2}^{2}=0$ GeV2. The cross-sections for $hh,\leavevmode\nobreak\ HH$ in the IHDM have peaks at $\sqrt{\hat{s}_{\gamma\gamma}}\sim 2M_{H^{\pm}}=400$ GeV. In the regions $\sqrt{\hat{s}_{\gamma\gamma}}\leq 750$ GeV, $\hat{\sigma}_{hh,HH}$ in the IHDM are larger than $\hat{\sigma}_{hh}$ in the SM. Beyond the regions of $\sqrt{\hat{s}_{\gamma\gamma}}\geq 750$ GeV, the cross-sections for $HH$ in the IHDM are suppressed in comparison with $hh$ productions in both the SM and the IHDM. It is interested in finding that the production cross-sections for $\gamma\gamma\rightarrow hh,\leavevmode\nobreak\ HH$ in the IHDM are dominant around the peaks compared with $\hat{\sigma}_{hh}$ in the SM. It indicates that the contributions from singly charged Higgs in the loop of $\gamma\gamma\rightarrow hh,\leavevmode\nobreak\ HH$ are massive attributions in the regions.
In the case of $\mu_{2}^{2}=200^{2}$ GeV2, we only observe a peak of $\hat{\sigma}_{HH}$ in the IHDM around $\sqrt{\hat{s}_{\gamma\gamma}}\sim 2M_{H^{\pm}}=400$ GeV. In further, the data shows that the cross-sections for $HH$ production are dominant in the regions $\sqrt{\hat{s}_{\gamma\gamma}}\leq 750$ GeV in contrast with the corresponding ones for $hh$ productions in both the SM and the IHDM. Beyond the regions $\sqrt{\hat{s}_{\gamma\gamma}}\geq 750$ GeV, $\hat{\sigma}_{HH}$ are suppressed. It is observed that the cross-sections for $hh$ production in the IHDM are smaller than $\hat{\sigma}_{hh}$ in the SM when $\sqrt{\hat{s}_{\gamma\gamma}}\leq 550$ GeV. In the regions of $\sqrt{\hat{s}_{\gamma\gamma}}\geq 550$ GeV, $\hat{\sigma}_{hh}$ in the IHDM tend to the cross-sections for $hh$ production in the SM. This can be explained as follows. Since the $hh$ productions in the IHDM are different from the ones in the SM by the contributions of charged Higgs in the loop of triangle $h^{*}$-pole and box diagrams. These contributions depend on $M_{H^{\pm}}$ and the vertices $hH^{\pm}H^{\mp}$, $hhH^{\pm}H^{\mp}$ expressing in terms of $\mu_{2}^{2}$. At the large value of $\mu_{2}^{2}$ these contributions may be cancelled out. As a result, cross-sections for $hh$ productions in the IHDM tend to the corresponding ones in the SM. Another case of $HH$ production, we have no couplings of $HH^{\pm}H^{\mp}$ due to the $Z_{2}$-symmetry and the vertex $HHH^{\pm}H^{\mp}$ depend on $\lambda_{2}$. Therefore, we have no such cancellations as mentioned. It is reasonable that the cross-sections for $HH$ production in the IHDM are dominant in the regions $\sqrt{\hat{s}_{\gamma\gamma}}\leq 550$ GeV and they also have peak at $\sqrt{\hat{s}_{\gamma\gamma}}\sim 2M_{H^{\pm}}=400$ GeV.
$\hat{\sigma}_{\phi_{i}\phi_{j}}(\mu_{2}^{2}=0)$ [pb] $\hat{\sigma}_{\phi_{i}\phi_{j}}(\mu_{2}^{2}=200^{2})$ [pb]
Refer to caption Refer to caption
                                                     $\sqrt{\hat{s}_{\gamma\gamma}}$ [GeV]                                                      $\sqrt{\hat{s}_{\gamma\gamma}}$ [GeV]
Figure 8: Total cross-sections for $\gamma\gamma\rightarrow hh,\leavevmode\nobreak\ HH$ in the SM and IHDM are presented as functions of $\sqrt{\hat{s}_{\gamma\gamma}}$. For the generated data, we select $M_{H^{\pm}}=200$ GeV, $M_{H}=150$ GeV in this case. We vary $350$ GeV $\leq\sqrt{\hat{s}_{\gamma\gamma}}\leq 1500$ GeV. In the below plots, we fix $\mu_{2}^{2}=0,200^{2}$ GeV2 and $\lambda_{2}=0.8$ for all cases. In the plots, the black line shows for $\hat{\sigma}_{hh}$ in the SM. Additionally, the blue (green) line presents for $\gamma\gamma\rightarrow hh\leavevmode\nobreak\ (HH)$ in THDM.
4.1.2 Enhancement factors
The enhancement factors given in Eq. 42 are examined in the IHDM. In Fig. 9, the factors for $\gamma\gamma\rightarrow hh,\leavevmode\nobreak\ HH$ are scanned in the parameter space of $M_{H^{\pm}},\leavevmode\nobreak\ \mu_{2}^{2}$. In the following scatter plots, singly charged Higgs masses are varied from $70$ GeV $\leq M_{H^{\pm}}\leq 600$ GeV and $-200$ GeV $\leq\mu_{2}\leq 200$ GeV. Furthermore, we fix $\lambda_{2}=0.8$ and $M_{H}=150$ GeV for all cases. The data is generated at $\sqrt{\hat{s}_{\gamma\gamma}}=500$ (all left panel scatter-plots) GeV and at $\sqrt{\hat{s}_{\gamma\gamma}}=1000$ GeV (all right panel scatter-plots).
The factors for $hh$ productions in the IHDM are first analyzed. Since the cross-sections for $hh$ productions are enhanced around the peaks at $\sqrt{\hat{s}_{\gamma\gamma}}\sim 2M_{H^{\pm}}$. Therefore, it is not surprised in finding that $\mu_{hh}^{\textrm{IHDM}}$ becomes largest at $\sqrt{\hat{s}_{\gamma\gamma}}\sim 2M_{H^{\pm}}=250$ GeV (for the left plots) and at $\sqrt{\hat{s}_{\gamma\gamma}}\sim 2M_{H^{\pm}}=500$ GeV (for the right plots). Around these peaks, the data indicates that the enhancement factors tend to about $\sim 1.5$ in the limit of $\mu_{2}^{2}\rightarrow M_{H^{\pm}}^{2}$. Since the contributions of charged Higgs in the loop being small when $\mu_{2}^{2}\rightarrow M_{H^{\pm}}^{2}$ (due to the fact that the couplings of $hH^{\pm}H^{\mp}$, $hhH^{\pm}H^{\mp}$ tend to zero in this limit). It is found that the enhancement factors can reach to factor of $6$ (for $500$ GeV of CoM) and factor of $13$ (for $1000$ GeV of CoM) around the peaks. Beyond the peaks, we observe that $1\leq\mu_{hh}^{\textrm{IHDM}}\leq 2$.
For the enhancement factors of $HH$ productions in the IHDM, we also find $\mu_{HH}^{\textrm{IHDM}}$ becomes largest contributions at $\sqrt{\hat{s}_{\gamma\gamma}}\sim 2M_{H^{\pm}}=250$ GeV (for the left plots) and at $\sqrt{\hat{s}_{\gamma\gamma}}\sim 2M_{H^{\pm}}=500$ GeV (for the right plots). It is attentive to realize that $\mu_{HH}^{\textrm{IHDM}}$ have different behavior in comparison with $\mu_{hh}^{\textrm{IHDM}}$. At the $500$ GeV of CoM, the factors develop to the peak and then are decreased rapidly beyond the peak. However, they grow up with the charged Higgs masses in the above regions of $M_{H^{\pm}}\geq\sim 300$ GeV. Because there aren’t couplings $HH^{\pm}H^{\mp}$ due to the $Z_{2}$-symmetry and the vertex $HHH^{\pm}H^{\mp}$ only depends on $\lambda_{2}$. Therefore, in the high regions of charged Higgs masses, the factors $\mu_{HH}^{\textrm{IHDM}}$ are increased with $M_{H^{\pm}}$.
$\mu_{hh}^{\textrm{IHDM}}$ $\mu_{hh}^{\textrm{IHDM}}$
Refer to caption Refer to caption
                                            $M_{H^{\pm}}$ [GeV]                                             $M_{H^{\pm}}$ [GeV]
$\mu_{HH}^{\textrm{IHDM}}$ $\mu_{HH}^{\textrm{IHDM}}$
Refer to caption Refer to caption
                                            $M_{H^{\pm}}$ [GeV]                                             $M_{H^{\pm}}$ [GeV]
Figure 9: The enhancement factors are presented in the parameter space of $M_{H^{\pm}},\leavevmode\nobreak\ \mu_{2}^{2}$. Charged Higgs masses are varied as $70$ GeV $\leq M_{H^{\pm}}\leq 1000$ GeV and $-200$ GeV $\leq\mu_{2}\leq 200$ GeV. We fix $\lambda_{2}=0.8$ and $M_{H}=150$ GeV for all cases. In the plots, we set $\sqrt{\hat{s}_{\gamma\gamma}}=500$ GeV (for the left panel-plots) and $\sqrt{\hat{s}_{\gamma\gamma}}=1000$ GeV (for the right panel-plots).
In Fig. 10, the enhancement factors for $\gamma\gamma\rightarrow hh,\leavevmode\nobreak\ HH$ are generated in the space of $M_{H^{\pm}},\leavevmode\nobreak\ \lambda_{2}$. The charged Higgs masses are varied as $70$ GeV $\leq M_{H^{\pm}}\leq 600$ GeV and $0\leq\lambda_{2}\leq 4$. We fix $\mu_{2}^{2}=200^{2}$ GeV2 and $M_{H}=150$ GeV for all cases. In the scatter plots, we set $\sqrt{\hat{s}_{\gamma\gamma}}=500$ (for all left panel plots) GeV and $\sqrt{\hat{s}_{\gamma\gamma}}=1000$ GeV (for all right panel plots). For the factors $\mu_{hh}^{\textrm{IHDM}}$ (as shown in all the above scatter plots), both the couplings $hH^{\pm}H^{\mp}$ and $hhH^{\pm}H^{\mp}$ are independent of $\lambda_{2}$. As a result, the factors only depend on $M_{H^{\pm}}$. For the factors $\mu_{HH}^{\textrm{IHDM}}$ (as shown in all the below scatter plots), it is found that the quadratic-coupling $HHH^{\pm}H^{\mp}$ depend on $\lambda_{2}$. As a result, the factors depend strongly on $\lambda_{2}$ and $M_{H^{\pm}}$. These massive contributions are mainly from the charged Higgs exchanging in the box diagrams.
$\mu_{hh}^{\textrm{IHDM}}$ $\mu_{hh}^{\textrm{IHDM}}$
Refer to caption Refer to caption
                                            $M_{H^{\pm}}$ [GeV]                                             $M_{H^{\pm}}$ [GeV]
$\mu_{HH}^{\textrm{IHDM}}$ $\mu_{HH}^{\textrm{IHDM}}$
Refer to caption Refer to caption
                                            $M_{H^{\pm}}$ [GeV]                                             $M_{H^{\pm}}$ [GeV]
Figure 10: The enhancement factors are scanned over the parameter space of $M_{H^{\pm}},\lambda_{2}$. Charged Higgs masses are in $70$ GeV $\leq M_{H^{\pm}}\leq 1000$ GeV and $0\leq\lambda_{2}\leq 4$. We fix $\mu_{2}^{2}=200^{2}$ GeV2 and $M_{H}=150$ GeV for all cases. In the plots, we set $\sqrt{\hat{s}_{\gamma\gamma}}=500$ GeV (for the above-Figures) and $\sqrt{\hat{s}_{\gamma\gamma}}=1000$ GeV (for the below-Figures), correspondingly.
4.2 THDM
The phenomenological results for the production processes $\gamma\gamma\rightarrow\phi_{i}\phi_{j}$ with CP-even Higgses $\phi_{i,j}\equiv h,H$ in the THDM are analysed in the following subsection.
4.2.1 Production cross-sections
[…]
We first inspect the data in the case of $M^{2}=0$. One notices that $\hat{\sigma}_{HH}$ become largest in the regions $\sqrt{\hat{s}_{\gamma\gamma}}\leq\sim 450$ GeV and they are decreased rapidly in the regions $\sqrt{\hat{s}_{\gamma\gamma}}\geq 450$ GeV. Moreover, $\hat{\sigma}_{hh}$ in the SM and the THDM are dominant in the regions of $\sqrt{\hat{s}_{\gamma\gamma}}\geq 450$ GeV contrasted to the ones for $\gamma\gamma\rightarrow hH,\leavevmode\nobreak\ HH$ in the THDM. Among the mentioned cross-sections, the $hh$ production in the THDM is largest in this case.
[…]
$\hat{\sigma}_{\phi_{i}\phi_{j}}(M^{2}=0)$ [pb] $\hat{\sigma}_{\phi_{i}\phi_{j}}(M^{2}=200^{2})$ [pb]
Refer to caption Refer to caption
                                                     $\sqrt{\hat{s}_{\gamma\gamma}}$ [GeV]                                                      $\sqrt{\hat{s}_{\gamma\gamma}}$ [GeV]
$\hat{\sigma}_{\phi_{i}\phi_{j}}(M^{2}=500^{2})$ [pb] $\hat{\sigma}_{\phi_{i}\phi_{j}}(M^{2}=700^{2})$ [pb]
Refer to caption Refer to caption
                                                     $\sqrt{\hat{s}_{\gamma\gamma}}$ [GeV]                                                      $\sqrt{\hat{s}_{\gamma\gamma}}$ [GeV]
Figure 11: Cross-sections for $\gamma\gamma\rightarrow\phi_{i}\phi_{j}$ in the THDM, $\gamma\gamma\rightarrow hh$ in the SM are shown as functions of $\sqrt{\hat{s}_{\gamma\gamma}}$. In the plots, we vary $350$ GeV $\leq\sqrt{\hat{s}_{\gamma\gamma}}\leq 1500$ GeV. We select $M_{H^{\pm}}=300$ GeV, $t_{\beta}=5$ in this case. Moreover, we fix $M_{H}=150$ GeV, $M^{2}=0,200^{2},500^{2},700^{2}$ GeV2 and take $c_{\beta-\alpha}=+0.2$ and $s_{\beta-\alpha}=+\sqrt{1-c_{\beta-\alpha}^{2}}$, accordingly.
Fermionphobic limit
The fermionphobic limit is studied in which the mixing angle is taken as $\alpha=\pm\pi/2$. For a typical example, we select $\alpha=+\pi/2$ in the following plots. Since, we have already checked that top quark propagating in the loop is dominant contributions versus other fermions. It is enough to take into account top quark in the loop for the present calculations. It means that the cross-sections are only contributed from $W$ boson and scalar particles in the loop when we consider the fermionphobic limit. Subsequently, we can examine the comparative sizes among these contributions. In Fig. 12 (for $hh$ production), Fig. 13 (for $hH$ production), Fig. 14 (for $HH$ production), the corresponding cross-sections for $\gamma\gamma\rightarrow\phi_{i}\phi_{j}$ in the THDM together with $\hat{\sigma}_{hh}$ in the SM as functions of $\sqrt{\hat{s}_{\gamma\gamma}}$, are analysed in the fermionphobic limit. The CoM energies are varied $350$ GeV $\leq\sqrt{\hat{s}_{\gamma\gamma}}\leq 1500$ GeV. We select $M_{H^{\pm}}=300$ GeV and apply $t_{\beta}=3$ (blue line), $t_{\beta}=5$ (green line), $t_{\beta}=7$ (red line), respectively. Moreover, we fix $M_{H}=150$ GeV for all cases. In further, the results are presented at $M^{2}=0$ GeV2 (for all the left Figures) and at $M^{2}=+200^{2}$ GeV2 (for all the right Figures).
The $hh$ productions in the SM and the THDM at $M^{2}=0$ GeV2 are first analysed. We notice that the cross-sections depend slightly on $t_{\beta}$ in the regions below the peak $\sqrt{\hat{s}_{\gamma\gamma}}\sim 2M_{H^{\pm}}=600$ GeV. In the regions above the peak, it is found that the cross-sections are more sensitive to $t_{\beta}$. For the case of $M^{2}=200^{2}$ GeV2, the cross-sections are proportional to $t_{\beta}$. Around the peak regions, the cross-sections are enhanced by charged Higgs loop.
$\hat{\sigma}_{hh}(M^{2}=0)$ [pb] $\hat{\sigma}_{hh}(M^{2}=200^{2})$ [pb]
Refer to caption Refer to caption
                                                     $\sqrt{\hat{s}_{\gamma\gamma}}$ [GeV]                                                      $\sqrt{\hat{s}_{\gamma\gamma}}$ [GeV]
Figure 12: Cross-sections for $\gamma\gamma\rightarrow hh$ in the SM and the THDM as functions $\sqrt{\hat{s}_{\gamma\gamma}}$ are shown in the fermionphobic limit. We show the data for $M^{2}=0$ GeV2 on the left panel and for $M^{2}=+200^{2}$ GeV2 on the right panel.
As mentioned in above, the cross-sections for $hH$ productions in the THDM are suppressed due to the softly breaking of the $Z_{2}$-symmetry. It is explainable for cross-sections for $hH$ productions are much smaller than the corresponding ones for $hh$ productions in the SM. However, at the peak of $\sqrt{\hat{s}_{\gamma\gamma}}=2M_{H^{\pm}}=600$ GeV, the cross-sections are enhanced and can reach to order of $\hat{\sigma}_{hh}$ in the SM. At $M^{2}=0$, $\hat{\sigma}_{\gamma\gamma\rightarrow hH}$ are more sensitive to $t_{\beta}$ in all range of CoM. Another case of $M^{2}=+200^{2}$ GeV2, the cross-sections are also more sensitive to $t_{\beta}$ in the regions below the peak $\sim 600$ GeV. But they are nearly proportional to $t_{\beta}$ beyond the peak.
$\hat{\sigma}_{hH}(M^{2}=0)$ [pb] $\hat{\sigma}_{hH}(M^{2}=200^{2})$ [pb]
Refer to caption Refer to caption
                                                     $\sqrt{\hat{s}_{\gamma\gamma}}$ [GeV]                                                      $\sqrt{\hat{s}_{\gamma\gamma}}$ [GeV]
Figure 13: Cross-sections for $\gamma\gamma\rightarrow hH$ in the SM and the THDM as functions of $\sqrt{\hat{s}_{\gamma\gamma}}$ in the fermionphobic limit. In the plots, we vary $350$ GeV $\leq\sqrt{\hat{s}_{\gamma\gamma}}\leq 1500$ GeV. We present the data for $M^{2}=0$ GeV2 on the left panel and for $M^{2}=+200^{2}$ GeV2 on the right panel.
For all cases of $M^{2}$ in $HH$ productions, it is interested in finding that the cross-sections are larger than $\hat{\sigma}_{hh}$ in the SM and they are proportional to $1/t_{\beta}$ for all range of CoM. We also observe that cross-sections are enhanced around the peak $\sim 2M_{H^{\pm}}=600$ GeV. The dominant contributions are from the singly charged Higgs exchanging in the loop.
$\hat{\sigma}_{HH}(M^{2}=0)$ [pb] $\hat{\sigma}_{HH}(M^{2}=200^{2})$ [pb]
Refer to caption Refer to caption
                                                     $\sqrt{\hat{s}_{\gamma\gamma}}$ [GeV]                                                      $\sqrt{\hat{s}_{\gamma\gamma}}$ [GeV]
Figure 14: Cross-sections for $\gamma\gamma\rightarrow HH$ in the SM and THDM as functions of $\sqrt{\hat{s}_{\gamma\gamma}}$ in the fermionphobic limit. In the plots, we vary $350$ GeV $\leq\sqrt{\hat{s}_{\gamma\gamma}}\leq 1500$ GeV. We show the data for $M^{2}=0$ GeV2 on the left panel and for $M^{2}=+200^{2}$ GeV2 on the right panel.
Decoupling limit
[…]
$\hat{\sigma}_{hh}(M^{2}=0)$ [pb] $\hat{\sigma}_{hh}(M^{2}=200^{2})$ [pb]
Refer to caption Refer to caption
                                                     $\sqrt{\hat{s}_{\gamma\gamma}}$ [GeV]                                                      $\sqrt{\hat{s}_{\gamma\gamma}}$ [GeV]
Figure 15: Cross-sections for $\gamma\gamma\rightarrow hh$ in the SM and THDM as functions of $\sqrt{\hat{s}_{\gamma\gamma}}$ are shown in the decoupling limit.
[…]
$\hat{\sigma}_{hH}(M^{2}=0)$ [pb] $\hat{\sigma}_{hH}(M^{2}=200^{2})$ [pb]
Refer to caption Refer to caption
                                                     $\sqrt{\hat{s}_{\gamma\gamma}}$ [GeV]                                                      $\sqrt{\hat{s}_{\gamma\gamma}}$ [GeV]
Figure 16: Cross-sections for $\gamma\gamma\rightarrow hH$ in the THDM as functions of $\sqrt{\hat{s}_{\gamma\gamma}}$ are shown in the decoupling limit.
$\hat{\sigma}_{HH}(M^{2}=0)$ [pb] $\hat{\sigma}_{HH}(M^{2}=200^{2})$ [pb]
Refer to caption Refer to caption
                                                     $\sqrt{\hat{s}_{\gamma\gamma}}$ [GeV]                                                      $\sqrt{\hat{s}_{\gamma\gamma}}$ [GeV]
Figure 17: Cross-sections for $\gamma\gamma\rightarrow HH$ in the THDM as functions of $\sqrt{\hat{s}_{\gamma\gamma}}$ are shown in the decoupling limit.
4.2.2 Enhancement factors
We pay attention to investigate the enhancement factors defined in Eq. 42 for $\gamma\gamma\rightarrow\phi_{i}\phi_{j}$ in the THDM. The factors scanned over parameter space of $M_{H^{\pm}},\leavevmode\nobreak\ t_{\beta}$ are first studied in this subsection. Two scenarios for $c_{\beta-\alpha}>0$ and $c_{\beta-\alpha}<0$ are examined in detail. In Figs. 18,  19, we fix $M^{2}=M_{H}^{2}=200^{2}$ GeV2. Moreover, we vary $100$ GeV $\leq M_{H^{\pm}}\leq 1000$ GeV and set $2\leq t_{\beta}\leq 10$ in the following plots. The factors $\mu_{\phi_{i}\phi_{j}}^{\textrm{THDM}}$ are generated at $\sqrt{\hat{s}_{\gamma\gamma}}=500$ GeV (for all above scatter-plots) and examined at $\sqrt{\hat{s}_{\gamma\gamma}}=1000$ GeV (for all below scatter-plots). In the left panel, we show for the enhancement factors for $hh$ productions. In the middle (right) panel, the enhancement factors for $hH,\leavevmode\nobreak\ (HH)$ productions are presented, respectively.
In Fig. 18, the first scenario for $c_{\beta-\alpha}>0$ is explored. In this scenario, we take $c_{\beta-\alpha}=+0.2$ for an example and $s_{\beta-\alpha}=+\sqrt{1-c_{\beta-\alpha}^{2}}$, correspondingly. At $\sqrt{\hat{s}_{\gamma\gamma}}=500$ GeV, $\mu_{hh}^{\textrm{THDM}}$ change from $1$ to $1.5$ for all range of $M_{H^{\pm}}$. The values of $\mu_{hh}^{\textrm{THDM}}$ are enhanced around the peak at $M_{H^{\pm}}=\sqrt{\hat{s}_{\gamma\gamma}}/2=250$ GeV. Predominantly, the factors are proportional to $t_{\beta}^{-1}$ in this case. Interestingly, we observe that the factors $\mu_{hH}^{\textrm{THDM}}$ change from 00 to $0.5$ for all range of $M_{H^{\pm}}$. The suppressed values of $\mu_{hH}^{\textrm{THDM}}$ are expected as explained in previous paragraphs due to the $Z_{2}$-symmetry. The $\mu_{hH}^{\textrm{THDM}}$ are the same behavior as $\mu_{hh}^{\textrm{THDM}}$ which they are inversely proportional to $t_{\beta}$. In the other hand, the enhancement factors for $HH$ productions in the THDM are strongly dependent of the charged Higgs mass but change slightly with $t_{\beta}$. In all range of $M_{H^{\pm}}$, the factors $\mu_{HH}^{\textrm{THDM}}$ are from $0.3$ to $1.2$.
At $\sqrt{\hat{s}_{\gamma\gamma}}=1000$ GeV, the factors $\mu_{hh}^{\textrm{THDM}}$ become biggest at the peak at $M_{H}^{\pm}=\sqrt{\hat{s}_{\gamma\gamma}}/2=500$ GeV. Around the peak, $\mu_{hh}^{\textrm{THDM}}$ change from $1.2$ to $1.8$. Beyond the peak, the factors are changed from $1.0$ to $1.2$ in all range of $M_{H^{\pm}}$. It is stress that $\mu_{hh}^{\textrm{THDM}}$ slightly change with $t_{\beta}$. For $HH$ productions, $\mu_{HH}^{\textrm{THDM}}$ are varied from $0.4$ to $2.5$ around the peak (at $M_{H^{\pm}}=500$ GeV) regions. It is realized that $\mu_{HH}^{\textrm{THDM}}$ slightly change with $t_{\beta}$. Otherwise, $\mu_{hH}^{\textrm{THDM}}$ are much smaller than $1$ and are inversely proportional to $t_{\beta}$.
$\mu_{hh}^{\textrm{THDM}}$ $\mu_{hH}^{\textrm{THDM}}$ $\mu_{HH}^{\textrm{THDM}}$
Refer to caption Refer to caption Refer to caption
                           $M_{H^{\pm}}$ [GeV]                            $M_{H^{\pm}}$ [GeV]                            $M_{H^{\pm}}$ [GeV]
$\mu_{hh}^{\textrm{THDM}}$ $\mu_{hH}^{\textrm{THDM}}$ $\mu_{HH}^{\textrm{THDM}}$
Refer to caption Refer to caption Refer to caption
                           $M_{H^{\pm}}$ [GeV]                            $M_{H^{\pm}}$ [GeV]                            $M_{H^{\pm}}$ [GeV]
Figure 18: The enhancement factors are presented in the parameter space of $M_{H^{\pm}},\leavevmode\nobreak\ t_{\beta}$. In the plots, we fix $M^{2}=M_{H}^{2}=200^{2}$ GeV2 and we take $c_{\beta-\alpha}=+0.2$ and $s_{\beta-\alpha}=+\sqrt{1-c_{\beta-\alpha}^{2}}$, accordingly. In these plots we set $\sqrt{\hat{s}_{\gamma\gamma}}=500$ (for the above-plots) GeV and $\sqrt{\hat{s}_{\gamma\gamma}}=1000$ GeV (for the below plots), respectively.
Another scenario for $c_{\beta-\alpha}<0$ is considered for examining how are the factors effect by setting different sign of $c_{\beta-\alpha}$ in this work. In Fig. 19, we take $c_{\beta-\alpha}=-0.2$ for an example and $s_{\beta-\alpha}=+\sqrt{1-c_{\beta-\alpha}^{2}}$, accordingly. At $\sqrt{\hat{s}_{\gamma\gamma}}=500$ GeV, it is excited in observing that $\mu_{hh}^{\textrm{THDM}}$ are different behavior in comparison with previous scenario. At this CoM energy, the factors $\mu_{hh}^{\textrm{THDM}}$ can reach to $1.5$ in the low region of $M_{H^{\pm}}<200$ GeV. They then are decreased to around $0.9$ when $M_{H^{\pm}}>200$ GeV. There isn’t peak of the factors observed in this scenario. Because the contributions of singly charged Higgs exchanging in the one-loop triangle diagrams may cancel out with the ones one-loop box diagrams in this scenario. Surprisingly, we find that the factors $\mu_{hh}^{\textrm{THDM}}$ are proportional to $t_{\beta}$ in this scenario. For the $hH$ productions, the factors are suppressed and they are in the range of [0.025,0.3][$\sim$ 0.025$\sim$ 0.3][ ∼ 0.025 , ∼ 0.3 ]. They are sensitive with $t_{\beta}^{-1}$ in all range of charged Higgs mass. In $HH$ productions, it is found that the factors develop to the peak around $M_{H^{\pm}}=500$ GeV. They reach to factor $2.5$ around the peak and they are in the ranges of [0.5,1.7][$\sim$ 0.5$\sim$ 1.7][ ∼ 0.5 , ∼ 1.7 ] beyond the peak regions. In all range of $M_{H^{\pm}}$, the factors $\mu_{HH}^{\textrm{THDM}}$ are proportional to $t_{\beta}^{-1}$ in this scenario.
The survey for all the enhancement factors at $\sqrt{\hat{s}_{\gamma\gamma}}=1000$ GeV are concerned in the next paragraphs. The factors $\mu_{hh}^{\textrm{THDM}}$ are large in the regions ($M_{H^{\pm}}\leq 200$ GeV) and they can reach to $1.5$. They then are decreased rapidly to the regions around $M_{H^{\pm}}\sim 300$ GeV and develop to the peak at $M_{H^{\pm}}=\sqrt{\hat{s}_{\gamma\gamma}}/2=500$ GeV. Around the peak, the enhancement factor is about $1.2$. In other ranges of charged Higgs mass, $\mu_{hh}^{\textrm{THDM}}\sim 0.9$. One also finds that $\mu_{hh}^{\textrm{THDM}}$ is inversely proportional to $t_{\beta}$ in this scenario. In $hH$ productions, the factors are increased to the peak $M_{H^{\pm}}=\sqrt{\hat{s}_{\gamma\gamma}}/2=500$ GeV and they are about $0.3$ around the peak. In all regions of $M_{H^{\pm}}$, the mentioned factors are in the ranges of [0.025,0.3][$\sim$ 0.025$\sim$ 0.3][ ∼ 0.025 , ∼ 0.3 ] and they are inversely proportional to $t_{\beta}$. In the last case, it is found that the factors $\mu_{HH}^{\textrm{THDM}}$ are same behavior as previous scenario. They are in the ranges of [0.5,4][$\sim$ 0.5$\sim$ 4][ ∼ 0.5 , ∼ 4 ] in all regions of $M_{H^{\pm}}$. However, the factors $\mu_{HH}^{\textrm{THDM}}$ depend slightly on $t_{\beta}$ in this scenario.
$\mu_{hh}^{\textrm{THDM}}$ $\mu_{hH}^{\textrm{THDM}}$ $\mu_{HH}^{\textrm{THDM}}$
Refer to caption Refer to caption Refer to caption
                           $M_{H^{\pm}}$ [GeV]                            $M_{H^{\pm}}$ [GeV]                            $M_{H^{\pm}}$ [GeV]
$\mu_{hh}^{\textrm{THDM}}$ $\mu_{hH}^{\textrm{THDM}}$ $\mu_{HH}^{\textrm{THDM}}$
Refer to caption Refer to caption Refer to caption
                           $M_{H^{\pm}}$ [GeV]                            $M_{H^{\pm}}$ [GeV]                            $M_{H^{\pm}}$ [GeV]
Figure 19: The enhancement factors are presented in the parameter space $M_{H^{\pm}},\leavevmode\nobreak\ t_{\beta}$. In the plots, we consider the scenario for $c_{\beta-\alpha}=-0.2<0$ and $s_{\beta-\alpha}=+\sqrt{1-c_{\beta-\alpha}^{2}}$, correspondingly. We also fix $M^{2}=M_{H}^{2}=200^{2}$ GeV2 and set $\sqrt{\hat{s}_{\gamma\gamma}}=500$ (for all the above-plots) GeV and $\sqrt{\hat{s}_{\gamma\gamma}}=1000$ GeV (for all the below-plots).
The enhancement factors scanned over the parameter space of $M_{H^{\pm}},\leavevmode\nobreak\ M^{2}$ in the THDM are also interested greatly in this work. Two scenarios for $c_{\beta-\alpha}>0$ and $c_{\beta-\alpha}<0$ are studied in detail in the following paragraphs. In Fig. 20 (for $c_{\beta-\alpha}>0$ scenario), Fig. 21 (for $c_{\beta-\alpha}<0$ scenario), we consider $\sqrt{\hat{s}_{\gamma\gamma}}=500$ GeV (for all the above scatter plots) and $\sqrt{\hat{s}_{\gamma\gamma}}=1000$ GeV (for all the below scatter plots). Moreover, we vary charged Higgs mass as $100$ GeV $\leq M_{H^{\pm}}\leq 1000$ GeV, the soft-breaking parameter as 00 GeV2 $\leq M^{2}\leq 200^{2}$ GeV2 and take $t_{\beta}=5$ for all cases.
In Fig. 20, the first scenario of $c_{\beta-\alpha}>0$ is examined. For this case, we take $c_{\beta-\alpha}=+0.2$ as an example and $s_{\beta-\alpha}=+\sqrt{1-c_{\beta-\alpha}^{2}}$, accordingly. For $hh$ production at $\sqrt{\hat{s}_{\gamma\gamma}}=500$ GeV, we observe the peak of $\mu_{hh}^{\textrm{THDM}}$ at $M_{H^{\pm}}=250$ GeV which is corresponding to the threshold of cross-sections $hh$ in the THDM at the peak $\sqrt{\hat{s}_{\gamma\gamma}}=2M_{H^{\pm}}$. Around the peak, $\mu_{hh}^{\textrm{THDM}}$ are varied from $1.0$ to $1.8$. Above the peak regions, the enhancement factors tend to $1.2$ and depend slightly on $M^{2}$. For $hH$ productions, $\mu_{hH}^{\textrm{THDM}}$ are more sensitive with $M^{2}$ in below the peak regions. The factors are in the ranges of $[0.07,1.5]$ in above the peak regions. Around the peak, $\mu_{hH}^{\textrm{THDM}}$ can reach to $0.3$. We also find the same behavior for $\mu_{HH}^{\textrm{THDM}}$. The factors for $HH$ productions are large in the low regions of $M_{H^{\pm}}$ and around the peak $M_{H^{\pm}}=250$ GeV. They are in the ranges of $[0.7,2.4]$ in the above the peak regions. Generally, we observe that $\mu_{\phi_{i}\phi_{j}}^{\textrm{THDM}}$ are proportional to $1/M$ at this CoM energy.
At $\sqrt{\hat{s}_{\gamma\gamma}}=1000$ GeV, we also find that $\mu_{hh}^{\textrm{THDM}}$ develop to the peak at $M_{H^{\pm}}=500$ GeV where the factors can reach to $2.2$ and are decreased rapidly beyond the peak. The factors depend slightly on $M^{2}$ and tend to $1$ beyond the peak regions. For $hH$ productions, $\mu_{hH}^{\textrm{THDM}}$ are sensitive with $M^{2}$ in the peak regions. They tend to $0.05$ and they are slightly dependent of $M^{2}$ in the above the peak regions. For $HH$ productions, the factors become large in the below the peak regions and they are inversely proportional to $M^{-1}$. Around the peak, the factors are enhanced by large values of $M^{2}$. Above the peak regions, $\mu_{HH}^{\textrm{THDM}}$ are varied around $1.0$.
$\mu_{hh}^{\textrm{THDM}}$ $\mu_{hH}^{\textrm{THDM}}$ $\mu_{HH}^{\textrm{THDM}}$
Refer to caption Refer to caption Refer to caption
                           $M_{H^{\pm}}$ [GeV]                            $M_{H^{\pm}}$ [GeV]                            $M_{H^{\pm}}$ [GeV]
$\mu_{hh}^{\textrm{THDM}}$ $\mu_{hH}^{\textrm{THDM}}$ $\mu_{HH}^{\textrm{THDM}}$
Refer to caption Refer to caption Refer to caption
                           $M_{H^{\pm}}$ [GeV]                            $M_{H^{\pm}}$ [GeV]                            $M_{H^{\pm}}$ [GeV]
Figure 20: The enhancement factors are presented in the parameter space of $M_{H^{\pm}},\leavevmode\nobreak\ M^{2}$. In the plots, we take $t_{\beta}=5$ and consider the first scenario of $c_{\beta-\alpha}=+0.2$ and $s_{\beta-\alpha}=+\sqrt{1-c_{\beta-\alpha}^{2}}$, accordingly. In these plots, we set $\sqrt{\hat{s}_{\gamma\gamma}}=500$ GeV (for the above Figures) and $\sqrt{\hat{s}_{\gamma\gamma}}=1000$ GeV (for the below Figures).
[…]
We next comment on physical results at $\sqrt{\hat{s}_{\gamma\gamma}}=1000$ GeV. The factors $\mu_{hh}^{\textrm{THDM}}$ are enhanced by the small values of $M$ in low regions of charged Higgs mass and they can reach $1.2$. They tend to $2.5$ around the peak. The factors are then varied around $\sim 0.9$. In general, the factors depend on $M^{-1}$. In the productions $hH$, $\mu_{hH}^{\textrm{THDM}}$ are more sensitive to $M^{-1}$ around the peak regions. They then tend to $0.2$ in the high mass regions of singly charged Higgs. For $HH$ productions, the factors $\mu_{HH}^{\textrm{THDM}}$ strongly depend on $M^{-1}$. At the peak, the factors are enhanced by the large value of $M$. Above the peak regions, $\mu_{HH}^{\textrm{THDM}}$ tend to $\sim 1$.
$\mu_{hh}^{\textrm{THDM}}$ $\mu_{hH}^{\textrm{THDM}}$ $\mu_{HH}^{\textrm{THDM}}$
Refer to caption Refer to caption Refer to caption
                           $M_{H^{\pm}}$ [GeV]                            $M_{H^{\pm}}$ [GeV]                            $M_{H^{\pm}}$ [GeV]
$\mu_{hh}^{\textrm{THDM}}$ $\mu_{hH}^{\textrm{THDM}}$ $\mu_{HH}^{\textrm{THDM}}$
Refer to caption Refer to caption Refer to caption
                           $M_{H^{\pm}}$ [GeV]                            $M_{H^{\pm}}$ [GeV]                            $M_{H^{\pm}}$ [GeV]
Figure 21: The enhancement factors are presented in the parameter space of $M_{H^{\pm}},\leavevmode\nobreak\ M^{2}$. In the plots, we take $t_{\beta}=5$ and $c_{\beta-\alpha}=-0.2$ and $s_{\beta-\alpha}=+\sqrt{1-c_{\beta-\alpha}^{2}}$, accordingly. In these plots, we set $\sqrt{\hat{s}_{\gamma\gamma}}=500$ GeV (the above Figures) and $\sqrt{\hat{s}_{\gamma\gamma}}=1000$ GeV (the below Figures).
5 Conclusions
In this paper, we have presented the results for one-loop induced processes $\gamma\gamma\rightarrow\phi_{i}\phi_{j}$ with CP-even Higgses $\phi_{i,j}\equiv h,\leavevmode\nobreak\ H$ at high energy photon-photon collision in the IHDM and the THDM. In the phenomenological results, we have shown the cross-sections at several center-of-mass energies. The results show that cross-sections for the computed processes in the models under investigations are enhanced at around the threshold of charged Higgs pair ($\sim 2M_{H^{\pm}}$). Furthermore, the enhancement factors for the processes are examined in parameter space of the models under consideration. In the IHDM, the factors are studied in the parameter space of $(M_{H^{\pm}},\leavevmode\nobreak\ \mu^{2}_{2})$ and $(M_{H^{\pm}},\leavevmode\nobreak\ \lambda_{2})$. In the the THDM, the factors are analysed in the planes of $(M_{H^{\pm}},\leavevmode\nobreak\ t_{\beta})$ and $(M_{H^{\pm}},\leavevmode\nobreak\ M^{2})$. Two scenarios of $c_{\beta-\alpha}>0$ and $c_{\beta-\alpha}<0$ have studied in further detail. The factors give a different behavior from considering these scenarios. As a result, discriminations for the above-mentioned scenarios can be performed at future colliders.
Acknowledgment:  This research is funded by Vietnam National Foundation for Science and Technology Development (NAFOSTED) under the grant number $103.01$-$2023.16$.
Appendix A: Effective Lagrangian in the IHDM
The kinematic terms of Lagrangian in the IHDM can be expanded as follows:
$\mathcal{L}_{K}^{\text{IHDM}}$ $\supset$ $\frac{2M_{W}^{2}}{v}W^{\pm}_{\mu}W^{\mp,\mu}h+\frac{M_{Z}^{2}}{v}hZ_{\mu}Z^{\mu}+i\frac{M_{Z}c_{2W}}{v}Z^{\mu}(H^{\mp}\partial_{\mu}H^{\pm}-H^{\pm}\partial_{\mu}H^{\mp})$ (43)
$+i\frac{M_{Z}s_{2W}}{v}A^{\mu}(H^{\mp}\partial_{\mu}H^{\pm}-H^{\pm}\partial_{\mu}H^{\mp})+i\frac{2M_{Z}^{2}c_{W}^{2}s_{W}}{v}A^{\mu}W_{\mu}^{\pm}G^{\pm}$
$+i\frac{M_{W}}{v}(HW^{\mp,\mu}\partial_{\mu}H^{\pm}-H^{\pm}W^{\mp,\mu}\partial_{\mu}H-HW^{\pm}_{\mu}\partial^{\mu}H^{\mp}+H^{\mp}W^{\pm}_{\mu}\partial^{\mu}H)$
$+\frac{M_{Z}^{2}s_{2W}^{2}}{v^{2}}A_{\mu}A^{\mu}H^{\pm}H^{\mp}+\frac{2M_{Z}^{2}c_{W}^{2}s_{W}}{v^{2}}HH^{\pm}W_{\mu}^{\mp}A^{\mu}+\frac{2M_{W}^{2}}{v^{2}}W^{\pm}_{\mu}W^{\mp,\mu}hh$
$+\frac{2M_{W}^{2}}{v^{2}}W^{\pm}_{\mu}W^{\mp,\mu}HH+\frac{M_{Z}^{2}s_{2W}^{2}}{v^{2}}A_{\mu}A^{\mu}G^{\pm}G^{\mp}$
$+i\frac{M_{Z}s_{2W}}{v}A^{\mu}(G^{\mp}\partial_{\mu}G^{\pm}-G^{\pm}\partial_{\mu}G^{\mp})+\cdots.$
We also expand the scalar Higgs potential of the IHDM and collect the terms involving to Higgs self-coupling as follows:
$-\mathcal{V}_{\text{IHDM}}(\phi_{1},\phi_{2})$ $\supset$ $-\frac{3M_{h}^{2}}{v}hhh+\frac{2(\mu_{2}^{2}-M_{H}^{2})}{v}hHH+\frac{2(\mu_{2}^{2}-M_{H^{\pm}}^{2})}{v}hH^{\pm}H^{\mp}$ (44)
$+\frac{M_{H^{\pm}}^{2}-M_{H}^{2}}{v}HH^{\pm}{G^{\mp}}+\frac{2(\mu_{2}^{2}-M_{H^{\pm}}^{2})}{v^{2}}hhH^{\pm}H^{\mp}-2\lambda_{2}HHH^{\pm}H^{\mp}$
$-\frac{M_{h}^{2}}{v^{2}}hhG^{\pm}G^{\mp}+\frac{2(\mu_{2}^{2}-M_{H^{\pm}}^{2})}{v^{2}}HHG^{\pm}G^{\mp}+\cdots.$
Appendix B: Effective Lagrangian in the THDM
We expand the kinematic terms of Lagrangian in the THDM as follows:
$\mathcal{L}_{K}^{\text{THDM}}$ $\supset$ $\frac{2M_{W}^{2}}{v}s_{\beta-\alpha}hW_{\mu}^{\pm}W^{\mp,\mu}+\frac{2M_{W}^{2}}{v}c_{\alpha-\beta}HW_{\mu}^{\pm}W^{\mp,\mu}+\frac{M_{Z}^{2}}{v}s_{\beta-\alpha}hZ_{\mu}Z^{\mu}$ (45)
$+\frac{M_{Z}^{2}}{v}c_{\alpha-\beta}HZ_{\mu}Z^{\mu}+i\frac{M_{Z}c_{2W}}{v}Z^{\mu}(H^{\mp}\partial_{\mu}{H^{\pm}}-H^{\pm}\partial_{\mu}H^{\mp})$
$+i\frac{M_{Z}s_{2W}}{v}A^{\mu}(H^{\mp}\partial_{\mu}{H^{\pm}}-H^{\pm}\partial_{\mu}H^{\mp})+\frac{4M_{W}^{2}s_{W}^{2}}{v^{2}}H^{\pm}H^{\mp}A_{\mu}{A^{\mu}}$
$-i\frac{M_{W}s_{\beta-\alpha}}{v}(HW^{\mp,\mu}\partial_{\mu}H^{\pm}-HW^{\pm,\mu}\partial_{\mu}H^{\mp}+H^{\mp}W^{\pm,\mu}\partial_{\mu}H-H^{\pm}W^{\mp,\mu}\partial_{\mu}H)$
$-i\frac{M_{W}c_{\beta-\alpha}}{v}(-hW^{\mp,\mu}\partial_{\mu}H^{\pm}+hW^{\pm,\mu}\partial_{\mu}H^{\mp}-H^{\mp}W^{\pm,\mu}\partial_{\mu}h+H^{\pm}W^{\mp,\mu}\partial_{\mu}h)$
$+\frac{2M_{W}^{2}s_{W}c_{\beta-\alpha}}{v^{2}}hH^{\mp}W_{\mu}^{\pm}A^{\mu}-\frac{2M_{W}^{2}s_{W}s_{\beta-\alpha}}{v^{2}}HH^{\mp}W^{\pm}_{\mu}A^{\mu}$
$+\dfrac{2M_{W}^{2}}{v^{2}}W^{\pm}_{\mu}W^{\mp,\mu}HH+\frac{2M_{W}^{2}}{v^{2}}W^{\pm}_{\mu}W^{\mp,\mu}hh+\cdots$
Expanding the scalar potential in the THDM, we then collect the terms involving to the Higgs self-couplings as
$-\mathcal{V}_{\text{THDM}}(\phi_{1},\phi_{2})$ $\supset$ $-\lambda_{hHH}hHH-\lambda_{Hhh}Hhh-\lambda_{hH^{\pm}H^{\mp}}hH^{\pm}H^{\mp}$ (46)
$-\lambda_{hH^{\pm}H^{\mp}}HH^{\pm}H^{\mp}-\lambda_{HhH^{\pm}H^{\mp}}HhH^{\pm}H^{\mp}+\cdots.$
All coefficients of the mentioned couplings are shown explicitly in terms of physical parameters as follows:
$-\lambda_{hHH}$ $=$ $\frac{3\lambda_{1}v}{2}s_{\alpha}c_{\alpha}^{2}c_{\beta}-\frac{3\lambda_{2}v}{2}s_{\beta}s_{\alpha}^{2}c_{\alpha}-\frac{\lambda_{345}}{2}v[c_{\beta}(2s_{\alpha}c_{\alpha}^{2}-s_{\alpha}^{3})+s_{\beta}(c_{\alpha}^{3}-2s_{\alpha}^{2}c_{\alpha})]$ (47)
$=$ $\frac{s_{\alpha-\beta}[s_{2\alpha}(3M^{2}-M_{h}^{2}-2M_{H}^{2})+M^{2}s_{2\beta}]}{v\;s_{2\beta}},$ (48)
$-\lambda_{Hhh}$ $=$ $-\frac{3\lambda_{1}v}{2}c_{\beta}c_{\alpha}s_{\alpha}^{2}-\frac{3\lambda_{2}v}{2}s_{\beta}c_{\alpha}^{2}s_{\alpha}-\frac{\lambda_{345}}{2}v[s_{\beta}(s_{\alpha}^{3}-2c_{\alpha}^{2}s_{\alpha})-c_{\beta}(2c_{\alpha}s_{\alpha}^{2}-c_{\alpha}^{3})]$ (49)
$=$ $\frac{c_{\alpha-\beta}[s_{2\alpha}(3M^{2}-M_{H}^{2}-2m_{h}^{2})-M^{2}s_{2\beta}]}{v\;s_{2\beta}},$ (50)
$-\lambda_{HH^{\pm}H^{\mp}}$ $=$ $-\lambda_{1}vc_{\beta}c_{\alpha}s_{\beta}^{2}-\lambda_{2}vs_{\beta}s_{\alpha}c_{\beta}^{2}-\lambda_{3}v(s_{\beta}s_{\alpha}s_{\beta}^{2}+c_{\beta}c_{\alpha}c_{\beta}^{2})+\frac{\lambda_{45}}{2}vs_{(2\beta)}\;s_{\beta+\alpha}$ (51)
$=$ $\frac{s_{\alpha+\beta}(4M^{2}-3M_{H}^{2}-2M_{H^{\pm}}^{2})+(2M_{H^{\pm}}^{2}-M_{H}^{2})s_{\alpha-3\beta}}{2vs_{(2\beta)}},$ (52)
$-\lambda_{hH^{\pm}H^{\mp}}$ $=$ $\lambda_{1}vc_{\beta}s_{\alpha}s_{\beta}^{2}-\lambda_{2}vs_{\beta}c_{\alpha}c_{\beta}^{2}-\lambda_{3}v(s_{\beta}c_{\alpha}s_{\beta}^{2}-c_{\beta}s_{\alpha}c_{\beta}^{2})+\frac{\lambda_{45}}{2}v\;s_{(2\beta)}\;c_{(\beta+\alpha)}$ (53)
$=$ $\frac{c_{\alpha+\beta}(4M^{2}-3M_{h}^{2}-2M_{H^{\pm}}^{2})+(2M_{H^{\pm}}^{2}-M_{h}^{2})c_{(\alpha-3\beta)}}{2vs_{2\beta}},$ (54)
and
$-\lambda_{HhH^{\pm}H^{\mp}}$ $=$ $\lambda_{1}s_{\beta}^{2}s_{\alpha}c_{\alpha}-\lambda_{2}c_{\beta}^{2}s_{\alpha}c_{\alpha}+\lambda_{3}s_{\alpha}c_{\alpha}c_{2\beta}+(\lambda_{4}+\lambda_{5})s_{\beta}c_{\beta}c_{2\alpha}$ (56)
$=$ $\frac{s_{2\alpha}(3c_{2\alpha}+c_{2(\alpha-2\beta)}-4c_{2\beta})}{4v^{2}s_{2\beta}^{2}}M_{H}^{2}-\frac{s_{2\alpha}(3c_{2\alpha}+c_{2(\alpha-2\beta)}+4c_{2\beta})}{4v^{2}s_{2\beta}^{2}}M_{h}^{2}$
$+\frac{s_{2(\alpha-\beta)}}{v^{2}}M_{H^{\pm}}^{2}+\frac{(s_{2(\alpha-3\beta)}+2s_{2(\alpha-\beta)}+5s_{2(\alpha+\beta)})}{4v^{2}s_{2\beta}^{2}}M^{2},$
$-\lambda_{HhG^{\pm}G^{\mp}}$ $=$ $\lambda_{1}c_{\beta}^{2}s_{\alpha}c_{\alpha}-\lambda_{2}s_{\beta}^{2}s_{\alpha}c_{\alpha}-\lambda_{3}s_{\alpha}c_{\alpha}c_{2\beta}-(\lambda_{4}+\lambda_{5})s_{\beta}c_{\beta}c_{2\alpha}$
$=$ $\frac{1}{2v^{2}s_{2\beta}}s_{2(\alpha-\beta)}[(M_{h}^{2}-M_{H}^{2})s_{2\alpha}+2(M^{2}-M_{H^{\pm}}^{2})s_{2\beta}].$ (57)
Furthermore, we have the following couplings:
$-\lambda_{hhh}$ $=$ $3v\Big{[}\lambda_{1}s_{\alpha}^{3}c_{\beta}-\lambda_{2}c_{\alpha}^{3}s_{\beta}+(\lambda_{3}+\lambda_{4}+\lambda_{5})(s_{\alpha}c_{\beta}c_{\alpha}^{2}-c_{\alpha}s_{\beta}s_{\alpha}^{2})\Big{]}$ (58)
$=$ $\dfrac{3e}{4M_{W}s_{W}s_{2\beta}}\Big{[}M^{2}c_{\alpha-3\beta}+(M^{2}-M_{h}^{2})c_{3\alpha-\beta}+(2M^{2}-3M_{h}^{2})c_{\alpha+\beta}\Big{]},$
$-\lambda_{HHH}$ $=$ $-3v\Big{[}\lambda_{1}c_{\alpha}^{3}c_{\beta}+\lambda_{2}s_{\alpha}^{3}s_{\beta}+(\lambda_{3}+\lambda_{4}+\lambda_{5})(c_{\alpha}c_{\beta}s_{\alpha}^{2}+s_{\alpha}s_{\beta}c_{\alpha}^{2})\Big{]}$ (59)
$=$ $\dfrac{3e}{4M_{W}s_{W}s_{2\beta}}\Big{[}M^{2}s_{\alpha-3\beta}+(M_{H}^{2}-M^{2})s_{3\alpha-\beta}+(2M^{2}-3M_{H}^{2})s_{\alpha+\beta}\Big{]},$
$-\lambda_{HHH^{\pm}{H^{\mp}}}$ $=$ $-\lambda_{1}c_{\alpha}^{2}s_{\beta}^{2}-\lambda_{2}s_{\alpha}^{2}c_{\beta}^{2}-\lambda_{3}(c_{\alpha}^{2}c_{\beta}^{2}+s_{\alpha}^{2}s_{\beta}^{2})+(\lambda_{4}+\lambda_{5})c_{\alpha}s_{\alpha}s_{2\beta}$ (60)
$=$ $-\dfrac{2c_{\alpha-\beta}^{2}}{v^{2}}M_{H^{\pm}}^{2}-\frac{s_{2\alpha}[3s_{2\alpha}+s_{2(\alpha-2\beta)}-2s_{2\beta}]}{4v^{2}s_{2\beta}^{2}}M_{h}^{2}$
$-\dfrac{[c_{\alpha}^{4}+c_{\alpha}^{3}s_{\alpha}\cot^{3}\beta+c_{\alpha}s_{\alpha}^{3}\cot\beta+s_{\alpha}^{4}\cot^{4}\beta]\tan^{2}\beta}{v^{2}}M_{H}^{2}$
$+\dfrac{s_{\beta}[4c_{\alpha}c_{\beta}s_{\alpha}+(1+\cot^{4}\beta)s_{\alpha}^{2}s_{\beta}+c_{\alpha}^{2}(1+\cot^{4}\beta)s_{\beta}\tan^{2}\beta]}{v^{2}}M^{2},$
$-\lambda_{hhH^{\pm}H^{\mp}}$ $=$ $-\lambda_{1}s_{\alpha}^{2}s_{\beta}^{2}-\lambda_{2}c_{\alpha}^{2}c_{\beta}^{2}-\lambda_{3}(s_{\alpha}^{2}c_{\beta}^{2}+c_{\alpha}^{2}s_{\beta}^{2})-(\lambda_{4}+\lambda_{5})c_{\alpha}s_{\alpha}s_{2\beta}$ (61)
$=$ $-\dfrac{2s_{\alpha-\beta}^{2}}{v^{2}}M_{H^{\pm}}^{2}-\dfrac{s_{2\alpha}[3s_{2\alpha}+s_{2(\alpha-2\beta)}+2s_{2\beta}]}{4v^{2}s_{2\beta}^{2}}M_{H}^{2}$
$+\dfrac{(-c_{\alpha}^{4}\cot^{2}{\beta}+c_{\alpha}s_{\alpha}^{3}\cot\beta+c_{\alpha}^{3}s_{\alpha}\tan\beta-s_{\alpha}^{4}\tan^{2}\beta)}{v^{2}}M_{h}^{2}$
$+\dfrac{s_{\beta}c_{\beta}[-4c_{\alpha}s_{\alpha}+(1+\cot^{4}\beta)s_{\alpha}^{2}\tan^{3}{\beta}+c_{\alpha}^{2}(\cot^{3}\beta+\tan\beta)]}{v^{2}}M^{2},$
$-\lambda_{hhG^{\pm}G^{\mp}}$ $=$ $-\lambda_{1}s_{\alpha}^{2}c_{\beta}^{2}-\lambda_{2}c_{\alpha}^{2}s_{\beta}^{2}-\lambda_{3}(c_{\alpha}^{2}c_{\beta}^{2}+s_{\alpha}^{2}s_{\beta}^{2})+(\lambda_{4}+\lambda_{5})c_{\alpha}s_{\alpha}s_{2\beta}$ (62)
$=$ $\dfrac{2c_{\alpha-\beta}^{2}}{v^{2}}(M^{2}-M_{H^{\pm}}^{2})-\dfrac{c_{\alpha-\beta}^{2}s_{2\alpha}}{v^{2}s_{2\beta}}M_{H}^{2}+\dfrac{-3s_{2\beta}+2s_{2\alpha}+s_{4\alpha-2\beta}}{4v^{2}s_{2\beta}}M_{h}^{2},$
$-\lambda_{HHG^{\pm}G^{\mp}}$ $=$ $-\lambda_{1}c_{\alpha}^{2}c_{\beta}^{2}-\lambda_{2}s_{\alpha}^{2}s_{\beta}^{2}-\lambda_{3}(s_{\alpha}^{2}c_{\beta}^{2}+c_{\alpha}^{2}s_{\beta}^{2})-(\lambda_{4}+\lambda_{5})c_{\alpha}s_{\alpha}s_{2\beta}$ (63)
$=$ $\dfrac{2s_{\alpha-\beta}^{2}}{v^{2}}(M^{2}-M_{H^{\pm}}^{2})+\dfrac{s_{\alpha-\beta}^{2}s_{2\alpha}}{v^{2}s_{2\beta}}M_{h}^{2}+\dfrac{-3s_{2\beta}-2s_{2\alpha}+s_{4\alpha-2\beta}}{4v^{2}s_{2\beta}}M_{H}^{2}.$
Additionally, we also derive the couplings relating to Goldstone bosons as follows:
$\mathcal{L}_{K}^{\text{THDM}}$ $\supset$ $\frac{2M_{W}^{2}s_{W}}{v}A_{\mu}W^{\pm,\mu}G^{\mp}+\frac{4M_{W}^{2}s_{W}^{2}}{v^{2}}A_{\mu}A^{\mu}G^{\pm}{G^{\mp}}$ (64)
$+i\frac{2M_{W}s_{W}}{v}A^{\mu}(G^{\mp}\partial_{\mu}G^{\pm}-G^{\pm}\partial_{\mu}G^{\mp})+\cdots$
From scalar potential, we have
$-\mathcal{V}(\phi_{1},\phi_{2})$ $\supset$ $-\lambda_{hH^{\pm}G^{\mp}}hH^{\pm}G^{\mp}-\lambda_{HH^{\pm}G^{\mp}}HH^{\pm}G^{\mp}+\cdots$ (65)
where the coefficients of the couplings are given by
$-\lambda_{hH^{\pm}G^{\mp}}$ $=$ $\frac{ec_{\alpha-\beta}}{2M_{W}s_{W}}(M_{H^{\pm}}^{2}-M_{h}^{2}),$ (66)
$-\lambda_{HH^{\pm}G^{\mp}}$ $=$ $\frac{es_{\alpha-\beta}}{2M_{W}s_{W}}(M_{H^{\pm}}^{2}-M_{H}^{2}).$ (67)