### v3

# Processes $\gamma\gamma\rightarrow\phi_{i}\phi_{j}$ in Inert Higgs Doublet Models and Two Higgs Doublet Models

Abstract
In this paper, we present a phenomenological analysis of one-loop induced processes $\gamma\gamma\rightarrow\phi_{i}\phi_{j}$, where the CP-even Higgs bosons are denoted as $\phi_{i,j}\equiv h,\leavevmode\nobreak\ H$, in high-energy photon-photon collisions, within the frameworks of the Inert Higgs Doublet Model and the Two Higgs Doublet Model. The total cross sections are evaluated as functions of the center-of-mass energy, finding that the cross sections for the considered processes in all the models under investigation are significantly enhanced around the threshold for charged Higgs boson pair production ($\sim 2M_{H^{\pm}}$). Furthermore, the enhancement factors, defined as the ratios of cross sections of $\gamma\gamma\rightarrow\phi_{i}\phi_{j}$ in the investigated models to those for $\gamma\gamma\rightarrow hh$ in the Standard Model, are examined in the relevant regions of the model’s parameter space. In the Inert Higgs Doublet Model, the factors are studied in the parameter space of $(M_{H^{\pm}},\leavevmode\nobreak\ \mu^{2}_{2})$ and $(M_{H^{\pm}},\leavevmode\nobreak\ \lambda_{2})$. In the Two Higgs Doublet Model, the factors are examined in the planes defined by $(M_{H^{\pm}},\leavevmode\nobreak\ t_{\beta})$ as well as in the space of the charged Higgs mass $M_{H^{\pm}}$ and the soft-breaking $Z_{2}$ parameter $m_{12}^{2}$. Two scenarios characterized by $c_{\beta-\alpha}>0$ and $c_{\beta-\alpha}<0$ are studied in further detail. The factors exhibit distinct behaviors between these two scenarios. As a result, it is possible to discriminate between them at future colliders. The dependence of the cross section for the process $\gamma\gamma\rightarrow hH$ on $m_{12}^{2}$ provides a potential probe of the soft $Z_{2}$-breaking scale in the Two Higgs Doublet Model.
keywords:
Higgs phenomenology, one-loop corrections, analytic methods for quantum field theory, dimensional regularization.
1 Introduction
Measuring scalar Higgs self-couplings, including Standard Model-like Higgs trilinear and quadratic couplings, as well as the couplings between scalar Higgses in many beyond-the-Standard-Model (BSM) scenarios, plays a key role in determining the scalar potential. We can subsequently answer the question of the origin of electroweak symmetry breaking (EWSB). In this scheme, Higgs boson pair production and multi-scalar Higgs production should be measured precisely at future colliders. Recently, searches for Higgs boson pair production in final states such as two bottom quarks associated with two photons, four bottom quarks, etc., in proton-proton collisions have been performed at the Large Hadron Collider (LHC), as in [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12]. It is well known that measurements of Higgs self-couplings are rather challenging at the LHC. The results from the study in [13, 14] show that the expected accuracy in the measurement of trilinear Higgs self-couplings would be about $20\%$$30\%$ at the high luminosity of $3000$ fb-1. We know that the physics program of future lepton colliders (LC) will be complementary to that of the LHC in many aspects, as studied in [15]. Furthermore, the LC can significantly improve LHC measurements in many cases. More importantly, photon-photon collisions are an option at the LC [16, 17], where scalar Higgs pair production ($\phi_{i}\phi_{j}$) can be measured via the channels $f\bar{f}\rightarrow f\bar{f}\gamma^{*}\gamma^{*}\rightarrow f\bar{f}\phi_{i}\phi_{j}$ for $f\equiv e,\mu$. In this regard, the LC could open a window for probing new physics signals through multi-scalar Higgs production.
From a theoretical calculation perspective, one-loop corrections to double Higgs production at the LHC have been calculated in the Standard Model (SM), the Higgs Extensions of the Standard Models (HESM), as well as in other BSM frameworks by many research groups. For example, it is worth mentioning notable works in [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78], for further reviews in [79] and the related references therein. Additionally, one-loop corrections for $hh$ production in high-energy $\gamma\gamma$ collisions in the SM, HESM, and other BSM models have been computed in Refs. [80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94]. Other computations for one-loop corrections to Higgs boson pair production at future linear lepton colliders, including multi-TeV muon colliders, have been performed in Refs. [95, 96, 97, 98] and the additional references therein. Furthermore, double pseudo-scalar Higgs $A^{0}A^{0}$ production at a $\gamma\gamma$ collider in the Two Higgs Doublet Model has been reported in Ref. [99]. One-loop contributions to the process $\gamma\gamma\rightarrow\phi_{i}\phi_{j}$ with CP-even Higgses $\phi_{i,j}\equiv h,\leavevmode\nobreak\ H$ in high-energy $\gamma\gamma$ collisions, valid for the Higgs Extensions of the Standard Models (HESM), are necessary. In our previous paper, analytic formulas for the concerned processes were reported [100].
In this paper, we present the phenomenological analysis of these processes at high-energy $\gamma\gamma$ collisions in the Inert Higgs Doublet Model (IHDM) and the Two Higgs Doublet Model (THDM). In the phenomenological analysis, cross sections are shown as functions of center-of-mass energies. Furthermore, the enhancement factors, defined as the ratios of cross sections of $\gamma\gamma\rightarrow\phi_{i}\phi_{j}$ in the HESMs to the corresponding ones for $\gamma\gamma\rightarrow hh$ in the Standard Model, are examined in the parameter space of the models under consideration. The paper is structured as follows. We briefly review the HESM in the next section. The phenomenological studies for the HESM are discussed in section 3. In section 4, we present the conclusion and outlook of the paper. In appendices A and B, we derive the couplings that appear in the calculations.
2 Higgs Extensions of the Standard Model
Two typical Higgs extensions of the Standard Model are studied in this paper. The first model is the Inert Higgs Doublet Model, which is reviewed in subsection 2.1. We then discuss the Two Higgs Doublet Model in subsection 2.2.
2.1 Inert Higgs Doublet Models
In the IHDM, an inert scalar $SU(2)_{L}$ doublet is included in the potential of the SM. The inert scalar particles serve as dark matter candidates. For a review of the theory and phenomenological studies of the IHDM in detail, we refer to the following papers [101, 102, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113]. The scalar potential of the IHDM is expressed as follows:
$\mathcal{V}_{\textrm{IHDM}}(\phi_{1},\phi_{2})$ $=$ $\mu_{1}^{2}|\phi_{1}|^{2}+\mu_{2}^{2}|\phi_{2}|^{2}+\lambda_{1}|\phi_{1}|^{4}+\lambda_{2}|\phi_{2}|^{4}+\lambda_{3}|\phi_{1}|^{2}|\phi_{2}|^{2}+\lambda_{4}|\phi_{1}^{\dagger}\phi_{2}|^{2}$ (1)
$+\frac{\lambda_{5}}{2}\left\{(\phi_{1}^{\dagger}\phi_{2})^{2}+{\rm h.c}\right\}.$
The potential is conserved with respect to the so-called global $Z_{2}$-symmetry, e.g., $\phi_{1}\leftrightarrow+\phi_{1}$, $\phi_{2}\leftrightarrow-\phi_{2}$. In this case, the scalar $\phi_{2}$ is odd, and $\phi_{1}$ and all particles in the SM are even under the $Z_{2}$-transformation. As mentioned, the $Z_{2}$-symmetry remains unbroken after the EWSB. The field $\phi_{2}$ then has a zero vacuum expectation value (VEV), while the field $\phi_{1}$ develops a non-zero VEV ($v$). The two scalar fields are then expanded around their VEVs for the EWSB as follows:
$\phi_{1}$ $=$ $\left(\begin{array}[]{c}G^{\pm}\\ \frac{1}{\sqrt{2}}(v+h+iG^{0})\\ \end{array}\right),$ (4)
$\phi_{2}$ $=$ $\left(\begin{array}[]{c}H^{\pm}\\ \frac{1}{\sqrt{2}}(H+iA^{0})\\ \end{array}\right).$ (7)
Where the VEV is fixed at $v\sim 246$ GeV (as in the SM case). The Goldstone bosons $G^{\pm},\;G^{0}$ give the masses for the $W^{\pm}$ boson and the $Z$ boson, respectively. There is no mixing between $h$ and $H$. The physical spectrum of the IHDM after the EWSB consists of three neutral scalar physical states: two CP-even Higgses $h,\leavevmode\nobreak\ H$ and a CP-odd Higgs $A_{0}$. Furthermore, we have a pair of singly charged Higgs bosons $H^{\pm}$ in this model. In the spectrum, the neutral scalar Higgs $h$ is identical to the SM-like Higgs boson discovered at the LHC. The masses of the scalar bosons are calculated from the model parameters as follows:
$M_{h}^{2}$ $=$ $-2\mu_{1}^{2}=2\lambda_{1}v^{2},$ (8)
$M_{H}^{2}$ $=$ $\mu_{2}^{2}+\frac{v^{2}}{2}\lambda_{L},$ (9)
$M_{A^{0}}^{2}$ $=$ $\mu_{2}^{2}+\frac{v^{2}}{2}\lambda_{R},$ (10)
$M_{H^{\pm}}^{2}$ $=$ $\mu_{2}^{2}+\frac{v^{2}}{2}\lambda_{3}.$ (11)
We have used $\lambda_{L/R}=\lambda_{3}+\lambda_{4}\pm\lambda_{5}$. As mentioned in the previous paragraphs, the global $Z_{2}$-symmetry is unbroken after the EWSB. The "inert" Higgs bosons such as $H^{\pm}$, $A^{0}$, and $H$ have an odd number under the $Z_{2}$-transformation. Subsequently, the "inert" Higgs bosons do not couple to the SM particles. Therefore, the lightest neutral scalar bosons may be considered as dark matter candidates.
As a consequence of the unbroken $Z_{2}$-symmetry, the Yukawa Lagrangian of the IHDM must be the same as that of the SM. In detail, the Yukawa Lagrangian is expressed as follows:
$\mathcal{L}^{\textrm{IHDM}}_{\rm Yukawa}=-\sum_{f=u,d,\ell}g_{hff}\cdot h\bar{f}f+\cdots,$ (12)
where the Yukawa coupling is given by $g_{hff}=m_{f}/v$ for fermion $f$. All the couplings involving the computed processes $\gamma\gamma\rightarrow\phi_{i}\phi_{j}$ in the IHDM are listed in Table 1 (for all physical couplings) and Table 2 (for unphysical particles). We emphasize that the processes $\gamma\gamma\rightarrow hH$ are forbidden in the IHDM due to the $Z_{2}$-symmetry. Therefore, we have only $\phi_{i}\phi_{j}\equiv hh,\leavevmode\nobreak\ HH$ in this case. The detailed expressions for all the concerned couplings are derived in appendix A.
Vertices Notations Coupling
$hW_{\mu}W_{\nu}$ $g_{hWW}\cdot g_{\mu\nu}$ $i\left(\dfrac{2M_{W}^{2}}{v}\right)\cdot g_{\mu\nu}$
$hZ_{\mu}Z_{\nu}$ $g_{hZZ}\cdot g_{\mu\nu}$ $i\left(\dfrac{M_{Z}^{2}}{v}\right)\cdot g_{\mu\nu}$
$hH^{\pm}{H^{\mp}}$ $g_{hH^{\pm}H^{\mp}}$ $i\dfrac{2(\mu_{2}^{2}-M_{H^{\pm}}^{2})}{v}$
$Z_{\mu}H^{\pm}(p_{1})H^{\mp}(p_{2})$ $g_{ZH^{\pm}H^{\mp}}\cdot(p_{1}-p_{2})_{\mu}$ $i\left(\dfrac{M_{Z}c_{2W}}{v}\right)\cdot(p_{1}-p_{2})_{\mu}$
$A_{\mu}H^{\pm}(p_{1})H^{\mp}(p_{2})$ $g_{AH^{\pm}H^{\mp}}\cdot(p_{1}-p_{2})_{\mu}$ $i\left(\dfrac{M_{Z}s_{2W}}{v}\right)\cdot(p_{1}-p_{2})_{\mu}$
$hhh$ $g_{hhh}$ $-i\left(\dfrac{3M_{h}^{2}}{v}\right)$
$hHH$ $g_{hHH}$ $i\dfrac{2(\mu_{2}^{2}-M_{H}^{2})}{v}$
$H(p_{2})H^{\pm}(p_{1})W^{\mp}_{\mu}$ $g_{HH^{\pm}W}\cdot(p_{1}-p_{2})_{\mu}$$\pm i\left(\dfrac{M_{W}}{v}\right)\cdot(p_{1}-p_{2})_{\mu}$
$H^{\pm}H^{\mp}A_{\mu}A_{\nu}$ $g_{AAH^{\pm}H^{\mp}}\cdot g_{\mu\nu}$ $i\left(\dfrac{M_{Z}^{2}s_{2W}^{2}}{v^{2}}\right)\cdot g_{\mu\nu}$
$HH^{\pm}W^{\mp}_{\nu}A_{\mu}$ $g_{HH^{\pm}WA}\cdot g_{\mu\nu}$ $i\left(\dfrac{2M_{Z}^{2}c_{W}^{2}s_{W}}{v^{2}}\right)\cdot g_{\mu\nu}$
$hhH^{\pm}H^{\mp}$ $g_{hhH^{\pm}H^{\mp}}$ $i\dfrac{2(\mu_{2}^{2}-M_{H^{\pm}}^{2})}{v^{2}}$
$HHH^{\pm}H^{\mp}$ $g_{HHH^{\pm}H^{\mp}}$ $-2i\lambda_{2}$
$hhW^{\pm}_{\mu}W^{\mp}_{\nu}$ $g_{hhWW}\cdot g_{\mu\nu}$ $i\left(\dfrac{2M_{Z}^{2}c_{W}^{2}}{v^{2}}\right)\cdot g_{\mu\nu}$
$HHW^{\pm}_{\mu}W^{\mp}_{\nu}$ $g_{HHWW}\cdot g_{\mu\nu}$ $i\left(\dfrac{2M_{Z}^{2}c_{W}^{2}}{v^{2}}\right)\cdot g_{\mu\nu}$
Table 1: We list all couplings (physical couplings) contributing to $\gamma\gamma\rightarrow\phi_{i}\phi_{j}$ for the IHDM.
Vertices Notations Coupling
$A_{\mu}W^{\pm}_{\nu}G^{\mp}$ $g_{AW^{\pm}G^{\mp}}\cdot g_{\mu\nu}$ $i\left(\dfrac{2M_{Z}^{2}c_{W}^{2}s_{W}}{v}\right)\cdot g_{\mu\nu}$
$HH^{\pm}G^{\mp}$ $g_{HH^{\pm}G^{\mp}}$ $i\dfrac{M_{H^{\pm}}^{2}-M_{H}^{2}}{v}$
$hhG^{\pm}G^{\mp}$ $g_{hhG^{\pm}G^{\mp}}$ $-i\left(\dfrac{M_{h}^{2}}{v^{2}}\right)$
$HHG^{\pm}G^{\mp}$ $g_{HHG^{\pm}G^{\mp}}$ $i\dfrac{2(\mu_{2}^{2}-M_{H^{\pm}}^{2})}{v^{2}}$
$A_{\mu}A_{\nu}G^{\pm}G^{\mp}$ $g_{AAG^{\pm}G^{\mp}}\cdot g_{\mu\nu}$ $i\left(\dfrac{M_{Z}^{2}s_{2W}^{2}}{v^{2}}\right)\cdot g_{\mu\nu}$
$A_{\mu}G^{\pm}(p_{1})G^{\mp}(p_{2})$ $g_{AG^{\pm}G^{\mp}}\cdot(p_{1}-p_{2})_{\mu}$ $i\left(\dfrac{M_{Z}s_{2W}}{v}\right)\cdot(p_{1}-p_{2})_{\mu}$
Table 2: We list all vertices (unphysical couplings) in processes $\gamma\gamma\rightarrow\phi_{i}\phi_{j}$ in the IHDM.
The parameter space of the IHDM for our analysis is included as follows:
$\mathcal{P}_{\rm IHDM}=\{\mu_{2}^{2},\lambda_{2}^{2},M_{h}^{2}\sim 125.\textrm{GeV},M_{H}^{2},M_{A^{0}}^{2},M^{2}_{H^{\pm}}\}.$ (13)
We are going to review the current constraints on the physical parameter space in the IHDM given in Eq. 13. The constraints for the physical parameter space can be obtained by including the theoretical conditions and the experimental data. In the perspective of the experimental data, we take into account the Electroweak Precision Tests (EWPT) of the IHDM, dark matter search at the LHC, as well as the LEP data. The topics have been studied in Refs. [101, 102, 103, 104]. Additionally, the implications for loop-induced decays of the SM-like Higgs ($h$) to $V\gamma$ with $V\equiv Z,\gamma$ in the IHDM framework, e.g. the decay process $h\rightarrow\gamma\gamma$, have been reported in [105, 112, 113], and decay channels $h\rightarrow Z\gamma$ have been examined in [112, 113]. Furthermore, searching for signals of the IHDM at future colliders has been performed in Refs. [106, 107, 108, 109, 110, 111]. In the theoretical side, the most important theoretical constraints are obtained from the conditions that the models follow tree-level unitarity, vacuum stability, and the perturbative regime. The theoretical constraints give the limitations on the Higgs self-couplings $\lambda_{i}$ for $i=1,2,\cdots,5$ and $\mu_{2}$. Taking the theoretical and experimental constraints from the above-mentioned papers, one can select the parameter space for the IHDM as follows: we can take $5$ GeV $\leq M_{H}\leq 150$ GeV, $70$ GeV $\leq M_{H^{\pm}},M_{A^{0}}\leq 1000$ GeV, $|\mu_{2}|\leq 500$ GeV, and $0\leq\lambda_{2}\leq 8\pi$.
2.2 Two Higgs Doublet Models
We now considerthe second kind of Higgs sector extension of the Standard Model, namely the Two Higgs Doublet Models, in this paper. For a comprehensive review of the theory and phenomenological studies of the THDM, we refer the reader to Ref. [114] for further information. The model is briefly summarized in this section. A complex Higgs doublet with hypercharge $Y=1/2$ is added to the scalar sector of the SM. The scalar potential takes the form:
$\mathcal{V}_{\textrm{THDM}}(\phi_{1},\phi_{2})$ $=$ $m_{11}^{2}\phi_{1}^{\dagger}\phi_{1}+m_{22}^{2}\phi_{2}^{\dagger}\phi_{2}-\Big{[}m_{12}^{2}\phi_{1}^{\dagger}\phi_{2}+{\rm h.c.}\Big{]}+\frac{\lambda_{1}}{2}(\phi_{1}^{\dagger}\phi_{1})^{2}+\frac{\lambda_{2}}{2}(\phi_{2}^{\dagger}\phi_{2})^{2}$ (14)
$+\lambda_{3}(\phi_{1}^{\dagger}\phi_{1})(\phi_{2}^{\dagger}\phi_{2})+\lambda_{4}(\phi_{1}^{\dagger}\phi_{2})(\phi_{2}^{\dagger}\phi_{1})+\frac{1}{2}[\lambda_{5}\leavevmode\nobreak\ (\phi_{1}^{\dagger}\phi_{2})^{2}+\leavevmode\nobreak\ {\rm h.c}].$
[…]
Two scalar doublets are expanded around their VEVs after electroweak symmetry breaking (EWSB) as
$\phi_{k}$ $=$ $\begin{bmatrix}\rho_{k}^{+}\\ (v_{k}+\eta_{k}+i\xi_{k})/\sqrt{2}\end{bmatrix}\quad\textrm{for}\quad k=1,2.$ (15)
The vacuum expectation value is then fixed at $v=\sqrt{v_{1}^{2}+v_{2}^{2}}\sim 246$ GeV, in agreement with the SM value. After the EWSB, the physical particles in the THDM consist of two CP-even Higgs bosons, $h$ and $H$, with one of them, $h$, being identified as the SM-like Higgs boson discovered at the LHC, a CP-odd Higgs boson ($A^{0}$), and a pair of charged Higgs bosons ($H^{\pm}$). To obtain the physical masses of the new scalar bosons, we perform the following rotations:
$\begin{pmatrix}\eta_{1}\\ \eta_{2}\end{pmatrix}$ $=$ $\begin{pmatrix}c_{\alpha}&-s_{\alpha}\\ s_{\alpha}&c_{\alpha}\end{pmatrix}\begin{pmatrix}H\\ h\end{pmatrix},$ (16)
$\begin{pmatrix}\rho_{1}^{\pm}\\ \rho_{2}^{\pm}\end{pmatrix}$ $=$ $\begin{pmatrix}c_{\beta}&-s_{\beta}\\ s_{\beta}&c_{\beta}\end{pmatrix}\begin{pmatrix}G^{\pm}\\ H^{\pm}\end{pmatrix},$ (17)
$\begin{pmatrix}\xi_{1}\\ \xi_{2}\end{pmatrix}$ $=$ $\begin{pmatrix}c_{\beta}&-s_{\beta}\\ s_{\beta}&c_{\beta}\end{pmatrix}\begin{pmatrix}G^{0}\\ A^{0}\end{pmatrix}.$ (18)
The mixing angle $\beta$ is defined by $t_{\beta}\equiv\tan\beta=v_{2}/v_{1}$. The physical masses of the Higgs bosons are then expressed in terms of the model parameters as follows:
$M_{H^{\pm}}^{2}$ $=$ $M^{2}-\frac{1}{2}\lambda_{45}v^{2},$ (19)
$M_{A^{0}}^{2}$ $=$ $M^{2}-\lambda_{5}v^{2},$ (20)
$M_{h}^{2}$ $=$ $M_{11}^{2}s_{\beta-\alpha}^{2}+M_{22}^{2}c_{\beta-\alpha}^{2}+M_{12}^{2}s_{2(\beta-\alpha)},$ (21)
$M_{H}^{2}$ $=$ $M_{11}^{2}c_{\beta-\alpha}^{2}+M_{22}^{2}s_{\beta-\alpha}^{2}-M_{12}^{2}s_{2(\beta-\alpha)}$ (22)
where the parameter $M^{2}$ is used as $M^{2}=m_{12}^{2}/(s_{\beta}c_{\beta})$. The elements $M_{ij}$ for $i,j=1,2$ are given by
$M_{11}^{2}$ $=$ $(\lambda_{1}c_{\beta}^{4}+\lambda_{2}s_{\beta}^{4})v^{2}+\frac{v^{2}}{2}\;\lambda_{345}\;s_{2\beta}^{2},$ (23)
$M_{22}^{2}$ $=$ $M^{2}+\frac{v^{2}}{4}\Big{[}\lambda_{12}-2\lambda_{345}\Big{]}s_{2\beta}^{2},$ (24)
$M_{12}^{2}$ $=$ $M_{21}^{2}=-\frac{v^{2}}{2}\Big{[}\lambda_{1}c_{\beta}^{2}-\lambda_{2}s_{\beta}^{2}-\lambda_{345}\;c_{2\beta}\Big{]}s_{2\beta}.$ (25)
Here, the shorten notation have used as $\lambda_{ij\cdots}=\lambda_{i}+\lambda_{j}+\cdots$.
We present the relevant couplings involved in the amplitude computations for the processes $\gamma\gamma\rightarrow\phi_{i}\phi_{j}$ in Tables 34 (for physical couplings) and in Table 5 (for unphysical couplings). These couplings are derived in Appendix B.
Vertices Notations Couplings
$hW_{\mu}W_{\nu}$ $g_{hWW}\cdot g_{\mu\nu}$ $i\left(\dfrac{2M_{W}^{2}}{v}\;s_{\beta-\alpha}\right)\cdot g_{\mu\nu}$
$HW_{\mu}W_{\nu}$ $g_{HWW}\cdot g_{\mu\nu}$ $i\left(\dfrac{2M_{W}^{2}}{v}\;c_{\beta-\alpha}\right)\cdot g_{\mu\nu}$
$hZ_{\mu}Z_{\nu}$ $g_{hZZ}\cdot g_{\mu\nu}$ $i\left(\dfrac{2M_{Z}^{2}}{v}\;s_{\beta-\alpha}\right)\cdot g_{\mu\nu}$
$HZ_{\mu}Z_{\nu}$ $g_{HZZ}\cdot g_{\mu\nu}$ $i\left(\dfrac{2M_{Z}^{2}}{v}\;c_{\beta-\alpha}\right)\cdot g_{\mu\nu}$
$hH^{\pm}H^{\mp}$ $g_{hH^{\pm}H^{\mp}}$ $i\Bigg{[}\dfrac{c_{\alpha+\beta}(4M^{2}-3M_{h}^{2}-2M_{H^{\pm}}^{2})}{2vs_{2\beta}}$
                  $+\dfrac{(2M_{H^{\pm}}^{2}-M_{h}^{2})c_{(\alpha-3\beta)}}{2vs_{2\beta}}\Bigg{]}$
$Z_{\mu}H^{\pm}(p_{1})H^{\mp}(p_{2})$ $g_{ZH^{\pm}H^{\mp}}\cdot(p_{1}-p_{2})_{\mu}$ $i\left(\dfrac{M_{Z}\;c_{2W}}{v}\right)\cdot(p_{1}-p_{2})_{\mu}$
$A_{\mu}H^{\pm}(p_{1})H^{\mp}(p_{2})$ $g_{AH^{\pm}H^{\mp}}\cdot(p_{1}-p_{2})_{\mu}$ $i\left(\dfrac{M_{Z}\;s_{2W}}{v}\right)\cdot(p_{1}-p_{2})_{\mu}$
$hhh$ $g_{hhh}$ $i\dfrac{3e}{4M_{W}s_{W}s_{2\beta}}\Bigg{[}M^{2}c_{\alpha-3\beta}+(M^{2}-M_{h}^{2})c_{3\alpha-\beta}$
                                    $+(2M^{2}-3M_{h}^{2})c_{\alpha+\beta}\Bigg{]}$
$HHH$ $g_{HHH}$ $i\dfrac{3e}{4M_{W}s_{W}s_{2\beta}}\Bigg{[}M^{2}s_{\alpha-3\beta}+(M_{H}^{2}-M^{2})s_{3\alpha-\beta}$
                                    $+(2M^{2}-3M_{H}^{2})s_{\alpha+\beta}\Bigg{]}$
$hHH$ $g_{hHH}$ $i\dfrac{\Big{[}s_{2\alpha}(3M^{2}-M_{h}^{2}-2M_{H}^{2})+M^{2}s_{2\beta}\Big{]}}{v\;s_{2\beta}}s_{\alpha-\beta}$
$Hhh$ $g_{Hhh}$ $i\dfrac{\Big{[}s_{2\alpha}(3M^{2}-M_{H}^{2}-2m_{h}^{2})-M^{2}s_{2\beta}\Big{]}}{2vs_{2\beta}}c_{\alpha-\beta}$
Table 3: The physical couplings contributing to the considered processes $\gamma\gamma\rightarrow\phi_{i}\phi_{j}$ in the THDM.
Vertices Notations Couplings
$HH^{\pm}H^{\mp}$ $g_{HH^{\pm}H^{\mp}}$ $i\Bigg{[}\dfrac{s_{\alpha+\beta}(4M^{2}-3M_{H}^{2}-2M_{H^{\pm}}^{2})}{2v\;s_{2\beta}}$
                  $+\dfrac{(2M_{H^{\pm}}^{2}-M_{H}^{2})s_{\alpha-3\beta}}{2v\;s_{2\beta}}\Bigg{]}$
$H(p_{1})H^{\pm}(p_{2})W^{\mp}_{\mu}$ $g_{HH^{\pm}W}\cdot(p_{1}-p_{2})_{\mu}$$\pm i\left(\dfrac{M_{W}}{v}s_{\beta-\alpha}\right)\cdot(p_{1}-p_{2})_{\mu}$
$h(p_{1})H^{\pm}(p_{2})W^{\mp}_{\mu}$ $g_{hH^{\pm}W}\cdot(p_{1}-p_{2})_{\mu}$ $\mp i\left(\dfrac{M_{W}}{v}c_{\alpha-\beta}\right)\cdot(p_{1}-p_{2})_{\mu}$
$H^{\pm}H^{\mp}A_{\mu}A_{\nu}$ $g_{H^{\pm}H^{\mp}AA}\cdot g_{\mu\nu}$ $i\left(\dfrac{4M_{W}^{2}s_{W}^{2}}{v^{2}}\right)\cdot g_{\mu\nu}$
$HH^{\pm}W^{\mp}_{\mu}A_{\nu}$ $g_{HH^{\pm}WA}\cdot g_{\mu\nu}$ $i\left(\dfrac{2M_{W}^{2}s_{W}}{v^{2}}s_{\alpha-\beta}\right)\cdot g_{\mu\nu}$
$hH^{\pm}W^{\mp}_{\mu}A_{\nu}$ $g_{hH^{\pm}WA}\cdot g_{\mu\nu}$ $i\left(\dfrac{2M_{W}^{2}s_{W}}{v^{2}}c_{\alpha-\beta}\right)\cdot g_{\mu\nu}$
$hHH^{\pm}H^{\mp}$ $g_{HhH^{\pm}H^{\mp}}$ $i\lambda_{HhH^{\pm}H^{\mp}}$ [in Eq. (42)]
$HHH^{\pm}H^{\mp}$ $g_{HHH^{\pm}H^{\mp}}$ $i\lambda_{HHH^{\pm}H^{\mp}}$ [in Eq. (46)]
$hhH^{\pm}H^{\mp}$ $g_{hhH^{\pm}H^{\mp}}$ $i\lambda_{hhH^{\pm}H^{\mp}}$ [in Eq. (47)]
$hhW^{\pm}_{\mu}W^{\mp}_{\nu}$ $g_{hhWW}\cdot g_{\mu\nu}$ $i\left(\dfrac{4M_{W}^{2}}{v^{2}}\right)\cdot g_{\mu\nu}$
$HHW^{\pm}_{\mu}W^{\mp}_{\nu}$ $g_{HHWW}\cdot g_{\mu\nu}$ $i\left(\dfrac{4M_{W}^{2}}{v^{2}}\right)\cdot g_{\mu\nu}$
Table 4: Additional, the physical couplings contributing to the considered processes $\gamma\gamma\rightarrow\phi_{i}\phi_{j}$ in the THDM.
Vertices Notations Couplings
$A_{\mu}W^{\pm}_{\nu}G^{\mp}$ $g_{AW^{\pm}G^{\mp}}\cdot g_{\mu\nu}$ $i\left(\dfrac{2M_{W}^{2}s_{W}}{v}\right)\cdot g_{\mu\nu}$
$HH^{\pm}G^{\mp}$ $g_{HH^{\pm}G^{\mp}}$ $i\left(\dfrac{e}{2M_{W}s_{W}}s_{\alpha-\beta}\right)(M_{H^{\pm}}^{2}-M_{H}^{2})$
$A_{\mu}A_{\nu}G^{\pm}G^{\mp}$ $g_{AAG^{\pm}G^{\mp}}\cdot g_{\mu\nu}$ $i\left(\dfrac{4M_{W}^{2}s_{W}^{2}}{v^{2}}\right)\cdot g_{\mu\nu}$
$A_{\mu}G^{\pm}(p_{1})G^{\mp}(p_{2})$ $g_{AG^{\pm}G^{\mp}}\cdot(p_{1}-p_{2})_{\mu}$ $i\left(\dfrac{2M_{W}s_{W}}{v}\right)\cdot(p_{1}-p_{2})_{\mu}$
$hHG^{\pm}G^{\mp}$ $g_{hHG^{\pm}G^{\mp}}$ $i\lambda_{hHG^{\pm}G^{\mp}}$ [in Eq. (43)]
$hhG^{\pm}G^{\mp}$ $g_{hhG^{\pm}G^{\mp}}$ $i\lambda_{hhG^{\pm}G^{\mp}}$ [in Eq. (48) ]
$HHG^{\pm}G^{\mp}$ $g_{HHG^{\pm}G^{\mp}}$ $i\lambda_{HHG^{\pm}G^{\mp}}$ [in Eq. (49)]
Table 5: Unphysical couplings involving to the processes under investigations are shown.
Finally, we pay attention to the Yukawa sector in the THDM. In order to avoid Tree-level Flavor-Changing Neutral Currents (FCNCs), the discrete $Z_{2}$-symmetry may be proposed in the THDM as in [115]. The $Z_{2}$-parity assignments for all fermions and the definition for four types of the THDM based on the couple of the scalar with fermions are shown in [125]. The Yukawa Lagrangian is then written in the mass eigenstates as in [114]
${\mathcal{L}}_{\rm Yukawa}=-\sum_{f=u,d,\ell}\left(g_{hff}\cdot\bar{f}fh+g_{Hff}\cdot\bar{f}fH-ig_{A^{0}ff}\cdot\bar{f}\gamma_{5}fA^{0}\right)+\cdots,$ (26)
We show the Yukawa couplings of CP-even with fermions for four types of the THDM in Table 6, see [114, 126] for further detail.
Type $g_{huu}$ $g_{hdd}$ $g_{h\ell\ell}$ $g_{Huu}$ $g_{Hdd}$ $g_{H\ell\ell}$
I $\dfrac{m_{u}}{\sqrt{2}v}\dfrac{c_{\alpha}}{s_{\beta}}$ $\dfrac{m_{d}}{\sqrt{2}v}\dfrac{c_{\alpha}}{s_{\beta}}$ $\dfrac{m_{\ell}}{\sqrt{2}v}\dfrac{c_{\alpha}}{s_{\beta}}$ $\dfrac{m_{u}}{\sqrt{2}v}\dfrac{s_{\alpha}}{s_{\beta}}$ $\dfrac{m_{d}}{\sqrt{2}v}\dfrac{s_{\alpha}}{s_{\beta}}$ $\dfrac{m_{\ell}}{\sqrt{2}v}\dfrac{s_{\alpha}}{s_{\beta}}$
II $\dfrac{m_{u}}{\sqrt{2}v}\dfrac{c_{\alpha}}{s_{\beta}}$ $-\dfrac{m_{d}}{\sqrt{2}v}\dfrac{s_{\alpha}}{c_{\beta}}$ $-\dfrac{m_{\ell}}{\sqrt{2}v}\dfrac{s_{\alpha}}{c_{\beta}}$ $\dfrac{m_{u}}{\sqrt{2}v}\dfrac{s_{\alpha}}{s_{\beta}}$ $-\dfrac{m_{d}}{\sqrt{2}v}\dfrac{c_{\alpha}}{c_{\beta}}$ $-\dfrac{m_{\ell}}{\sqrt{2}v}\dfrac{c_{\alpha}}{c_{\beta}}$
X $\dfrac{m_{u}}{\sqrt{2}v}\dfrac{c_{\alpha}}{s_{\beta}}$ $\dfrac{m_{d}}{\sqrt{2}v}\dfrac{c_{\alpha}}{s_{\beta}}$ $-\dfrac{m_{\ell}}{\sqrt{2}v}\dfrac{s_{\alpha}}{c_{\beta}}$ $\dfrac{m_{u}}{\sqrt{2}v}\dfrac{s_{\alpha}}{s_{\beta}}$ $\dfrac{m_{d}}{\sqrt{2}v}\dfrac{s_{\alpha}}{s_{\beta}}$ $-\dfrac{m_{\ell}}{\sqrt{2}v}\dfrac{c_{\alpha}}{c_{\beta}}$
Y $\dfrac{m_{u}}{\sqrt{2}v}\dfrac{c_{\alpha}}{s_{\beta}}$ $-\dfrac{m_{d}}{\sqrt{2}v}\dfrac{s_{\alpha}}{c_{\beta}}$ $\dfrac{m_{\ell}}{\sqrt{2}v}\dfrac{c_{\alpha}}{s_{\beta}}$ $\dfrac{m_{u}}{\sqrt{2}v}\dfrac{s_{\alpha}}{s_{\beta}}$ $-\dfrac{m_{d}}{\sqrt{2}v}\dfrac{c_{\alpha}}{c_{\beta}}$ $\dfrac{m_{\ell}}{\sqrt{2}v}\dfrac{s_{\alpha}}{s_{\beta}}$
Table 6: We show all the Yukawa couplings of CP-even Higges to fermions for all types of THDMs.
The parameter space $\mathcal{P}_{\rm THDM}$ for THDM is include as follows
$\mathcal{P}_{\rm THDM}=\{M_{h}^{2}\sim 125.\textrm{GeV},M_{H}^{2},M_{A^{0}}^{2},M^{2}_{H^{\pm}},m_{12}^{2},t_{\beta},s_{\beta-\alpha}\}.$ (27)
As in the case of the IHDM, we first summarize the current constraints on the parameter space of the THDM, given in Eq. 27. Both theoretical conditions and experimental data are taken into account, from which the allowed regions of the parameter space of the THDM are determined. Theoretical constraints arise from the requirements that the model remains within the perturbative regime, satisfies tree-level unitarity, and ensures vacuum stability of the scalar potential. These topics have been discussed in detail in Refs. [116, 117, 118, 119, 121] and references therein. We also consider electroweak precision tests (EWPT) in the context of the THDM, based on experimental data. Relevant implications for these constraints from LEP are reported in Refs. [122, 123]. The allowed mass ranges of the scalar particles in the THDM have been investigated at LEP, the Tevatron, and the LHC, as reviewed in Ref. [120]. Furthermore, one-loop induced decays of the SM-like Higgs boson, such as $h\rightarrow\gamma\gamma$ and $h\rightarrow Z\gamma$, have also been analyzed within the THDM in Refs. [112, 113] and references therein. By combining all of the above constraints, the physical parameter ranges can be chosen as follows: $126\leavevmode\nobreak\ \text{GeV}\leq M_{H}\leq 1000\leavevmode\nobreak\ \text{GeV}$, $60\leavevmode\nobreak\ \text{GeV}\leq M_{A^{0}}\leq 1000\leavevmode\nobreak\ \text{GeV}$, and $80\leavevmode\nobreak\ \text{GeV}\leq M_{H^{\pm}}\leq 1000\leavevmode\nobreak\ \text{GeV}$ for Type-I and Type-X THDMs. For Type-II and Type-Y, the physical parameters are scanned in the ranges: $500\leavevmode\nobreak\ \text{GeV}\leq M_{H}\leq 1000\leavevmode\nobreak\ \text{GeV}$, $500\leavevmode\nobreak\ \text{GeV}\leq M_{A^{0}}\leq 1000\leavevmode\nobreak\ \text{GeV}$, and $580\leavevmode\nobreak\ \text{GeV}\leq M_{H^{\pm}}\leq 1500\leavevmode\nobreak\ \text{GeV}$. The soft $Z_{2}$-breaking parameter is chosen as $m_{12}^{2}=M_{H}^{2}s_{\beta}c_{\beta}$. Finally, additional constraints in the $(t_{\beta},M_{H^{\pm}})$ plane are examined by incorporating flavor physics data, as shown in Ref. [124]. The results in Ref. [124] indicate that small values of $t_{\beta}$ are favored to be consistent with flavor experimental constraints. In our complementary analysis, we also explore the phenomenology associated with low $t_{\beta}$ scenarios.
3 Phenomenological results
The detailed calculations for $\gamma\gamma\rightarrow\phi_{i}\phi_{j}$ with CP-even Higgses $\phi_{i,j}\equiv h,H$ in the HESM have been reported in our previous work [100]. In this paper, we focus on the phenomenological analysis of the concerned processes in the IHDM and the THDM. We are interested in examining the enhancement factors $\mu_{\phi_{i}\phi_{j}}^{\textrm{NP}}$ with NP standing for the THDM and the IHDM, respectively, defined as the ratio of the cross-sections of $\gamma\gamma\rightarrow\phi_{i}\phi_{j}$ in the HESM to the corresponding ones for $\gamma\gamma\rightarrow hh$ in the SM. The factors are given explicitly by
$\mu_{\phi_{i}\phi_{j}}^{\textrm{NP}}=\frac{\hat{\sigma}^{\textrm{NP}}_{\gamma\gamma\rightarrow\phi_{i}\phi_{j}}}{\hat{\sigma}^{\textrm{SM}}_{\gamma\gamma\rightarrow hh}}(\mathcal{P}_{\textrm{NP}}).$ (28)
In this work, the enhancement factors are examined within the parameter space of the THDM and the IHDM. Forthe phenomenological results, all physical input parameters in the SM are taken to be the same as those in [125, 126].
3.1 IHDM
Phenomenological studies for the processes $\gamma\gamma\rightarrow\phi_{i}\phi_{j}$ in the IHDM are presented in this subsection. In the IHDM, the process $\gamma\gamma\rightarrow hH$ is forbidden by the $Z_{2}$-symmetry. For this reason, we only focus on physical results for the processes $\gamma\gamma\rightarrow hh,\leavevmode\nobreak\ HH$ in the IHDM. We note that all physical results in the IHDM presented in the following subsections can be considered as the first outcomes from this work.
3.1.1 Production cross-sections
In Fig. 1, we show cross-sections for $\gamma\gamma\rightarrow hh,\leavevmode\nobreak\ HH$ in the IHDM, together with those for $hh$ production in the SM, as functions of center-of-mass energy (CoM, or $\sqrt{\hat{s}_{\gamma\gamma}}$). For the generated data, we select the following parameter space in the IHDM: $M_{H^{\pm}}=200$ GeV, $M_{H}=150$ GeV, and fix $\lambda_{2}=0.8$ for all cases. We vary $350$ GeV $\leq\sqrt{\hat{s}_{\gamma\gamma}}\leq 1500$ GeV in the plots. Cross-sections are presented for $\mu_{2}^{2}=0$ GeV2 on the left panel and for $200^{2}$ GeV2 on the right panel, respectively. In the plots, the black line represents the cross-sections for $\gamma\gamma\rightarrow hh$ in the SM, and the blue (green) line represents the cross-sections for $\gamma\gamma\rightarrow hh,\leavevmode\nobreak\ (HH)$ in the IHDM, respectively.
We first comment on the results in the case of $\mu_{2}^{2}=0$ GeV2. The cross-sections for $hh,\leavevmode\nobreak\ HH$ in the IHDM have peaks at $\sqrt{\hat{s}_{\gamma\gamma}}\sim 2M_{H^{\pm}}=400$ GeV. In the regions $\sqrt{\hat{s}_{\gamma\gamma}}\leq 750$ GeV, $\hat{\sigma}_{hh,HH}$ in the IHDM are larger than $\hat{\sigma}_{hh}$ in the SM. Beyond the regions of $\sqrt{\hat{s}_{\gamma\gamma}}\geq 750$ GeV, the cross-sections for $HH$ in the IHDM are suppressed in comparison with $hh$ production in both the SM and the IHDM. It is interesting to find that the production cross-sections for $\gamma\gamma\rightarrow hh,\leavevmode\nobreak\ HH$ in the IHDM are dominant around the peaks compared with $\hat{\sigma}_{hh}$ in the SM. This indicates that the contributions from singly charged Higgs in the loop of $\gamma\gamma\rightarrow hh,\leavevmode\nobreak\ HH$ are significant in these regions.
In the case of $\mu_{2}^{2}=200^{2}$ GeV2, we only observe a peak of $\hat{\sigma}_{HH}$ in the IHDM around $\sqrt{\hat{s}_{\gamma\gamma}}\sim 2M_{H^{\pm}}=400$ GeV. Furthermore, the data shows that the cross-sections for $HH$ production are dominant in the regions $\sqrt{\hat{s}_{\gamma\gamma}}\leq 750$ GeV, in contrast with the corresponding ones for $hh$ production in both the SM and the IHDM. Beyond the regions $\sqrt{\hat{s}_{\gamma\gamma}}\geq 750$ GeV, $\hat{\sigma}_{HH}$ is suppressed. It is observed that the cross-sections for $hh$ production in the IHDM are smaller than $\hat{\sigma}_{hh}$ in the SM when $\sqrt{\hat{s}_{\gamma\gamma}}\leq 550$ GeV. In the regions of $\sqrt{\hat{s}_{\gamma\gamma}}\geq 550$ GeV, $\hat{\sigma}_{hh}$ in the IHDM tends to the cross-sections for $hh$ production in the SM. This can be explained as follows. Since the $hh$ production in the IHDM are different from those in the SM by the contributions of charged Higgs in the loop of triangle $h^{*}$-pole and box diagrams. These contributions depend on $M_{H^{\pm}}$ and the vertices $hH^{\pm}H^{\mp}$, $hhH^{\pm}H^{\mp}$, expressed in terms of $\mu_{2}^{2}$. At large values of $\mu_{2}^{2}$, these contributions may be cancelled out. As a result, cross-sections for $hh$ production in the IHDM tend to the corresponding ones in the SM. In another case of $HH$ production, we have no couplings of $HH^{\pm}H^{\mp}$ due to the $Z_{2}$-symmetry and the vertex $HHH^{\pm}H^{\mp}$ depends on $\lambda_{2}$. Therefore, we have no such cancellations as mentioned. It is reasonable that the cross-sections for $HH$ production in the IHDM are dominant in the regions $\sqrt{\hat{s}_{\gamma\gamma}}\leq 550$ GeV and they also have a peak at $\sqrt{\hat{s}_{\gamma\gamma}}\sim 2M_{H^{\pm}}=400$ GeV.
$\hat{\sigma}_{\phi_{i}\phi_{j}}(\mu_{2}^{2}=0)$ [pb] $\hat{\sigma}_{\phi_{i}\phi_{j}}(\mu_{2}^{2}=200^{2})$ [pb]
Refer to caption Refer to caption
                                                     $\sqrt{\hat{s}_{\gamma\gamma}}$ [GeV]                                                      $\sqrt{\hat{s}_{\gamma\gamma}}$ [GeV]
Figure 1: Total cross-sections for $\gamma\gamma\rightarrow hh,\leavevmode\nobreak\ HH$ in the SM and IHDM are presented as functions of $\sqrt{\hat{s}_{\gamma\gamma}}$. For the generated data, we select $M_{H^{\pm}}=200$ GeV, $M_{H}=150$ GeV in this case. We vary $350$ GeV $\leq\sqrt{\hat{s}_{\gamma\gamma}}\leq 1500$ GeV. In the plots below, we fix $\mu_{2}^{2}=0,200^{2}$ GeV2 and $\lambda_{2}=0.8$ for all cases. In the plots, the black line shows $\hat{\sigma}_{hh}$ in the SM. Additionally, the blue (green) line presents $\gamma\gamma\rightarrow hh\leavevmode\nobreak\ (HH)$ in the THDM.
3.1.2 Enhancement factors
The enhancement factors given in Eq. 28 are examined in the IHDM. In Fig. 2, the factors for $\gamma\gamma\rightarrow hh,\leavevmode\nobreak\ HH$ are scanned in the parameter space of $M_{H^{\pm}},\leavevmode\nobreak\ \mu_{2}^{2}$. In the following scatter plots, singly charged Higgs masses are varied from $70$ GeV $\leq M_{H^{\pm}}\leq 600$ GeV and $-200$ GeV $\leq\mu_{2}\leq 200$ GeV. Furthermore, we fix $\lambda_{2}=0.8$ and $M_{H}=150$ GeV for all cases. The data are generated at $\sqrt{\hat{s}_{\gamma\gamma}}=500$ (all left panel scatter-plots) GeV and at $\sqrt{\hat{s}_{\gamma\gamma}}=1000$ GeV (all right panel scatter-plots).
The factors for $hh$ production in the IHDM are first analyzed. Since the cross-sections for $hh$ production are enhanced around the peaks at $\sqrt{\hat{s}_{\gamma\gamma}}\sim 2M_{H^{\pm}}$, it is not surprising to find that $\mu_{hh}^{\textrm{IHDM}}$ becomes largest at $\sqrt{\hat{s}_{\gamma\gamma}}\sim 2M_{H^{\pm}}=250$ GeV (for the left plots) and at $\sqrt{\hat{s}_{\gamma\gamma}}\sim 2M_{H^{\pm}}=500$ GeV (for the right plots). Around these peaks, the data indicate that the enhancement factors tend to about $\sim 1.5$ in the limit of $\mu_{2}^{2}\rightarrow M_{H^{\pm}}^{2}$, since the contributions of the charged Higgs in the loop are small when $\mu_{2}^{2}\rightarrow M_{H^{\pm}}^{2}$ (due to the fact that the couplings of $hH^{\pm}H^{\mp}$ and $hhH^{\pm}H^{\mp}$ tend to zero in this limit). It is found that the enhancement factors can reach a factor of $6$ (for $500$ GeV of CoM) and a factor of $13$ (for $1000$ GeV of CoM) around the peaks. Beyond the peaks, we observe that $1\leq\mu_{hh}^{\textrm{IHDM}}\leq 2$.
For the enhancement factors of $HH$ production in the IHDM, we also find that $\mu_{HH}^{\textrm{IHDM}}$ becomes largest at $\sqrt{\hat{s}_{\gamma\gamma}}\sim 2M_{H^{\pm}}=250$ GeV (for the left plots) and at $\sqrt{\hat{s}_{\gamma\gamma}}\sim 2M_{H^{\pm}}=500$ GeV (for the right plots). It is important to realize that $\mu_{HH}^{\textrm{IHDM}}$ has different behavior in comparison with $\mu_{hh}^{\textrm{IHDM}}$. At $500$ GeV of CoM, the factors rise to the peak and then decrease rapidly beyond the peak. However, they grow with the charged Higgs masses in the above regions of $M_{H^{\pm}}\geq\sim 300$ GeV. Because there are no couplings of $HH^{\pm}H^{\mp}$ due to the $Z_{2}$-symmetry and the vertex $HHH^{\pm}H^{\mp}$ only depends on $\lambda_{2}$, the factors $\mu_{HH}^{\textrm{IHDM}}$ increase with $M_{H^{\pm}}$ in the high regions of charged Higgs masses.
$\mu_{hh}^{\textrm{IHDM}}$ $\mu_{hh}^{\textrm{IHDM}}$
Refer to caption Refer to caption
                                            $M_{H^{\pm}}$ [GeV]                                             $M_{H^{\pm}}$ [GeV]
$\mu_{HH}^{\textrm{IHDM}}$ $\mu_{HH}^{\textrm{IHDM}}$
Refer to caption Refer to caption
                                            $M_{H^{\pm}}$ [GeV]                                             $M_{H^{\pm}}$ [GeV]
Figure 2: The enhancement factors are presented in the parameter space of $M_{H^{\pm}},\leavevmode\nobreak\ \mu_{2}^{2}$. Charged Higgs masses are varied as $70$ GeV $\leq M_{H^{\pm}}\leq 1000$ GeV and $-200$ GeV $\leq\mu_{2}\leq 200$ GeV. We fix $\lambda_{2}=0.8$ and $M_{H}=150$ GeV for all cases. In the plots, we set $\sqrt{\hat{s}_{\gamma\gamma}}=500$ GeV (for the left panel-plots) and $\sqrt{\hat{s}_{\gamma\gamma}}=1000$ GeV (for the right panel-plots).
In Fig. 3, the enhancement factors for $\gamma\gamma\rightarrow hh,\leavevmode\nobreak\ HH$ are generated in the space of $M_{H^{\pm}},\leavevmode\nobreak\ \lambda_{2}$. The charged Higgs masses are varied as $70$ GeV $\leq M_{H^{\pm}}\leq 600$ GeV and $0\leq\lambda_{2}\leq 4$. We fix $\mu_{2}^{2}=200^{2}$ GeV2 and $M_{H}=150$ GeV for all cases. In the scatter plots, we set $\sqrt{\hat{s}_{\gamma\gamma}}=500$ (for all left panel plots) GeV and $\sqrt{\hat{s}_{\gamma\gamma}}=1000$ GeV (for all right panel plots). For the factors $\mu_{hh}^{\textrm{IHDM}}$ (as shown in all the above scatter plots), both the couplings $hH^{\pm}H^{\mp}$ and $hhH^{\pm}H^{\mp}$ are independent of $\lambda_{2}$. As a result, the factors only depend on $M_{H^{\pm}}$. For the factors $\mu_{HH}^{\textrm{IHDM}}$ (as shown in all the below scatter plots), it is found that the quadratic-coupling $HHH^{\pm}H^{\mp}$ depends on $\lambda_{2}$. As a result, the factors depend strongly on $\lambda_{2}$ and $M_{H^{\pm}}$. These massive contributions are mainly from the charged Higgs exchanging in the box diagrams.
$\mu_{hh}^{\textrm{IHDM}}$ $\mu_{hh}^{\textrm{IHDM}}$
Refer to caption Refer to caption
                                            $M_{H^{\pm}}$ [GeV]                                             $M_{H^{\pm}}$ [GeV]
$\mu_{HH}^{\textrm{IHDM}}$ $\mu_{HH}^{\textrm{IHDM}}$
Refer to caption Refer to caption
                                            $M_{H^{\pm}}$ [GeV]                                             $M_{H^{\pm}}$ [GeV]
Figure 3: The enhancement factors are scanned over the parameter space of $M_{H^{\pm}},\lambda_{2}$. Charged Higgs masses are in $70$ GeV $\leq M_{H^{\pm}}\leq 1000$ GeV and $0\leq\lambda_{2}\leq 4$. We fix $\mu_{2}^{2}=200^{2}$ GeV2 and $M_{H}=150$ GeV for all cases. In the plots, we set $\sqrt{\hat{s}_{\gamma\gamma}}=500$ GeV (for the above figures) and $\sqrt{\hat{s}_{\gamma\gamma}}=1000$ GeV (for the below figures), correspondingly.
3.2 THDM
The phenomenological results for the production processes $\gamma\gamma\rightarrow\phi_{i}\phi_{j}$ with CP-even Higgses $\phi_{i,j}\equiv h,H$ in the THDM are analyzed in the following subsection.
3.2.1 Production cross-sections
[…]
We inspect the data in the case of $M^{2}=0$. One notices that $\hat{\sigma}_{HH}$ becomes largest in the regions $\sqrt{\hat{s}_{\gamma\gamma}}\leq\sim 450$ GeV and they decrease rapidly in the regions $\sqrt{\hat{s}_{\gamma\gamma}}\geq 450$ GeV. Moreover, $\hat{\sigma}_{hh}$ in the SM and the THDM are dominant in the regions of $\sqrt{\hat{s}_{\gamma\gamma}}\geq 450$ GeV contrasted to the ones for $\gamma\gamma\rightarrow hH,\leavevmode\nobreak\ HH$ in the THDM. Among the mentioned cross-sections, the $hh$ production in the THDM is largest in this case.
[…]
$\hat{\sigma}_{\phi_{i}\phi_{j}}(M^{2}=0)$ [pb] $\hat{\sigma}_{\phi_{i}\phi_{j}}(M^{2}=200^{2})$ [pb]
Refer to caption Refer to caption
                                                     $\sqrt{\hat{s}_{\gamma\gamma}}$ [GeV]                                                      $\sqrt{\hat{s}_{\gamma\gamma}}$ [GeV]
$\hat{\sigma}_{\phi_{i}\phi_{j}}(M^{2}=500^{2})$ [pb] $\hat{\sigma}_{\phi_{i}\phi_{j}}(M^{2}=700^{2})$ [pb]
Refer to caption Refer to caption
                                                     $\sqrt{\hat{s}_{\gamma\gamma}}$ [GeV]                                                      $\sqrt{\hat{s}_{\gamma\gamma}}$ [GeV]
Figure 4: Cross-sections for $\gamma\gamma\rightarrow\phi_{i}\phi_{j}$ in the THDM, $\gamma\gamma\rightarrow hh$ in the SM are shown as functions of $\sqrt{\hat{s}_{\gamma\gamma}}$. In the plots, we vary $350$ GeV $\leq\sqrt{\hat{s}_{\gamma\gamma}}\leq 1500$ GeV. We select $M_{H^{\pm}}=300$ GeV, $t_{\beta}=5$ in this case. Moreover, we fix $M_{H}=150$ GeV, $M^{2}=0,200^{2},500^{2},700^{2}$ GeV2 and take $c_{\beta-\alpha}=+0.2$ and $s_{\beta-\alpha}=+\sqrt{1-c_{\beta-\alpha}^{2}}$, accordingly.
3.2.2 Enhancement factors
We focus on investigating the enhancement factors defined in Eq. 28 for $\gamma\gamma\rightarrow\phi_{i}\phi_{j}$ in the THDM. The factors are first studied over the parameter space of $M_{H^{\pm}},\leavevmode\nobreak\ t_{\beta}$ in this subsection. Two scenarios for $c_{\beta-\alpha}>0$ and $c_{\beta-\alpha}<0$ are examined in detail. In Figs. 5,  6, we fix $M^{2}=M_{H}^{2}=200^{2}$ GeV2. Moreover, we vary $100$ GeV $\leq M_{H^{\pm}}\leq 1000$ GeV and set $2\leq t_{\beta}\leq 10$ in the following plots. The factors $\mu_{\phi_{i}\phi_{j}}^{\textrm{THDM}}$ are generated at $\sqrt{\hat{s}_{\gamma\gamma}}=500$ GeV (for all the above scatter-plots) and examined at $\sqrt{\hat{s}_{\gamma\gamma}}=1000$ GeV (for all the below scatter-plots). In the left panel, we show the enhancement factors for $hh$ production. In the middle (right) panel, the enhancement factors for $hH$ and $(HH)$ production are presented, respectively.
In Fig. 5, the first scenario for $c_{\beta-\alpha}>0$ is explored. In this scenario, we take $c_{\beta-\alpha}=+0.2$ as an example and $s_{\beta-\alpha}=+\sqrt{1-c_{\beta-\alpha}^{2}}$, correspondingly. At $\sqrt{\hat{s}_{\gamma\gamma}}=500$ GeV, $\mu_{hh}^{\textrm{THDM}}$ changes from $1$ to $1.5$ for the entire range of $M_{H^{\pm}}$. The values of $\mu_{hh}^{\textrm{THDM}}$ are enhanced around the peak at $M_{H^{\pm}}=\sqrt{\hat{s}_{\gamma\gamma}}/2=250$ GeV. Predominantly, the factors are proportional to $t_{\beta}^{-1}$ in this case. Interestingly, we observe that the factors $\mu_{hH}^{\textrm{THDM}}$ change from 00 to $0.5$ for the entire range of $M_{H^{\pm}}$. The suppressed values of $\mu_{hH}^{\textrm{THDM}}$ are expected, as explained in the previous paragraph, due to the $Z_{2}$-symmetry. The $\mu_{hH}^{\textrm{THDM}}$ exhibit the same behavior as $\mu_{hh}^{\textrm{THDM}}$, in that they are inversely proportional to $t_{\beta}$. On the other hand, the enhancement factors for $HH$ productions in the THDM are strongly dependent on the charged Higgs mass but change slightly with $t_{\beta}$. In the entire range of $M_{H^{\pm}}$, the factors $\mu_{HH}^{\textrm{THDM}}$ range from $0.3$ to $1.2$.
At $\sqrt{\hat{s}_{\gamma\gamma}}=1000$ GeV, the factors $\mu_{hh}^{\textrm{THDM}}$ become the largest at the peak at $M_{H^{\pm}}=\sqrt{\hat{s}_{\gamma\gamma}}/2=500$ GeV. Around the peak, $\mu_{hh}^{\textrm{THDM}}$ changes from $1.2$ to $1.8$. Beyond the peak, the factors change from $1.0$ to $1.2$ in the entire range of $M_{H^{\pm}}$. It is important to note that $\mu_{hh}^{\textrm{THDM}}$ slightly changes with $t_{\beta}$. For $HH$ productions, $\mu_{HH}^{\textrm{THDM}}$ varies from $0.4$ to $2.5$ around the peak (at $M_{H^{\pm}}=500$ GeV) regions. It is noted that $\mu_{HH}^{\textrm{THDM}}$ slightly changes with $t_{\beta}$. Otherwise, $\mu_{hH}^{\textrm{THDM}}$ is much smaller than $1$ and is inversely proportional to $t_{\beta}$.
$\mu_{hh}^{\textrm{THDM}}$ $\mu_{hH}^{\textrm{THDM}}$ $\mu_{HH}^{\textrm{THDM}}$
Refer to caption Refer to caption Refer to caption
                           $M_{H^{\pm}}$ [GeV]                            $M_{H^{\pm}}$ [GeV]                            $M_{H^{\pm}}$ [GeV]
$\mu_{hh}^{\textrm{THDM}}$ $\mu_{hH}^{\textrm{THDM}}$ $\mu_{HH}^{\textrm{THDM}}$
Refer to caption Refer to caption Refer to caption
                           $M_{H^{\pm}}$ [GeV]                            $M_{H^{\pm}}$ [GeV]                            $M_{H^{\pm}}$ [GeV]
Figure 5: The enhancement factors are presented in the parameter space of $M_{H^{\pm}},\leavevmode\nobreak\ t_{\beta}$. In the plots, we fix $M^{2}=M_{H}^{2}=200^{2}$ GeV2 and take $c_{\beta-\alpha}=+0.2$ and $s_{\beta-\alpha}=+\sqrt{1-c_{\beta-\alpha}^{2}}$, accordingly. In these plots, we set $\sqrt{\hat{s}_{\gamma\gamma}}=500$ GeV (for the above plots) and $\sqrt{\hat{s}_{\gamma\gamma}}=1000$ GeV (for the below plots), respectively.
Another scenario for $c_{\beta-\alpha}<0$ is considered for examining how the factors are affected by setting a different sign of $c_{\beta-\alpha}$ in this work. In Fig. 6, we take $c_{\beta-\alpha}=-0.2$ for an example and $s_{\beta-\alpha}=+\sqrt{1-c_{\beta-\alpha}^{2}}$, accordingly. At $\sqrt{\hat{s}_{\gamma\gamma}}=500$ GeV, it is interesting to observe that $\mu_{hh}^{\textrm{THDM}}$ shows different behavior in comparison with the previous scenario. At this CoM energy, the factors $\mu_{hh}^{\textrm{THDM}}$ can reach up to $1.5$ in the low region of $M_{H^{\pm}}<200$ GeV. They then decrease to around $0.9$ when $M_{H^{\pm}}>200$ GeV. There is no peak of the factors observed in this scenario because the contributions of singly charged Higgs exchanging in the one-loop triangle diagrams may cancel out with the ones from the one-loop box diagrams in this scenario. Surprisingly, we find that the factors $\mu_{hh}^{\textrm{THDM}}$ are proportional to $t_{\beta}$ in this scenario. For the $hH$ productions, the factors are suppressed and they are in the range of [0.025,0.3][$\sim$ 0.025$\sim$ 0.3][ ∼ 0.025 , ∼ 0.3 ]. They are sensitive to $t_{\beta}^{-1}$ in all ranges of charged Higgs mass. In $HH$ productions, it is found that the factors develop to the peak around $M_{H^{\pm}}=500$ GeV. They reach a factor of $2.5$ around the peak and they are in the ranges of [0.5,1.7][$\sim$ 0.5$\sim$ 1.7][ ∼ 0.5 , ∼ 1.7 ] beyond the peak regions. In all ranges of $M_{H^{\pm}}$, the factors $\mu_{HH}^{\textrm{THDM}}$ are proportional to $t_{\beta}^{-1}$ in this scenario.
The survey for all the enhancement factors at $\sqrt{\hat{s}_{\gamma\gamma}}=1000$ GeV is presented in the next paragraphs. The factors $\mu_{hh}^{\textrm{THDM}}$ are large in the regions ($M_{H^{\pm}}\leq 200$ GeV) and they can reach up to $1.5$. They then decrease rapidly to the regions around $M_{H^{\pm}}\sim 300$ GeV and develop a peak at $M_{H^{\pm}}=\sqrt{\hat{s}_{\gamma\gamma}}/2=500$ GeV. Around the peak, the enhancement factor is about $1.2$. In other ranges of charged Higgs mass, $\mu_{hh}^{\textrm{THDM}}\sim 0.9$. One also observes that $\mu_{hh}^{\textrm{THDM}}$ is inversely proportional to $t_{\beta}$ in this scenario. In $hH$ productions, the factors increase to the peak at $M_{H^{\pm}}=\sqrt{\hat{s}_{\gamma\gamma}}/2=500$ GeV and they are about $0.3$ around the peak. In all regions of $M_{H^{\pm}}$, the mentioned factors are in the ranges of [0.025,0.3][$\sim$ 0.025$\sim$ 0.3][ ∼ 0.025 , ∼ 0.3 ] and they are inversely proportional to $t_{\beta}$. In the last case, it is found that the factors $\mu_{HH}^{\textrm{THDM}}$ show the same behavior as the previous scenario. They are in the ranges of [0.5,4][$\sim$ 0.5$\sim$ 4][ ∼ 0.5 , ∼ 4 ] in all regions of $M_{H^{\pm}}$. However, the factors $\mu_{HH}^{\textrm{THDM}}$ depend only slightly on $t_{\beta}$ in this scenario.
$\mu_{hh}^{\textrm{THDM}}$ $\mu_{hH}^{\textrm{THDM}}$ $\mu_{HH}^{\textrm{THDM}}$
Refer to caption Refer to caption Refer to caption
                           $M_{H^{\pm}}$ [GeV]                            $M_{H^{\pm}}$ [GeV]                            $M_{H^{\pm}}$ [GeV]
$\mu_{hh}^{\textrm{THDM}}$ $\mu_{hH}^{\textrm{THDM}}$ $\mu_{HH}^{\textrm{THDM}}$
Refer to caption Refer to caption Refer to caption
                           $M_{H^{\pm}}$ [GeV]                            $M_{H^{\pm}}$ [GeV]                            $M_{H^{\pm}}$ [GeV]
Figure 6: The enhancement factors are presented in the parameter space $M_{H^{\pm}},\leavevmode\nobreak\ t_{\beta}$. In the plots, we consider the scenario for $c_{\beta-\alpha}=-0.2<0$ and $s_{\beta-\alpha}=+\sqrt{1-c_{\beta-\alpha}^{2}}$, correspondingly. We also fix $M^{2}=M_{H}^{2}=200^{2}$ GeV2 and set $\sqrt{\hat{s}_{\gamma\gamma}}=500$ (for all the above-plots) GeV and $\sqrt{\hat{s}_{\gamma\gamma}}=1000$ GeV (for all the below-plots).
[…]
In Fig. 7, the first scenario of $c_{\beta-\alpha}>0$ is examined. For this case, we take $c_{\beta-\alpha}=+0.2$ as an example and $s_{\beta-\alpha}=+\sqrt{1-c_{\beta-\alpha}^{2}}$, accordingly. For $hh$ production at $\sqrt{\hat{s}_{\gamma\gamma}}=500$ GeV, we observe the peak of $\mu_{hh}^{\textrm{THDM}}$ at $M_{H^{\pm}}=250$ GeV, which is corresponding to the threshold of cross-sections for $hh$ in the THDM at the peak $\sqrt{\hat{s}_{\gamma\gamma}}=2M_{H^{\pm}}$. Around the peak, $\mu_{hh}^{\textrm{THDM}}$ varies from $1.0$ to $1.8$. Above the peak regions, the enhancement factors tend to $1.2$ and depend slightly on $M^{2}$. For $hH$ production, $\mu_{hH}^{\textrm{THDM}}$ is more sensitive to $M^{2}$ below the peak regions. The factors are in the range of $[0.07,1.5]$ above the peak regions. Around the peak, $\mu_{hH}^{\textrm{THDM}}$ can reach up to $0.3$. We also find the same behavior for $\mu_{HH}^{\textrm{THDM}}$. The factors for $HH$ production are large in the low regions of $M_{H^{\pm}}$ and around the peak $M_{H^{\pm}}=250$ GeV. They are in the range of $[0.7,2.4]$ above the peak regions. Generally, we observe that $\mu_{\phi_{i}\phi_{j}}^{\textrm{THDM}}$ is proportional to $1/M$ at this CoM energy.
At $\sqrt{\hat{s}_{\gamma\gamma}}=1000$ GeV, we also find that $\mu_{hh}^{\textrm{THDM}}$ develops to the peak at $M_{H^{\pm}}=500$ GeV, where the factors can reach up to $2.2$ and decrease rapidly beyond the peak. The factors depend slightly on $M^{2}$ and tend to $1$ beyond the peak regions. For $hH$ production, $\mu_{hH}^{\textrm{THDM}}$ is sensitive to $M^{2}$ in the peak regions. They tend to $0.05$ and are slightly dependent on $M^{2}$ in the regions above the peak. For $HH$ production, the factors become large in the regions below the peak and they are inversely proportional to $M^{-1}$. Around the peak, the factors are enhanced by large values of $M^{2}$. Above the peak regions, $\mu_{HH}^{\textrm{THDM}}$ varies around $1.0$.
$\mu_{hh}^{\textrm{THDM}}$ $\mu_{hH}^{\textrm{THDM}}$ $\mu_{HH}^{\textrm{THDM}}$
Refer to caption Refer to caption Refer to caption
                           $M_{H^{\pm}}$ [GeV]                            $M_{H^{\pm}}$ [GeV]                            $M_{H^{\pm}}$ [GeV]
$\mu_{hh}^{\textrm{THDM}}$ $\mu_{hH}^{\textrm{THDM}}$ $\mu_{HH}^{\textrm{THDM}}$
Refer to caption Refer to caption Refer to caption
                           $M_{H^{\pm}}$ [GeV]                            $M_{H^{\pm}}$ [GeV]                            $M_{H^{\pm}}$ [GeV]
Figure 7: The enhancement factors are presented in the parameter space of $M_{H^{\pm}},\leavevmode\nobreak\ M^{2}$. In the plots, we take $t_{\beta}=5$ and consider the first scenario of $c_{\beta-\alpha}=+0.2$ and $s_{\beta-\alpha}=+\sqrt{1-c_{\beta-\alpha}^{2}}$, accordingly. In these plots, we set $\sqrt{\hat{s}_{\gamma\gamma}}=500$ GeV (for the above Figures) and $\sqrt{\hat{s}_{\gamma\gamma}}=1000$ GeV (for the below Figures).
[…]
We next comment on the physical results at $\sqrt{\hat{s}_{\gamma\gamma}}=1000$ GeV. The factors $\mu_{hh}^{\textrm{THDM}}$ are enhanced by the small values of $M$ in the low regions of charged Higgs mass and they can reach up to $1.2$. They tend to $2.5$ around the peak. The factors are then varied around $\sim 0.9$. In general, the factors depend on $M^{-1}$. In the $hH$ production, $\mu_{hH}^{\textrm{THDM}}$ is more sensitive to $M^{-1}$ around the peak regions. They then tend to $0.2$ in the high mass regions of singly charged Higgs. For $HH$ production, the factors $\mu_{HH}^{\textrm{THDM}}$ strongly depend on $M^{-1}$. At the peak, the factors are enhanced by the large value of $M$. Above the peak regions, $\mu_{HH}^{\textrm{THDM}}$ tend to $\sim 1$.
$\mu_{hh}^{\textrm{THDM}}$ $\mu_{hH}^{\textrm{THDM}}$ $\mu_{HH}^{\textrm{THDM}}$
Refer to caption Refer to caption Refer to caption
                           $M_{H^{\pm}}$ [GeV]                            $M_{H^{\pm}}$ [GeV]                            $M_{H^{\pm}}$ [GeV]
$\mu_{hh}^{\textrm{THDM}}$ $\mu_{hH}^{\textrm{THDM}}$ $\mu_{HH}^{\textrm{THDM}}$
Refer to caption Refer to caption Refer to caption
                           $M_{H^{\pm}}$ [GeV]                            $M_{H^{\pm}}$ [GeV]                            $M_{H^{\pm}}$ [GeV]
Figure 8: The enhancement factors are presented in the parameter space of $M_{H^{\pm}},\leavevmode\nobreak\ M^{2}$. In the plots, we take $t_{\beta}=5$ and $c_{\beta-\alpha}=-0.2$ and $s_{\beta-\alpha}=+\sqrt{1-c_{\beta-\alpha}^{2}}$, accordingly. In these plots, we set $\sqrt{\hat{s}_{\gamma\gamma}}=500$ GeV (the above Figures) and $\sqrt{\hat{s}_{\gamma\gamma}}=1000$ GeV (the below Figures).
4 Conclusions
In this paper, we have presented the phenomenological results for one-loop induced processes $\gamma\gamma\rightarrow\phi_{i}\phi_{j}$ with CP-even Higgses $\phi_{i,j}\equiv h,\leavevmode\nobreak\ H$ at high-energy photon-photon collisions in the IHDM and the THDM. In the phenomenological results, we have shown the cross sections at several center-of-mass energies. The results show that cross sections for the computed processes in the models under investigation are enhanced around the threshold of charged Higgs pair production ($\sim 2M_{H^{\pm}}$). Furthermore, the enhancement factors for the processes are examined in the parameter space of the models under consideration. In the IHDM, the factors are studied in the parameter space of $(M_{H^{\pm}},\leavevmode\nobreak\ \mu_{2}^{2})$ and $(M_{H^{\pm}},\leavevmode\nobreak\ \lambda_{2})$. In the THDM, the factors are analyzed in the planes of $(M_{H^{\pm}},\leavevmode\nobreak\ t_{\beta})$ and $(M_{H^{\pm}},\leavevmode\nobreak\ M^{2})$. Two scenarios of $c_{\beta-\alpha}>0$ and $c_{\beta-\alpha}<0$ have been studied in further detail. The factors give a different behavior when considering these scenarios. As a result, discriminations for the above-mentioned scenarios can be performed at future colliders. The dependence of the cross section for the process $\gamma\gamma\rightarrow hH$ on $m_{12}^{2}$ provides a potential probe of the soft $Z_{2}$-breaking scale in the Two Higgs Doublet Model.
Acknowledgment:  This research is funded by Vietnam National Foundation for Science and Technology Development (NAFOSTED) under the grant number $103.01$-$2023.16$.
Appendix A: Effective Lagrangian in the IHDM
The kinematic terms of Lagrangian in the IHDM can be expanded as follows:
$\mathcal{L}_{K}^{\text{IHDM}}$ $\supset$ $\frac{2M_{W}^{2}}{v}W^{\pm}_{\mu}W^{\mp,\mu}h+\frac{M_{Z}^{2}}{v}hZ_{\mu}Z^{\mu}+i\frac{M_{Z}c_{2W}}{v}Z^{\mu}(H^{\mp}\partial_{\mu}H^{\pm}-H^{\pm}\partial_{\mu}H^{\mp})$ (29)
$+i\frac{M_{Z}s_{2W}}{v}A^{\mu}(H^{\mp}\partial_{\mu}H^{\pm}-H^{\pm}\partial_{\mu}H^{\mp})+i\frac{2M_{Z}^{2}c_{W}^{2}s_{W}}{v}A^{\mu}W_{\mu}^{\pm}G^{\pm}$
$+i\frac{M_{W}}{v}(HW^{\mp,\mu}\partial_{\mu}H^{\pm}-H^{\pm}W^{\mp,\mu}\partial_{\mu}H-HW^{\pm}_{\mu}\partial^{\mu}H^{\mp}+H^{\mp}W^{\pm}_{\mu}\partial^{\mu}H)$
$+\frac{M_{Z}^{2}s_{2W}^{2}}{v^{2}}A_{\mu}A^{\mu}H^{\pm}H^{\mp}+\frac{2M_{Z}^{2}c_{W}^{2}s_{W}}{v^{2}}HH^{\pm}W_{\mu}^{\mp}A^{\mu}+\frac{2M_{W}^{2}}{v^{2}}W^{\pm}_{\mu}W^{\mp,\mu}hh$
$+\frac{2M_{W}^{2}}{v^{2}}W^{\pm}_{\mu}W^{\mp,\mu}HH+\frac{M_{Z}^{2}s_{2W}^{2}}{v^{2}}A_{\mu}A^{\mu}G^{\pm}G^{\mp}$
$+i\frac{M_{Z}s_{2W}}{v}A^{\mu}(G^{\mp}\partial_{\mu}G^{\pm}-G^{\pm}\partial_{\mu}G^{\mp})+\cdots.$
We also expand the scalar Higgs potential of the IHDM and collect the terms involving to Higgs self-coupling as follows:
$-\mathcal{V}_{\text{IHDM}}(\phi_{1},\phi_{2})$ $\supset$ $-\frac{3M_{h}^{2}}{v}hhh+\frac{2(\mu_{2}^{2}-M_{H}^{2})}{v}hHH+\frac{2(\mu_{2}^{2}-M_{H^{\pm}}^{2})}{v}hH^{\pm}H^{\mp}$ (30)
$+\frac{M_{H^{\pm}}^{2}-M_{H}^{2}}{v}HH^{\pm}{G^{\mp}}+\frac{2(\mu_{2}^{2}-M_{H^{\pm}}^{2})}{v^{2}}hhH^{\pm}H^{\mp}-2\lambda_{2}HHH^{\pm}H^{\mp}$
$-\frac{M_{h}^{2}}{v^{2}}hhG^{\pm}G^{\mp}+\frac{2(\mu_{2}^{2}-M_{H^{\pm}}^{2})}{v^{2}}HHG^{\pm}G^{\mp}+\cdots.$
Appendix B: Effective Lagrangian in the THDM
We expand the kinematic terms of Lagrangian in the THDM as follows:
$\mathcal{L}_{K}^{\text{THDM}}$ $\supset$ $\frac{2M_{W}^{2}}{v}s_{\beta-\alpha}hW_{\mu}^{\pm}W^{\mp,\mu}+\frac{2M_{W}^{2}}{v}c_{\alpha-\beta}HW_{\mu}^{\pm}W^{\mp,\mu}+\frac{M_{Z}^{2}}{v}s_{\beta-\alpha}hZ_{\mu}Z^{\mu}$ (31)
$+\frac{M_{Z}^{2}}{v}c_{\alpha-\beta}HZ_{\mu}Z^{\mu}+i\frac{M_{Z}c_{2W}}{v}Z^{\mu}(H^{\mp}\partial_{\mu}{H^{\pm}}-H^{\pm}\partial_{\mu}H^{\mp})$
$+i\frac{M_{Z}s_{2W}}{v}A^{\mu}(H^{\mp}\partial_{\mu}{H^{\pm}}-H^{\pm}\partial_{\mu}H^{\mp})+\frac{4M_{W}^{2}s_{W}^{2}}{v^{2}}H^{\pm}H^{\mp}A_{\mu}{A^{\mu}}$
$-i\frac{M_{W}s_{\beta-\alpha}}{v}(HW^{\mp,\mu}\partial_{\mu}H^{\pm}-HW^{\pm,\mu}\partial_{\mu}H^{\mp}+H^{\mp}W^{\pm,\mu}\partial_{\mu}H-H^{\pm}W^{\mp,\mu}\partial_{\mu}H)$
$-i\frac{M_{W}c_{\beta-\alpha}}{v}(-hW^{\mp,\mu}\partial_{\mu}H^{\pm}+hW^{\pm,\mu}\partial_{\mu}H^{\mp}-H^{\mp}W^{\pm,\mu}\partial_{\mu}h+H^{\pm}W^{\mp,\mu}\partial_{\mu}h)$
$+\frac{2M_{W}^{2}s_{W}c_{\beta-\alpha}}{v^{2}}hH^{\mp}W_{\mu}^{\pm}A^{\mu}-\frac{2M_{W}^{2}s_{W}s_{\beta-\alpha}}{v^{2}}HH^{\mp}W^{\pm}_{\mu}A^{\mu}$
$+\dfrac{2M_{W}^{2}}{v^{2}}W^{\pm}_{\mu}W^{\mp,\mu}HH+\frac{2M_{W}^{2}}{v^{2}}W^{\pm}_{\mu}W^{\mp,\mu}hh+\cdots$
Expanding the scalar potential in the THDM, we then collect the terms involving to the Higgs self-couplings as
$-\mathcal{V}_{\text{THDM}}(\phi_{1},\phi_{2})$ $\supset$ $-\lambda_{hHH}hHH-\lambda_{Hhh}Hhh-\lambda_{hH^{\pm}H^{\mp}}hH^{\pm}H^{\mp}$ (32)
$-\lambda_{hH^{\pm}H^{\mp}}HH^{\pm}H^{\mp}-\lambda_{HhH^{\pm}H^{\mp}}HhH^{\pm}H^{\mp}+\cdots.$
All coefficients of the mentioned couplings are shown explicitly in terms of physical parameters as follows:
$-\lambda_{hHH}$ $=$ $\frac{3\lambda_{1}v}{2}s_{\alpha}c_{\alpha}^{2}c_{\beta}-\frac{3\lambda_{2}v}{2}s_{\beta}s_{\alpha}^{2}c_{\alpha}-\frac{\lambda_{345}}{2}v[c_{\beta}(2s_{\alpha}c_{\alpha}^{2}-s_{\alpha}^{3})+s_{\beta}(c_{\alpha}^{3}-2s_{\alpha}^{2}c_{\alpha})]$ (33)
$=$ $\frac{s_{\alpha-\beta}[s_{2\alpha}(3M^{2}-M_{h}^{2}-2M_{H}^{2})+M^{2}s_{2\beta}]}{v\;s_{2\beta}},$ (34)
$-\lambda_{Hhh}$ $=$ $-\frac{3\lambda_{1}v}{2}c_{\beta}c_{\alpha}s_{\alpha}^{2}-\frac{3\lambda_{2}v}{2}s_{\beta}c_{\alpha}^{2}s_{\alpha}-\frac{\lambda_{345}}{2}v[s_{\beta}(s_{\alpha}^{3}-2c_{\alpha}^{2}s_{\alpha})-c_{\beta}(2c_{\alpha}s_{\alpha}^{2}-c_{\alpha}^{3})]$ (35)
$=$ $\frac{c_{\alpha-\beta}[s_{2\alpha}(3M^{2}-M_{H}^{2}-2m_{h}^{2})-M^{2}s_{2\beta}]}{v\;s_{2\beta}},$ (36)
$-\lambda_{HH^{\pm}H^{\mp}}$ $=$ $-\lambda_{1}vc_{\beta}c_{\alpha}s_{\beta}^{2}-\lambda_{2}vs_{\beta}s_{\alpha}c_{\beta}^{2}-\lambda_{3}v(s_{\beta}s_{\alpha}s_{\beta}^{2}+c_{\beta}c_{\alpha}c_{\beta}^{2})+\frac{\lambda_{45}}{2}vs_{(2\beta)}\;s_{\beta+\alpha}$ (37)
$=$ $\frac{s_{\alpha+\beta}(4M^{2}-3M_{H}^{2}-2M_{H^{\pm}}^{2})+(2M_{H^{\pm}}^{2}-M_{H}^{2})s_{\alpha-3\beta}}{2vs_{(2\beta)}},$ (38)
$-\lambda_{hH^{\pm}H^{\mp}}$ $=$ $\lambda_{1}vc_{\beta}s_{\alpha}s_{\beta}^{2}-\lambda_{2}vs_{\beta}c_{\alpha}c_{\beta}^{2}-\lambda_{3}v(s_{\beta}c_{\alpha}s_{\beta}^{2}-c_{\beta}s_{\alpha}c_{\beta}^{2})+\frac{\lambda_{45}}{2}v\;s_{(2\beta)}\;c_{(\beta+\alpha)}$ (39)
$=$ $\frac{c_{\alpha+\beta}(4M^{2}-3M_{h}^{2}-2M_{H^{\pm}}^{2})+(2M_{H^{\pm}}^{2}-M_{h}^{2})c_{(\alpha-3\beta)}}{2vs_{2\beta}},$ (40)
and
$-\lambda_{HhH^{\pm}H^{\mp}}$ $=$ $\lambda_{1}s_{\beta}^{2}s_{\alpha}c_{\alpha}-\lambda_{2}c_{\beta}^{2}s_{\alpha}c_{\alpha}+\lambda_{3}s_{\alpha}c_{\alpha}c_{2\beta}+(\lambda_{4}+\lambda_{5})s_{\beta}c_{\beta}c_{2\alpha}$ (42)
$=$ $\frac{s_{2\alpha}(3c_{2\alpha}+c_{2(\alpha-2\beta)}-4c_{2\beta})}{4v^{2}s_{2\beta}^{2}}M_{H}^{2}-\frac{s_{2\alpha}(3c_{2\alpha}+c_{2(\alpha-2\beta)}+4c_{2\beta})}{4v^{2}s_{2\beta}^{2}}M_{h}^{2}$
$+\frac{s_{2(\alpha-\beta)}}{v^{2}}M_{H^{\pm}}^{2}+\frac{(s_{2(\alpha-3\beta)}+2s_{2(\alpha-\beta)}+5s_{2(\alpha+\beta)})}{4v^{2}s_{2\beta}^{2}}M^{2},$
$-\lambda_{HhG^{\pm}G^{\mp}}$ $=$ $\lambda_{1}c_{\beta}^{2}s_{\alpha}c_{\alpha}-\lambda_{2}s_{\beta}^{2}s_{\alpha}c_{\alpha}-\lambda_{3}s_{\alpha}c_{\alpha}c_{2\beta}-(\lambda_{4}+\lambda_{5})s_{\beta}c_{\beta}c_{2\alpha}$
$=$ $\frac{1}{2v^{2}s_{2\beta}}s_{2(\alpha-\beta)}[(M_{h}^{2}-M_{H}^{2})s_{2\alpha}+2(M^{2}-M_{H^{\pm}}^{2})s_{2\beta}].$ (43)
Furthermore, we have the following couplings:
$-\lambda_{hhh}$ $=$ $3v\Big{[}\lambda_{1}s_{\alpha}^{3}c_{\beta}-\lambda_{2}c_{\alpha}^{3}s_{\beta}+(\lambda_{3}+\lambda_{4}+\lambda_{5})(s_{\alpha}c_{\beta}c_{\alpha}^{2}-c_{\alpha}s_{\beta}s_{\alpha}^{2})\Big{]}$ (44)
$=$ $\dfrac{3e}{4M_{W}s_{W}s_{2\beta}}\Big{[}M^{2}c_{\alpha-3\beta}+(M^{2}-M_{h}^{2})c_{3\alpha-\beta}+(2M^{2}-3M_{h}^{2})c_{\alpha+\beta}\Big{]},$
$-\lambda_{HHH}$ $=$ $-3v\Big{[}\lambda_{1}c_{\alpha}^{3}c_{\beta}+\lambda_{2}s_{\alpha}^{3}s_{\beta}+(\lambda_{3}+\lambda_{4}+\lambda_{5})(c_{\alpha}c_{\beta}s_{\alpha}^{2}+s_{\alpha}s_{\beta}c_{\alpha}^{2})\Big{]}$ (45)
$=$ $\dfrac{3e}{4M_{W}s_{W}s_{2\beta}}\Big{[}M^{2}s_{\alpha-3\beta}+(M_{H}^{2}-M^{2})s_{3\alpha-\beta}+(2M^{2}-3M_{H}^{2})s_{\alpha+\beta}\Big{]},$
$-\lambda_{HHH^{\pm}{H^{\mp}}}$ $=$ $-\lambda_{1}c_{\alpha}^{2}s_{\beta}^{2}-\lambda_{2}s_{\alpha}^{2}c_{\beta}^{2}-\lambda_{3}(c_{\alpha}^{2}c_{\beta}^{2}+s_{\alpha}^{2}s_{\beta}^{2})+(\lambda_{4}+\lambda_{5})c_{\alpha}s_{\alpha}s_{2\beta}$ (46)
$=$ $-\dfrac{2c_{\alpha-\beta}^{2}}{v^{2}}M_{H^{\pm}}^{2}-\frac{s_{2\alpha}[3s_{2\alpha}+s_{2(\alpha-2\beta)}-2s_{2\beta}]}{4v^{2}s_{2\beta}^{2}}M_{h}^{2}$
$-\dfrac{[c_{\alpha}^{4}+c_{\alpha}^{3}s_{\alpha}\cot^{3}\beta+c_{\alpha}s_{\alpha}^{3}\cot\beta+s_{\alpha}^{4}\cot^{4}\beta]\tan^{2}\beta}{v^{2}}M_{H}^{2}$
$+\dfrac{s_{\beta}[4c_{\alpha}c_{\beta}s_{\alpha}+(1+\cot^{4}\beta)s_{\alpha}^{2}s_{\beta}+c_{\alpha}^{2}(1+\cot^{4}\beta)s_{\beta}\tan^{2}\beta]}{v^{2}}M^{2},$
$-\lambda_{hhH^{\pm}H^{\mp}}$ $=$ $-\lambda_{1}s_{\alpha}^{2}s_{\beta}^{2}-\lambda_{2}c_{\alpha}^{2}c_{\beta}^{2}-\lambda_{3}(s_{\alpha}^{2}c_{\beta}^{2}+c_{\alpha}^{2}s_{\beta}^{2})-(\lambda_{4}+\lambda_{5})c_{\alpha}s_{\alpha}s_{2\beta}$ (47)
$=$ $-\dfrac{2s_{\alpha-\beta}^{2}}{v^{2}}M_{H^{\pm}}^{2}-\dfrac{s_{2\alpha}[3s_{2\alpha}+s_{2(\alpha-2\beta)}+2s_{2\beta}]}{4v^{2}s_{2\beta}^{2}}M_{H}^{2}$
$+\dfrac{(-c_{\alpha}^{4}\cot^{2}{\beta}+c_{\alpha}s_{\alpha}^{3}\cot\beta+c_{\alpha}^{3}s_{\alpha}\tan\beta-s_{\alpha}^{4}\tan^{2}\beta)}{v^{2}}M_{h}^{2}$
$+\dfrac{s_{\beta}c_{\beta}[-4c_{\alpha}s_{\alpha}+(1+\cot^{4}\beta)s_{\alpha}^{2}\tan^{3}{\beta}+c_{\alpha}^{2}(\cot^{3}\beta+\tan\beta)]}{v^{2}}M^{2},$
$-\lambda_{hhG^{\pm}G^{\mp}}$ $=$ $-\lambda_{1}s_{\alpha}^{2}c_{\beta}^{2}-\lambda_{2}c_{\alpha}^{2}s_{\beta}^{2}-\lambda_{3}(c_{\alpha}^{2}c_{\beta}^{2}+s_{\alpha}^{2}s_{\beta}^{2})+(\lambda_{4}+\lambda_{5})c_{\alpha}s_{\alpha}s_{2\beta}$ (48)
$=$ $\dfrac{2c_{\alpha-\beta}^{2}}{v^{2}}(M^{2}-M_{H^{\pm}}^{2})-\dfrac{c_{\alpha-\beta}^{2}s_{2\alpha}}{v^{2}s_{2\beta}}M_{H}^{2}+\dfrac{-3s_{2\beta}+2s_{2\alpha}+s_{4\alpha-2\beta}}{4v^{2}s_{2\beta}}M_{h}^{2},$
$-\lambda_{HHG^{\pm}G^{\mp}}$ $=$ $-\lambda_{1}c_{\alpha}^{2}c_{\beta}^{2}-\lambda_{2}s_{\alpha}^{2}s_{\beta}^{2}-\lambda_{3}(s_{\alpha}^{2}c_{\beta}^{2}+c_{\alpha}^{2}s_{\beta}^{2})-(\lambda_{4}+\lambda_{5})c_{\alpha}s_{\alpha}s_{2\beta}$ (49)
$=$ $\dfrac{2s_{\alpha-\beta}^{2}}{v^{2}}(M^{2}-M_{H^{\pm}}^{2})+\dfrac{s_{\alpha-\beta}^{2}s_{2\alpha}}{v^{2}s_{2\beta}}M_{h}^{2}+\dfrac{-3s_{2\beta}-2s_{2\alpha}+s_{4\alpha-2\beta}}{4v^{2}s_{2\beta}}M_{H}^{2}.$
Additionally, we also derive the couplings relating to Goldstone bosons as follows:
$\mathcal{L}_{K}^{\text{THDM}}$ $\supset$ $\frac{2M_{W}^{2}s_{W}}{v}A_{\mu}W^{\pm,\mu}G^{\mp}+\frac{4M_{W}^{2}s_{W}^{2}}{v^{2}}A_{\mu}A^{\mu}G^{\pm}{G^{\mp}}$ (50)
$+i\frac{2M_{W}s_{W}}{v}A^{\mu}(G^{\mp}\partial_{\mu}G^{\pm}-G^{\pm}\partial_{\mu}G^{\mp})+\cdots$
From scalar potential, we have
$-\mathcal{V}(\phi_{1},\phi_{2})$ $\supset$ $-\lambda_{hH^{\pm}G^{\mp}}hH^{\pm}G^{\mp}-\lambda_{HH^{\pm}G^{\mp}}HH^{\pm}G^{\mp}+\cdots$ (51)
where the coefficients of the couplings are given by
$-\lambda_{hH^{\pm}G^{\mp}}$ $=$ $\frac{ec_{\alpha-\beta}}{2M_{W}s_{W}}(M_{H^{\pm}}^{2}-M_{h}^{2}),$ (52)
$-\lambda_{HH^{\pm}G^{\mp}}$ $=$ $\frac{es_{\alpha-\beta}}{2M_{W}s_{W}}(M_{H^{\pm}}^{2}-M_{H}^{2}).$ (53)